### v2

# Discrete-to-continuum limits of semilinear stochastic evolution equations in Banach spaces

Abstract.
We study the convergence of semilinear parabolic stochastic evolution equations, posed on a sequence of Banach spaces approximating a limiting space and driven by additive white noise projected onto the former spaces. Under appropriate uniformity and convergence conditions on the linear operators, nonlinear drifts and initial data, we establish convergence of the associated mild solution processes when lifted to a common state space. Our framework is applied to the case where the limiting problem is a stochastic partial differential equation whose linear part is a generalized Whittle–Matérn operator on a manifold $\mathcal{M}$, discretized by a sequence of graphs constructed from a (random) point cloud. In this setting we obtain various discrete-to-continuum convergence results for solutions lifted to $L^{q}(\mathcal{M})$ for $q\in[2,\infty]$, one of which recovers the $L^{\infty}$-convergence of a finite-difference discretization of certain (fractional) stochastic Allen–Cahn equations.
Key words and phrases:
Discrete-to-continuum convergence, stochastic evolution equations, generalized Whittle–Matérn operators, geometric graphs
2020 Mathematics Subject Classification:
Primary: 60H15, 35K58, 65M12; secondary: 35R11, 35R01, 35R02.
1. Introduction
1.1. Background and motivation
We establish discrete-to-continuum limits of stochastic evolution equations of the form (1.1), i.e., semilinear parabolic stochastic partial diffential equations (SPDEs) driven by Gaussian white noise. Such SPDEs of evolution play an important role in the modeling of physical and other systems, such as fluid dynamics [5, 26, 31, 62], quantum optics [13], phase separation [18], diffusion in random media [44, 53], and population dynamics [87]. Given their significance, there has been a considerable interest in the analysis and numerical analysis of SPDEs; see the introductory textbooks [60] and [61], respectively.
We consider the convergence of a sequence of abstract continuous-time equations, each posed on a different Banach space in order to model the approximation of an evolution SPDE by equations that are continuous in time and discrete in space. This framework covers a typical setting where the spatial domains are finite graphs and the limiting differential operator in space is approximated by the corresponding graphical variants. If the finite graphs approximate an underlying manifold, then the graphical differential operators are related to a finite-difference approximation of the differential operator. As an example, we study a class of semilinear SPDEs whose linear part is given by a fractional-order elliptic differential operator on a manifold (known as a Whittle–Matérn operator), discretized by a sequence of graphs constructed from a (possibly randomly sampled) cloud of points. Linear equations of this type have previously been studied in the context of statistics and machine learning [65, 71, 72]. By verifying in detail that the hypotheses of our abstract framework are satisfied in this situation, we establish the discrete-to-continuum convergence of this scheme. Although the main advantage of our general results is their applicability to highly unstructured discretizations, we additionally show that they recover the $L^{\infty}$-convergence of finite-difference discretizations of the fractional stochastic Allen–Cahn equation on the one-dimensional flat torus. To further illustrate the significance of results of this form, we now discuss a few examples of semidiscrete SPDEs as well as their continuum limits.
Semidiscrete models appear frequently in the numerical analysis of (S)PDEs of evolution; since the challenges of spatial, temporal and spatiotemporal discretization are different, these settings are often analyzed separately. This leads to thorough studies of (S)PDEs that are discretized in space but not in time. In the context of SPDEs of evolution, we refer to [8, 40, 54] for examples.
Stochastic PDEs on graphs also appear naturally as models in the physical sciences, e.g., for interacting particle systems [17] or the representation of disordered media [44]. In the former case, the continuum limit represents the large particle limit in the interacting particle system.
In the data science literature, (S)PDEs on graphs have recently gained popularity as semi-supervised learning techniques. In a semi-supervised learning problem we are given a set of labeled features as well as a set of unlabeled features, and the goal is to use the former features to recover the labels of the latter. Features are, for example, images, text, or voice recordings; corresponding labels may be descriptors of the content of the images, the author of the text, or a transcript of the voice recording, respectively. Given an appropriate similarity measure on the space of features, an edge-weighted graph can be constructed in which nodes representing similar features are connected by highly weighted edges. The unknown labels can then be estimated by space-discretized PDEs on this graph, as in [6, 10, 79, 89]. The PDEs often describe gradient flows that minimize a variational functional. Stochastic PDEs appear in this setting if, in addition to finding an estimate for the labels, the uncertainty in the labels is to be quantified as well [7, 71]. The SPDEs of evolution here either form the basis of Markov chain Monte Carlo sampling algorithms [16, 42, 43, 69] or of a randomized global optimization scheme [14, 15] for the solution of the variational problem in the deterministic setting. In this semi-supervised learning setting, discrete-to-continuum limits are of interest because they establish the consistency of the models in the large-data limit. For deterministic PDEs, the literature has grown to encompass pointwise limits of operators, as in [45], $\Gamma$-limits of the functionals that underlie the dynamics [33, 58, 76, 78, 85], and more recently, discrete-to-continuum limits for the dynamics themselves [27, 30, 36, 47, 59, 63, 84]. For a more in-depth overview of the literature of discrete-to-continuum limits, we refer to [80].
1.2. Main results
We will now summarize the abstract setting and main discrete-to-continuum convergence results from Sections 46, which will already be applied in Section 3 to the Whittle–Matérn and stochastic fractional Allen–Cahn equations described in Section 1.1.
Given a probability space $(\Omega,\mathcal{F},\mathbb{P})$ and some $T\in(0,\infty)$ (called the time horizon), we consider a sequence of semilinear parabolic stochastic evolution equations
[…]
indexed by $n\in\overline{\mathbb{N}}\coloneqq\{1,2,\dots\}\cup\{\infty\}$. This problem will be rigorously formulated as a stochastic differential equation taking values in a real Banach space $E_{n}$ (or a smaller embedded space $B_{n}\hookrightarrow E_{n}$) called the state space. In general, we assume that the terms appearing in (1.1) are as follows:
  • $-A_{n}$ is the generator of a bounded analytic semigroup of bounded linear operators on $E_{n}$ or $B_{n}$;
  • $u_{n}\mapsto F_{n}(\omega,t,u_{n})$ is a possibly random and nonlinear drift operator on $E_{n}$ or $B_{n}$ for all $(\omega,t)\in\Omega\times[0,T]$;
  • $(W_{n}(t))_{t\geq 0}$ is the projection, onto an appropriate subspace, of an $H$-valued cylindrical Wiener process $(W(t))_{t\geq 0}$ for a real and separable Hilbert space $H$ (more details are specified in Theorem 1.1 below);
  • $\xi_{n}$ is a possibly random initial datum with values in $E_{n}$ or $B_{n}$.
The precise assumptions (in particular, whether the operators and initial data are $E_{n}$-valued or $B_{n}$-valued) vary throughout Sections 36; an overview is provided in Table 1. Depending on the setting, the mild solutions to (1.1) are either well defined on the whole of $[0,T]$ almost surely, or cease to exist at a time $t<T$ with nonzero probability; such solutions are said to be global or local, respectively.
The aim of this work is to establish conditions on the data of (1.1) under which the corresponding solutions $(X_{n})_{n\in\mathbb{N}}$ converge to $X_{\infty}$ as $n\to\infty$. In order to compare processes which take their values in different Banach spaces, we need to assume that each of the families $(E_{n})_{n\in\overline{\mathbb{N}}}$, $(H_{n})_{n\in\overline{\mathbb{N}}}$ and $(B_{n})_{n\in\overline{\mathbb{N}}}$ embeds uniformly into a common space—namely into $E_{\infty}$, $H_{\infty}$ and a closed subspace ${\widetilde{B}\subseteq B_{\infty}}$, respectively—which they approximate in some appropriate sense as $n\to\infty$. In particular, we shall assume that they share a common sequence of (linear) lifting operators $(\Lambda_{n})_{n\in\overline{\mathbb{N}}}$ such that each $\Lambda_{n}$ maps $E_{n}$ (resp. $H_{n}$, $B_{n}$) boundedly into $E_{\infty}$ (resp. $H_{\infty}$, $\widetilde{B}$), as well as a sequence of projection operators $(\Pi_{n})_{n\in\overline{\mathbb{N}}}$ which are left-inverses to the respective lifting operators. That is, each sequence satisfies Assumption 2.1 below with the same lifting and projection operators. As an example (see Section 3), one can think of $E_{\infty}\coloneqq L^{q}(\mathcal{D})$ for $q\in[2,\infty)$, $H_{\infty}\coloneqq L^{2}(\mathcal{D})$, $\widetilde{B}\coloneqq L^{\infty}(\mathcal{D})$ (Lebesgue spaces) and $B_{\infty}\coloneqq C(\mathcal{D})$ (continuous functions) for some spatial domain $\mathcal{D}$, along with $E_{n}\coloneqq L^{q}(\mathcal{D}_{n})$, $H_{n}\coloneqq L^{2}(\mathcal{D}_{n})$ and $B_{n}\coloneqq L^{\infty}(\mathcal{D}_{n})$ for some approximations $(\mathcal{D}_{n})_{n\in\mathbb{N}}$ of the domain $\mathcal{D}$.
The projection and lifting operators allow us to compare the ($E_{n}$- or $B_{n}$-valued) solution processes $X_{n}$ by instead considering convergence of the lifted processes $\widetilde{X}_{n}\coloneqq\Lambda_{n}X_{n}$ to $X_{\infty}$ as $n\to\infty$, which we call discrete-to-continuum convergence. Moreover, they allow us to formulate assumptions under which this occurs in terms of conditions imposed on the lifted resolvents $\widetilde{R}_{n}\coloneqq\Lambda_{n}(A_{n}+\operatorname{Id}_{n})^{-1}\Pi_{n}$ of the linear operators $A_{n}$, the lifted drift operators $\widetilde{F}_{n}(\omega,t,u)\coloneqq\Lambda_{n}F_{n}(t,\omega,\Pi_{n}u)$, and the lifted initial data $\widetilde{\xi}_{n}\coloneqq\Lambda_{n}\xi_{n}$. Roughly speaking, we assume that
  • $\widetilde{F}_{n}\to F_{\infty}$ ‘pointwise’ (see (F2) in Section 5.1 or (F2-B) in Section 6.1);
  • $\widetilde{R}_{n}\to R_{\infty}$ ‘pointwise’ and there exists a small enough $\beta\in[0,\frac{1}{2})$ such that the fractional powers $\widetilde{R}_{n}^{\beta}$ converge to $R_{\infty}^{\beta}$ in an appropriate operator norm (see (A3) in Section 4 or (A3-B) in Section 6.1)
  • $\widetilde{\xi}_{n}\to\xi_{\infty}$ in $L^{p}(\Omega;E_{\infty})$ or $L^{p}(\Omega;\widetilde{B})$ for some $p\in[1,\infty)$ (see (IC) in Section 5.1 or (IC-B) in Section 6.1).
Again we refer to Table 1 for an overview of the different settings and types of solutions, with references to the precise formulations of the corresponding assumptions; the setting in the first row (i.e., Section 3.2) covers the (fractional) stochastic Allen–Cahn equations announced in Section 1.1, see Example 3.8 below. The following theorem is a summary of the discrete-to-continuum approximation results for solutions to the abstract equations (1.1) in these respective settings.
Theorem 1.1 (Discrete-to-continuum convergence—summarized).
Let $(\Omega,\mathcal{F},\mathbb{P})$ be a probability space and $T\in(0,\infty)$ be a terminal time. Consider equations (1.1), where the state spaces, linear operators, drift operators and initial data are as in one of the rows of Table 1. Let $p\in[1,\infty)$ be the stochastic integrability of the initial data and let $W_{n}\coloneqq\Pi_{n}W$, where $(W(t))_{t\geq 0}$ is an $H$-valued cylindrical Wiener process. For all $n\in\overline{\mathbb{N}}$, there exists a unique (local or global, see Table 1) mild solution $X_{n}$ to (1.1), and the lifted solution processes $\widetilde{X}_{n}\coloneqq\Lambda_{n}X_{n}$ satisfy the following:
  1. (i)
    If the solutions are global and $p>1$, then for all $p^{-}\in[1,p)$ we have
    $\widetilde{X}_{n}\to X_{\infty}\quad\text{as }n\to\infty$ (1.2)
    in $L^{p^{-}}\!(\Omega;C([0,T];E_{\infty}))$ (resp. in $L^{p^{-}}\!(\Omega;C([0,T];\widetilde{B}))$). In the (semi)linear settings with globally Lipschitz drifts of linear growth, the same in fact holds with $p^{-}\coloneqq p$ for any $p\in[1,\infty)$.
  2. (ii)
    If the solutions are local, with associated explosion times $\sigma_{n}\colon\Omega\to(0,T]$ (precisely defined in (5.41) below), then we have
    $\widetilde{X}_{n}\mathbf{1}_{[0,\sigma_{\infty}\wedge\sigma_{n})}\to X_{\infty}\mathbf{1}_{[0,\sigma_{\infty})}\quad\text{as }n\to\infty$ (1.3)
    in $L^{0}(\Omega\times[0,T];E_{\infty})$ (resp. in $L^{0}(\Omega\times[0,T];\widetilde{B})$), where $L^{0}$ indicates convergence in measure.
The full convergence statement for each setting is given in the corresponding part of Sections 46. To be precise, Theorem 1.1 is comprised of the following results, in order of appearance: Theorem 3.10, Proposition 4.5, Theorems 5.4 and 5.9, Proposition 6.1, Theorems 6.6 and 6.7 and Corollary 6.10.
Sec. Description Assumptions Sol. type
§3.2 graph-based approximation of Whittle–Matérn operators on a manifold $A_{n}\coloneqq[\mathcal{L}_{n}^{\tau,\kappa}]^{s}$ (Whittle–Matérn operators) $[F_{n}(t,u)](x)\coloneqq f_{n}(t,u(x))$ (Nemytskii drift) Assumption 3.7 (on the functions $(f_{n})_{n\in\overline{\mathbb{N}}}$) global
§4 $E_{n}$-valued linear (A1)(A3) (linear operators) $F_{n}\coloneqq 0$ and $\xi_{n}\coloneqq 0$ global
§5.1 $E_{n}$-valued semilinear;
globally Lipschitz drifts of linear growth
(A1)(A3) (linear operators) (F1)(F2) (drift operators) (IC) (initial data) global
§5.2 $E_{n}$-valued semilinear;
locally Lipschitz and locally bounded drifts
(A1)(A3) (linear operators) (F1) and (F2) (drift operators) (IC) (initial data) local
§6.1 $B_{n}$-valued semilinear;
globally Lipschitz drifts of linear growth
(A1-B)(A4-B) with $\theta+2\beta<1$ (lin. ops.) (F1-B)(F2-B) (drift operators) (IC-B) (initial data) global
§6.2 $B_{n}$-valued semilinear;
locally Lipschitz and locally bounded drifts
(A1-B)(A4-B) with $\theta+2\beta<1$ (lin. ops.) (F1-B) and (F2-B) (drift operators) (IC-B) (initial data) local
§6.3 $B_{n}$-valued semilinear;
dissipative drifts
(A1-B)(A4-B) with $\theta+2\beta<1$ (lin. ops.) (F1′′-B) and (F2-B) (drift operators) (IC-B) (initial data) global
Table 1. Overview of the types of SPDEs considered in the different (sub)sections comprising this work. The row in which an assumption appears for the first time also indicates the (sub)section where its definition can be found.
1.3. Contributions
The abstract discrete-to-continuum approximation theorems for stochastic semilinear parabolic evolution equations driven by additive cylindrical Wiener noise—summarized in Theorem 1.1 and proved in Sections 46—complement the results from [55, 56], which establish the continuous dependence on the coefficients of semilinear equations driven by multiplicative noise in a state space with unconditional martingale differences (UMD), i.e., convergence in the case where $E_{n}=E$ (and $B_{n}=B$) for all $n\in\overline{\mathbb{N}}$. Given the motivating applications and our aim to provide a self-contained exposition of the proofs, we make the simplifying assumptions that the UMD spaces $(E_{n})_{n\in\overline{\mathbb{N}}}$ have Rademacher type 2 (which was also assumed for $E$ in [56] but not in [55]) and the noise is additive.
Under these conditions, we provide a direct proof of convergence of the $E_{n}$-valued stochastic convolutions solving the linear parts of (1.1) using a Da Prato–Kwapień–Zabczyk factorization argument [19] and a discrete-to-continuum analog of the Trotter–Kato approximation theorem [51, Theorem 2.1], see Proposition 4.5. We extend it to the semilinear $E_{n}$-valued settings described in Table 1 by adapting the arguments from [55, Sections 3 and 4] and [56, Subsection 3.1] to incorporate the discrete-to-continuum projection and lifting operators, yielding Theorems 5.4 and 5.9, respectively. In order to state and prove the analogous Theorems 6.6 and 6.7 for the $B_{n}$-valued settings, we impose a uniform ultracontractivity condition on the semigroups which replaces the restriction in [56, Section 3] that the fractional domain spaces $\dot{E}_{n}^{\alpha}\coloneqq\mathsf{D}((\operatorname{Id}_{n}+A_{n})^{\alpha/2})$ also coincide for all $n\in\overline{\mathbb{N}}$.
Theorem 3.10, regarding the graph discretization of equations whose linear operators are of generalized Whittle–Matérn type on a manifold $\mathcal{M}$, extends analogous convergence results for linear equations on a spatial domain (cf. [71, Theorem 4.2] and [71, Theorem 7] in $L^{2}(\mathcal{M})$ and $L^{\infty}(\mathcal{M})$, respectively) to spatiotemporal and semilinear equations. Like the cited theorems, its proof relies on recent spectral convergence results for graph Laplacians (see [11] and [12] for convergence of eigenfunctions in $L^{2}$ and $L^{\infty}$, respectively), which we use to verify that these SPDEs fit into the abstract framework from Sections 46.
1.4. Outline
The remainder of this article is structured as follows. In Section 2, we establish some notational conventions and collect preliminaries regarding the (deterministic) discrete-to-continuum Trotter–Kato approximation theorem and stochastic integration in UMD-type-2 Banach spaces. We demonstrate in Section 3 how the results summarized by Theorem 1.1 can be applied to graph discretizations of stochastic parabolic evolution equations whose linear part is a generalized Whittle–Matérn operator on a manifold. In Section 4, we consider the linear $E_{n}$-valued version of (1.1), whose solutions are also known as infinite-dimensional Ornstein–Uhlenbeck processes. These results are extended in Section 5 to allow for semilinear $E_{n}$-valued drift operators under (local or global) Lipschitz continuity and boundedness assumptions. In Section 6 we first treat the analogous results in the semilinear $B_{n}$-valued setting, and then establish global well-posedness and convergence for dissipative drifts. Finally, in Section 7 we discuss some potential directions for further research. This work is complemented by three appendices: Appendix A consists of postponed proofs of some intermediate results from Section 3. Appendices B and C are concerned with fractional parabolic integration and (uniformly) sectorial linear operators, respectively. See Figure 1 for a schematic overview of the relations between Sections 36 and the appendices.
Section 4 Appendix B Appendix C Subsection 5.1 Subsection 5.2 Subsection 6.1 Subsection 6.2 Subsection 6.3 Section 3 Appendix A
Figure 1. Relations between the sections comprising the main part of this article and the appendix. Arrows indicate when the results of one section are applied in another.
2. Preliminaries
2.1. Notation
Elementary sets and operations
$\mathbb{N}$ positive integers
$\mathbb{N}_{0}$ nonnegative integers
$\overline{\mathbb{N}}$$\mathbb{N}\cup\{\infty\}$
$\operatorname{Id}_{D}$ identity map on a set $D$
$\mathbf{1}_{D_{0}}$ indicator map on $D_{0}\subseteq D$
$s\wedge t$ minimum of $s,t\in\mathbb{R}$
Bounded linear operators
$H$, $K$ separable Hilbert spaces
$E$, $F$ arbitrary Banach spaces
$\langle\,\cdot\,,\,\cdot\,\rangle_{H}$ inner product of $H$
$\|\,\cdot\,\|_{E}$ norm of $E$
$E^{*}$ dual space of $E$
$\mathscr{L}(E;F)$ bounded linear operators (from $E$ to $F$)
$\mathscr{L}(E)$ abbreviation of $\mathscr{L}(E;E)$
$\gamma(H;E)$ $\gamma$-radonifying operators
$\mathscr{L}_{2}(H;K)$ Hilbert–Schmidt operators
Function spaces
$C(K;E)$ continuous functions from a compact space $K$ to $E$
$C(K)$ abbreviation of $C(K;\mathbb{R})$
$L^{p}(S;E)$ Bochner space of $p$-integrable functions from a measure space $(S,\mathscr{A}\!,\nu)$ to $E$
$L^{p}(S)$ Lebesgue space $L^{p}(S;\mathbb{R})$
Graph discretization
$\mathcal{M}$ manifold from Assumption 3.1
$\mathbb{T}^{m}$ $m$-dimensional flat torus
$d_{\mathcal{M}}$ geodesic metric on $\mathcal{M}$
$\mu$ volume measure on $\mathcal{M}$
$\mathcal{M}_{n}$ point cloud $(x_{n}^{(j)})_{j=1}^{n}\subset\mathcal{M}$
$(\widetilde{\Omega},\widetilde{\mathcal{F}},\widetilde{\mathbb{P}})$ probability space of random point cloud from Example 3.3
$\mu_{n}$ empirical measure on $\mathcal{M}_{n}$
$T_{n}$ transport map from $\mathcal{M}$ to $\mathcal{M}_{n}$
$\varepsilon_{n}$ $\sup_{x\in\mathcal{M}}d_{\mathcal{M}}(x,T_{n}(x))$
$h_{n}$ graph connectivity length scale, see (3.10)
$\mathcal{L}_{n}^{\tau,\kappa}$ (discretized) Whittle–Matérn operator with coefficient functions $\tau,\kappa\colon\mathcal{M}\to[0,\infty)$
$(\lambda_{n}^{(j)})_{j=1}^{n}$ eigenvalues of $\mathcal{L}_{n}^{\tau,\kappa}$
$(\psi_{n}^{(j)})_{j=1}^{n}$ $L^{2}(\mathcal{M}_{n})$-normalized eigenfunctions of $\mathcal{L}_{n}^{\tau,\kappa}$
$M_{\psi,\infty}$ uniform $L^{\infty}$-bound of the eigenfunctions, see Assumption 3.9
$M_{S,q}$ uniform-ultracontractivity constant, see (3.36)
‘Discrete-to-continuum’ spaces
$(E_{n})_{n\in\overline{\mathbb{N}}}$, $\widetilde{E}$ Banach spaces from Assumption 2.1 or (A1)
$(H_{n})_{n\in\overline{\mathbb{N}}}$, $\widetilde{H}$ Hilbert spaces from (A1)
$(B_{n})_{n\in\overline{\mathbb{N}}}$, $\widetilde{B}$ Banach spaces from (A1-B)
Projection and lifting
$\Pi_{n}$ projection operator from $E_{n}$ (resp. $H_{n}$ or $B_{n}$) to $\widetilde{E}$ (resp. $\widetilde{H}$ or $\widetilde{B}$)
$\Lambda_{n}$ lifting operator from $\widetilde{E}$ (resp. $\widetilde{H}$ or $\widetilde{B}$) to $E_{n}$ (resp. $H_{n}$ or $B_{n}$)
$\widetilde{T}_{n}$ lifted version $\Lambda_{n}T_{n}\Pi_{n}$ on $\widetilde{E}$ (resp. $\widetilde{H}$ or $\widetilde{B}$) of operator $T_{n}$ on $E_{n}$ (resp. $H_{n}$ or $B_{n}$)
$\widetilde{Y}_{n}$ lifted version $\Lambda_{n}Y_{n}$ on $\widetilde{E}$ (resp. $\widetilde{H}$ or $\widetilde{B}$) of process $Y_{n}$ on $E_{n}$ (resp. $H_{n}$ or $B_{n}$)
$M_{\Pi}$ $\sup_{n\in\mathbb{N}}\|\Pi_{n}\|_{\mathscr{L}(\widetilde{E};E_{n})}$
$\widetilde{M}_{\Pi}$ $\sup_{n\in\mathbb{N}}\|\Pi_{n}\|_{\mathscr{L}(\widetilde{B};B_{n})}$
$M_{\Lambda}$ $\sup_{n\in\mathbb{N}}\|\Lambda_{n}\|_{\mathscr{L}(E_{n};\widetilde{E})}$
$\widetilde{M}_{\Lambda}$ $\sup_{n\in\mathbb{N}}\|\Lambda_{n}\|_{\mathscr{L}(B_{n};\widetilde{B})}$
Linear operators in evolution equations
$A_{n}$ linear operator on $E_{n}$ with domain $\mathsf{D}(A_{n})$
$S_{n}$ semigroup generated by $-A_{n}$
$M_{S}$ uniform-boundedness constant of $(S_{n})_{n\in\overline{\mathbb{N}}}$ in $(E_{n})_{n\in\overline{\mathbb{N}}}$, see (2.1)
$\widetilde{M}_{S}$ uniform-boundedness constant of $(S_{n})_{n\in\overline{\mathbb{N}}}$ in $(B_{n})_{n\in\overline{\mathbb{N}}}$, see (6.4)
$\rho(A_{n})$ resolvent set of $A_{n}$
$R_{n}^{\beta}$ $(A_{n}+\operatorname{Id}_{n})^{-\beta}$
$\mathfrak{I}_{A_{n}}^{s}$ fractional parabolic integration operator, see Appendix B
Stochastic evolution equations
$Q_{n}$ covariance $\Pi_{n}\Pi_{n}^{*}\in\mathscr{L}(H_{n})$
$\mathop{}\!\mathrm{d}W_{n}$ $H_{n}$-valued $Q_{n}$-cylindrical Wiener noise on $(\Omega,\mathcal{F},\mathbb{P})$
$W_{A_{n}}$ stochastic convolution, see (4.2)
$\xi_{n}$ initial datum
$F_{n}$ drift operator
$L_{F}$, $C_{F}$ Lipschitz and growth constants of $E_{n}$-valued drifts, see (F1)
$\widetilde{L}_{F}$, $\widetilde{C}_{F}$ Lipschitz and growth constants of $B_{n}$-valued drifts, see (F1-B)
$L_{F}^{(r)}$, $C_{F,0}$ local Lipschitz and growth constants of $E_{n}$-valued drifts, see (F1)
$f_{n}$ real-valued drift coefficient function
$\widetilde{L}_{f}$, $\widetilde{C}_{f}$ Lipschitz and growth constants of $(f_{n})_{n\in\overline{\mathbb{N}}}$ see Assumption 3.7(i)
Table 2. A selection of notation which is used repeatedly throughout this work.
Table 2 lists some of the most important notation which recurs throughout this work.
The Cartesian product $\prod_{j\in\mathcal{I}}B_{j}$ of an indexed family of sets $(B_{j})_{j\in\mathcal{I}}$ is comprised of all functions $f\colon\mathcal{I}\to\bigcup_{j\in\mathcal{I}}B_{j}$ satisfying $f(j)\in B_{j}$ for all $j\in\mathcal{I}$. Given two sets $\mathscr{P},\mathscr{Q}$ and maps $\mathscr{F},\mathscr{G}\colon\mathscr{P}\times\mathscr{Q}\to\mathbb{R}$, we write $\mathscr{F}(p,q)\lesssim_{q}\mathscr{G}(p,q)$ to indicate the existence of some $C\colon\mathscr{Q}\to[0,\infty)$ such that $\mathscr{F}(p,q)\leq C(q)\,\mathscr{G}(p,q)$ for all $(p,q)\in\mathscr{P}\times\mathscr{Q}$. We write $\mathscr{F}(p,q)\eqsim_{q}\mathscr{G}(p,q)$ if both $\mathscr{F}(p,q)\lesssim_{q}\mathscr{G}(p,q)$ and $\mathscr{G}(p,q)\lesssim_{q}\mathscr{F}(p,q)$ hold.
All normed spaces will be considered over the real or complex scalar field. Results concerning spectra and complex interpolation are naturally formulated for complex spaces; if the space is real, then we implicitly apply them to the complexification and subsequently restrict back to the original space.
The space $\mathscr{L}(E;F)$ of bounded linear operators from $E$ to $F$ shall be equipped with the norm $\|T\|_{\mathscr{L}(E;F)}\coloneqq\sup_{\|x\|_{E}=1}\|Tx\|_{F}$. We call $T\in\mathscr{L}(E;F)$ a contraction if $\|T\|_{\mathscr{L}(E;F)}\leq 1$; in particular, the inequality need not be strict. An operator $T\in\mathscr{L}(H)$ is said to be positive definite if there exists $\theta\in(0,\infty)$ such that $\langle Tx,x\rangle_{H}\geq\theta\|x\|_{H}^{2}$ for all $x\in H$, and nonnegative definite if $\langle Tx,x\rangle_{H}\geq 0$. Given $T,S\in\mathscr{L}(H;K)$, we set $\langle T,S\rangle_{\mathscr{L}_{2}(H;K)}\coloneqq\sum_{j=1}^{\infty}\langle Te_{j},Se_{j}\rangle_{K}$, where $(e_{j})_{j\in\mathbb{N}}$ is any orthonormal basis of $H$; the space $\mathscr{L}_{2}(H;K)$ of Hilbert–Schmidt operators consists of those $T$ for which $\|T\|_{\mathscr{L}_{2}(H;K)}^{2}\coloneqq\langle T,T\rangle_{\mathscr{L}_{2}(H;K)}<\infty$.
Given a measure space $(S,\mathscr{A}\!,\nu)$, a function $f\colon S\to E$ is said to be strongly measurable if it can be approximated $\nu$-a.e. by simple functions. For $p\in[1,\infty]$, let $L^{p}(S;E)\coloneqq L^{p}(S,\mathscr{A}\!,\nu;E)$ denote the Bochner space of (equivalence classes of) strongly measurable and $p$-integrable functions from $S$ to $E$, with norm
$\|f\|_{L^{p}(S;E)}\coloneqq\begin{cases}(\int_{S}\|f(s)\|_{E}^{p}\mathop{}\!\mathrm{d}\nu(s))^{\frac{1}{p}},&\text{if }p\in[1,\infty);\\ \operatorname{ess\,sup}_{s\in S}\|f(s)\|_{E},&\text{if }p=\infty.\end{cases}$
Sub-intervals $J\subseteq\mathbb{R}$ are equipped with the Lebesgue $\sigma$-algebra and measure. The Banach space of (bounded) continuous functions $u\colon J\to E$, endowed with the supremum norm, is denoted by $C(J;E)$.
The meaning of the tensor symbol $\otimes$ will depend on the context: Given a map $\Phi\colon(a,b)\to\mathscr{L}(E;F)$ and some $x\in E$, we define the function $\Phi\otimes x\colon(a,b)\to F$ by $[\Phi\otimes x](t)\coloneqq\Phi(t)x$. If instead an $h\in H$ is given, we define $h\otimes x\in\mathscr{L}(H;E)$ to be the rank-one operator $[h\otimes x](u)\coloneqq\langle h,u\rangle_{H}x$. The space of all (finite) linear combinations of such operators is denoted by $H\otimes E$. We define the convolution $\Psi*f\colon[0,T]\to F$ of the functions $\Psi\colon[0,T]\to\mathscr{L}(E;F)$ and $f\colon[0,T]\to E$ by $[\Psi*f](t)\coloneqq\int_{0}^{t}\Psi(t-s)f(s)\mathop{}\!\mathrm{d}s$.
2.2. Discrete-to-continuum Trotter–Kato approximation theorem
We encode the discrete-to-continuum setting in the following way:
Assumption 2.1.
Let $(E_{n},\|\,\cdot\,\|_{E_{n}})_{n\in\overline{\mathbb{N}}}$ and $(\widetilde{E},\|\,\cdot\,\|_{\widetilde{E}})$ be real or complex Banach spaces and suppose that $E_{\infty}$ is a closed linear subspace of $\widetilde{E}$. We assume that there exist operators $\Pi_{n}\in\mathscr{L}(\widetilde{E};E_{n})$ and $\Lambda_{n}\in\mathscr{L}(E_{n};\widetilde{E})$ for all $n\in\mathbb{N}$ which satisfy
  1. (i)
    $M_{\Pi}\coloneqq\sup_{n\in\mathbb{N}}\|\Pi_{n}\|_{\mathscr{L}(\widetilde{E};E_{n})}<\infty$ and $M_{\Lambda}\coloneqq\sup_{n\in\mathbb{N}}\|\Lambda_{n}\|_{\mathscr{L}(E_{n};\widetilde{E})}<\infty$;
  2. (ii)
    $\Lambda_{n}\Pi_{n}x\to x$ in $\widetilde{E}$ as $n\to\infty$ for all $x\in E_{\infty}$;
  3. (iii)
    $\Pi_{n}\Lambda_{n}=\operatorname{Id}_{E_{n}}$ for all $n\in\mathbb{N}$.
In addition, we denote $\Pi_{\infty}=\Lambda_{\infty}\coloneqq\operatorname{Id}_{\widetilde{E}}$ for convenience.
Note that parts (i) and (iii) together imply that the lifting operators are continuous embeddings $\Lambda_{n}\colon E_{n}\hookrightarrow\widetilde{E}$. In applications, they will typically be nested (in the sense that $E_{n}\hookrightarrow E_{n+1}$ for all $n\in\mathbb{N}$) and finite-dimensional, but neither of these assumptions is strictly necessary in the abstract theory. Moreover, we will often have $\widetilde{E}=E_{\infty}$, but not always; see Section 6.
Now we consider the following sequence of linear operators on $(E_{n})_{n\in\overline{\mathbb{N}}}$:
Assumption 2.2.
For all $n\in\overline{\mathbb{N}}$, let $-A_{n}\colon\mathsf{D}(A_{n})\subseteq E_{n}\to E_{n}$ be a linear operator generating a strongly continuous semigroup $(S_{n}(t))_{t\geq 0}\subseteq\mathscr{L}(E_{n})$, meaning that $S_{n}(0)=\operatorname{Id}_{E_{n}}$, $S_{n}(t+s)=S_{n}(t)S_{n}(s)$ for all $t,s\geq 0$, and $S_{n}(t)x\to x$ as $t\downarrow 0$ for all $x\in E_{n}$. Suppose that there exist $M_{S}\in[1,\infty)$ and $w\in\mathbb{R}$ such that
$\|S_{n}(t)\|_{\mathscr{L}(E_{n})}\leq M_{S}e^{-wt}\quad\text{for all }n\in\overline{\mathbb{N}}\text{ and }t\in[0,\infty).$ (2.1)
Given a linear operator $A\colon\mathsf{D}(A)\subseteq E\to E$ on a Banach space $(E,\|\,\cdot\,\|_{E})$, where $\mathsf{D}(A)$ denotes its domain, we say that $\lambda\in\mathbb{C}$ belongs to the resolvent set $\rho(A)$ of $A$ if the corresponding resolvent operator $R(\lambda,A):=(\lambda\operatorname{Id}_{E}-A)^{-1}$ exists in $\mathscr{L}(E)$. Given a sequence $(A_{n})_{n\in\overline{\mathbb{N}}}$ of such operators, generating strongly continuous semigroups with uniform growth bounds, the Trotter–Kato approximation theorem (see, e.g., [29, Chapter III, Theorem 4.8]) establishes a link between the strong convergence of resolvents and uniform convergence of the semigroups on compact subintervals of $[0,\infty)$. The following discrete-to-continuum analog of this result was proved by Ito and Kappel [51, Theorem 2.1]:
Theorem 2.3 (Discrete-to-continuum Trotter–Kato approximation).
Let Assumptions 2.1 and 2.2 be satisfied, with $w\in\mathbb{R}$. The following statements are equivalent:
  1. (a)
    There exists a $\lambda\in\bigcap_{n\in\overline{\mathbb{N}}}\rho(A_{n})$ such that, for every $x\in E_{\infty}$,
    $\Lambda_{n}R(\lambda,A_{n})\Pi_{n}x\to R(\lambda,A_{\infty})x\quad\text{in }\widetilde{E}\quad\text{as }n\to\infty.$ (2.2)
  2. (b)
    For all $x\in E_{\infty}$ and $T\in(0,\infty)$ it holds that
    $\Lambda_{n}S_{n}\Pi_{n}\otimes x\to S_{\infty}\otimes x\quad\text{in }C([0,T];\widetilde{E})\quad\text{as }n\to\infty.$ (2.3)
If a holds for some $\lambda\in\bigcap_{n\in\overline{\mathbb{N}}}\rho(A_{n})$ (or, equivalently, if b holds), then a holds in fact for every $\lambda\in\mathbb{C}$ such that $\operatorname{Re}\lambda<w$.
2.3. Stochastic integration in UMD-type-2 Banach spaces
Given a separable Hilbert space $(H,\langle\,\cdot\,,\,\cdot\,\rangle_{H})$ over the real scalar field, let $(W(t))_{t\geq 0}$ be an $H$-valued cylindrical Wiener process with respect to some filtered probability space $(\Omega,(\mathcal{F}_{t})_{t\geq 0},\mathcal{F},\mathbb{P})$. A rigorous definition can be found in [60, Section 2.5.1]. For our purpose of constructing the stochastic integral below, it suffices to define $\langle W(t),h\rangle_{H}\coloneqq\sum_{j=1}^{\infty}\beta_{j}(t)\langle e_{j},h\rangle_{H}$ for all $t\geq 0$ and $h\in H$, where $(e_{j})_{j\in\mathbb{N}}$ is an orthonormal basis of $H$ and $(\beta_{j}(\,\cdot\,))_{j\in\mathbb{N}}$ is a sequence of independent (real-valued) Brownian motions on $(\Omega,(\mathcal{F}_{t})_{t\geq 0},\mathcal{F},\mathbb{P})$. Then $(\langle W(t),h\rangle_{H})_{t\geq 0}$ is a well-defined Brownian motion for every $h\in H$, and intuitively one can think of $(W(t))_{t\geq 0}$ as being given by $W(t)=\sum_{j=1}^{\infty}\beta_{j}(t)e_{j}$.
Let $(E,\|\,\cdot\,\|_{E})$ be a real Banach space, and let $(\gamma_{j})_{j\in\mathbb{N}}$ be a sequence of independent (real-valued) standard normal random variables on a probability space $(\Omega^{\prime},\mathcal{F}^{\prime},\mathbb{P}^{\prime})$, independent of the probability spaces $(\Omega,\mathcal{F},\mathbb{P})$ and $(\widetilde{\Omega},\widetilde{\mathcal{F}},\widetilde{\mathbb{P}})$ used in the rest of this work. We define the space $\gamma(H;E)$ of $\gamma$-radonifying operators from $H$ to $E$ as the completion of the finite-rank operators $H\otimes E$ with respect to the norm $\lVert\sum_{j=1}^{n}h_{j}\otimes x_{j}\rVert_{\gamma(H;E)}\coloneqq\lVert\sum_{j=1}^{n}\gamma_{j}x_{j}\rVert_{L^{2}(\Omega^{\prime};E)},$ where we assume that the $(h_{j})_{j=1}^{n}$ are $H$-orthonormal. This norm is well-defined, i.e., it can be checked that the right-hand side is independent of the choice of representation. An important feature of $\gamma(H;E)$ is its ideal property (in the algebraic sense) [49, Theorem 9.1.10], which states that for all $T\in\gamma(H;E)$, $U\in\mathscr{L}(E;F)$ and $S\in\mathscr{L}(K;H)$, we have
$UTS\in\gamma(K;F)\quad\text{with}\quad\|UTS\|_{\gamma(K;F)}\leq\|U\|_{\mathscr{L}(E;F)}\|T\|_{\gamma(H;E)}\|S\|_{\mathscr{L}(K;H)}.$ (2.4)
For any rank-one operator $h\otimes x\in\mathscr{L}(H;E)$, we have $h\otimes x\in\gamma(H;E)$ with
$\|h\otimes x\|_{\gamma(H;E)}=\|h\|_{H}\|x\|_{E}.$ (2.5)
The stochastic integral of an elementary integrand $\Phi\colon(0,\infty)\to H\otimes E$, i.e., a function of the form $\Phi(t)=\sum_{j=1}^{n}\mathbf{1}_{(a_{j},b_{j}]}(t)h_{j}\otimes x_{j}$, is defined by
$\int_{0}^{\infty}\Phi(t)\mathop{}\!\mathrm{d}W(t)\coloneqq\sum_{j=1}^{n}\bigl{(}\langle h_{j},W(b_{j})\rangle_{H}-\langle h_{j},W(a_{j})\rangle_{H}\bigr{)}x_{j}\in L^{2}(\Omega;E).$ (2.6)
In order to extend the definition of the stochastic integral beyond elementary integrands, one needs to impose further geometric assumptions on the Banach space $E$. In this article we work in one of the standard settings, namely that of spaces with unconditional martingale differences and Rademacher type $2$ (abbreviated to UMD-type-2). Definitions of these notions can be found in [48, Section 4.2] and [49, Section 7.1], respectively, but we will only use them to ensure existence of stochastic integrals. In this case, one can establish the Itô inequality
$\bigg{\lVert}\int_{0}^{\infty}\Phi(t)\mathop{}\!\mathrm{d}W(t)\bigg{\rVert}_{L^{2}(\Omega;E)}\lesssim_{E}\|\Phi\|_{L^{2}(0,\infty;\gamma(H;E))}$ (2.7)
for elementary integrands [82, Proposition 4.2], and use it to extend the definition of the stochastic integral to all $\Phi\in L^{2}(0,\infty;\gamma(H;E))$. In fact, since we are only concerned with deterministic integrands in this work, we could suffice with the type $2$ assumption. Despite this, we additionally impose the UMD assumption for the sake of compatibility with some of the literature, and because the concrete examples of Banach spaces in which we are interested (such as the Lebesgue $L^{q}$-spaces for $q\in[2,\infty)$) satisfy both properties. For more details on stochastic integration in Banach spaces, we refer to the survey article [82].
The exponent $2$ in $L^{2}$ appearing on both sides of (2.7) can be replaced by any other $p\in[1,\infty)$ at the cost of a $p$-dependent constant, see for instance [82, Theorem 4.7]. If $E$ is also a Hilbert space, then $\gamma(H;E)$ is isometrically isomorphic to the space $\mathscr{L}_{2}(H;E)$, see [49, Proposition 9.1.9], and instead of the inequality (2.7) we have the Itô isometry between $L^{2}(0,\infty;\mathscr{L}_{2}(H;E))$ and $L^{2}(\Omega;E)$.
3. Graph-based semilinear stochastic evolution equations with Whittle–Matérn linear operators
Before developing the general discrete-to-continuum convergence results summarized by Theorem 1.1 in the upcoming sections, in this section we demonstrate how they can be applied to the particular case of equations whose linear parts are graph discretizations of a generalized Whittle–Matérn operator on a manifold. In the spatial and linear case, such convergence results have been proven in [71, 72]. We also mention the work [65], in which the statistical properties of the spatiotemporal linear equation were investigated for fixed $n\in\mathbb{N}$.
3.1. Geometric graphs and generalized Whittle–Matérn operators
Assumption 3.1 (Manifold assumption).
Given $m,d\in\mathbb{N}$, suppose that $\mathcal{M}$ is an $m$-dimensional smooth, connected, compact Riemannian manifold without boundary, embedded smoothly and isometrically into $\mathbb{R}^{d}$. Let $\mu$ and $d_{\mathcal{M}}$ denote the normalized volume measure and geodesic metric on $\mathcal{M}$, respectively.
For each $n\in\mathbb{N}$, let a point cloud $\mathcal{M}_{n}\coloneqq(x_{n}^{(j)})_{j=1}^{n}\subseteq\mathcal{M}$ be given. We suppose that $\mathcal{M}$ can be partitioned into $n$ regions of mass $1/n$, which can be transported to the corresponding $n$ points comprising $\mathcal{M}_{n}$, in such a way that the maximal geodetic displacement tends to zero as $n\to\infty$. More precisely, we assume that there exists a sequence $(T_{n})_{n\in\mathbb{N}}$ of transport maps $T_{n}\colon\mathcal{M}\to\mathcal{M}_{n}$ such that
$\mu_{n}={}$ $T_{n\sharp}\mu$ for all $\text{for all }n$ $\in\mathbb{N},\quad\text{and}$ (3.1)
[…]
Here, $\mu_{n}\coloneqq\frac{1}{n}\sum_{j=1}^{n}\delta_{x_{n}^{(j)}}$ is the empirical measure on $\mathcal{M}$ associated to $\mathcal{M}_{n}$, and $T_{n\sharp}\mu$ denotes the pushforward measure $T_{n\sharp}\mu(B)\coloneqq\mu(\{T_{n}\in B\})$ on $\mathcal{M}_{n}$. Two different examples in which this assumption is satisfied are presented in Settings 3.2 and 3.3 below.
Given $u_{n}\colon\mathcal{M}_{n}\to\mathbb{R}$ for $n\in\overline{\mathbb{N}}$, these transport maps enable us to define the functions $\Lambda_{u}u_{n}\colon\mathcal{M}\to\mathbb{R}$ and $\Pi_{n}u\colon\mathcal{M}_{n}\to\mathbb{R}$ by setting
$\Lambda_{n}u_{n}(x)\coloneqq u_{n}(T_{n}(x))\quad\text{and}\quad\Pi_{n}u(x_{n}^{(j)})\coloneqq n\int_{V_{n}^{(j)}}u(x)\mathop{}\!\mathrm{d}\mu(x),$ (3.3)
respectively, for all $x\in\mathcal{M}$ and $j\in\{1,\dots,n\}$, where $V_{n}^{(j)}\coloneqq\{T_{n}=x_{n}^{(j)}\}\subseteq\mathcal{M}$.
It turns out that the operations defined in (3.3) satisfy Assumption 2.1 with respect to the following function spaces: Given $q\in[1,\infty]$ and $n\in\mathbb{N}$, we set $E_{n}\coloneqq L^{q}(\mathcal{M}_{n})\coloneqq L^{q}(\mathcal{M},\mu_{n})$, as well as $\widetilde{E}\coloneqq L^{q}(\mathcal{M})$ and
$E_{\infty}\coloneqq\begin{cases}L^{q}(\mathcal{M}),&\text{if }q\in[1,\infty);\\ C(\mathcal{M}),&\text{if }q=\infty.\end{cases}$ (3.4)
Later on, we will need $q\in[2,\infty)$, so that $E_{n}$ is a UMD-type-2 space for use in stochastic integration, but the statements here hold for all $q\in[1,\infty]$.
For these spaces, Assumption 2.1(i) is satisfied with $M_{\Lambda}=1$ and $M_{\Pi}\leq 1$. Indeed, the fact that $\Lambda_{n}$ is an isometry follows from (3.1) if $q\in[1,\infty)$, whereas for $q=\infty$ we see directly from the definition that
$\|\Lambda_{n}u_{n}\|_{L^{\infty}(\mathcal{M})}=\sup_{x\in\mathcal{M}}\lvert u_{n}(T_{n}(x))\rvert=\max_{j=1}^{n}\lvert u_{n}(x_{n}^{(j)})\rvert=\|u_{n}\|_{L^{\infty}(\mathcal{M}_{n})}.$
To show that $\Pi_{n}$ is a contraction, we first apply Hölder’s inequality in (3.3) with $\frac{1}{q}+\frac{1}{q^{\prime}}=1$ to find $\lvert\Pi_{n}u(x_{n}^{(j)})\rvert\leq n\|u\|_{L^{q}(V_{n}^{(j)})}\mu(V_{n}^{(j)})^{\frac{1}{q^{\prime}}}=n^{\frac{1}{q}}\|u\|_{L^{q}(V_{n}^{(j)})}$, so that
$\|\Pi_{n}u\|_{L^{q}(\mathcal{M}_{n})}^{q}=\frac{1}{n}\sum_{j=1}^{n}\lvert\Pi_{n}u(x_{n}^{(j)})\rvert^{q}\leq\sum_{j=1}^{n}\|u\|_{L^{q}(V_{n}^{(j)}\!\!,\,\mu)}^{q}=\|u\|_{L^{q}(\mathcal{M})}^{q}.$ (3.5)
Assumption 2.1(ii) is a consequence of (3.2), from which it follows that $\Lambda_{n}\Pi_{n}u\to u$ in $L^{q}(\mathcal{M})$ for any $u\in C(\mathcal{M})$. For $q=\infty$, this is what we wanted to show; if $q\in[1,\infty)$, then $C(\mathcal{M})$ is dense in $L^{q}(\mathcal{M})$, and the fact that $(\Lambda_{n}\Pi_{n})_{n\in\mathbb{N}}$ is uniformly bounded in $\mathscr{L}(L^{q}(\mathcal{M}))$ by Assumption 2.1(i) which we have just proven to hold, yields $\Lambda_{n}\Pi_{n}u\to u$ for all $u\in L^{q}(\mathcal{M})$ as desired. Assumption 2.1(iii) can be verified via direct computation using the definitions. Finally, we have
$\int_{\mathcal{M}}\Lambda_{n}u_{n}(x)$ $v(x)\mathop{}\!\mathrm{d}\mu(x)=\int_{\mathcal{M}}u_{n}(T_{n}(x))v(x)\mathop{}\!\mathrm{d}\mu(x)=\sum_{j=1}^{n}\int_{V_{n}^{(j)}}u_{n}(x_{n}^{(j)})v(x)\mathop{}\!\mathrm{d}\mu(x)$ (3.6)
$=\frac{1}{n}\sum_{j=1}^{n}u_{n}(x_{n}^{(j)})\Pi_{n}v(x_{n}^{(j)})=\int_{\mathcal{M}_{n}}u_{n}(x)\Pi_{n}v(x)\mathop{}\!\mathrm{d}\mu_{n}(x).$ (3.7)
which shows that the adjoint of $\Pi_{n}\in\mathscr{L}(L^{q}(\mathcal{M});L^{q}(\mathcal{M}_{n}))$, where $q\in[1,\infty)$, is given by $\Pi_{n}^{*}=\Lambda_{n}\in\mathscr{L}(L^{q^{\prime}}(\mathcal{M}_{n});L^{q^{\prime}}(\mathcal{M}))$.
The concrete choices of $\mathcal{M}$ and their discretizations which we will consider in this section are the following two:
Setting 3.2 (Square grid on $\mathbb{T}^{m}$).
Let $\mathcal{M}\coloneqq\mathbb{T}^{m}$ be the $m$-dimensional flat torus, which we view as the cube $[0,1]^{m}$ endowed with periodic boundary conditions. For notational convenience, we will index our sequence of discretizations of $\mathbb{T}^{m}$ only by the natural numbers $n$ such that $n^{1/m}\in\mathbb{N}$, for which we define the following square equidistant grid with mesh size $h_{n}\coloneqq n^{-1/m}$:
$\mathcal{M}_{n}\coloneqq\{\tfrac{1}{2}n^{-1/m},\tfrac{3}{2}n^{-1/m},\dots,1-\tfrac{1}{2}n^{-1/m}\}^{m}.$
Then the grid points of $\mathcal{M}_{n}$ can be written as $\mathbf{x}_{n}^{(\mathbf{j})}=n^{-1/m}(\mathbf{j}-\frac{1}{2}\mathbf{1})$ for some $m$-tuple $\mathbf{j}\in\{1,\dots,n^{1/m}\}^{m}$, where $\mathbf{1}=(1,\dots,1)\in\mathbb{R}^{m}$. To each of these points $\mathbf{x}_{n}^{(\mathbf{j})}\in\mathcal{M}_{n}$, we associate the half-open cube $U_{n}^{(\mathbf{j})}\coloneqq\prod_{k=1}^{m}\bigl{[}n^{-1/m}(j_{k}-1),n^{-1/m}j_{k}\bigr{)}$. Since these cubes form a partition of $\mathcal{M}$ (recalling that opposite sides are identified), we can define the transport map $T_{n}\colon\mathcal{M}\to\mathcal{M}_{n}$ by $T_{n}(x)\coloneqq\mathbf{x}_{n}^{(\mathbf{j})}$ whenever $x\in U_{n}^{(\mathbf{j})}$. It readily follows that (3.1) holds, as does (3.2), since $\varepsilon_{n}=\frac{1}{2}\sqrt{m}n^{-1/m}$ for all $n\in\mathbb{N}$.
Setting 3.3 (Randomly sampled point cloud).
Let $\mathcal{M}$ be any manifold satisfying Assumption 3.1, and let $(x^{(n)})_{n\in\mathbb{N}}\subseteq\mathcal{M}$ be a sequence of points independently sampled from $\mu$. This sequence can be viewed as a sample from the product probability space $(\widetilde{\Omega},\widetilde{\mathcal{F}},\widetilde{\mathbb{P}})\coloneqq\prod_{n\in\mathbb{N}}(\mathcal{M},\mathcal{B}(\mathcal{M}),\mu)$, where $\mathcal{B}(\mathcal{M})$ denotes the Borel $\sigma$-algebra on $\mathcal{M}$. If we set $\mathcal{M}_{n}\coloneqq(x^{(j)})_{j=1}^{n}$ for all $n\in\mathbb{N}$, then [71, Proposition 4.1] states that, $\widetilde{\mathbb{P}}$-a.s., there exists a sequence $(T_{n})_{n\in\mathbb{N}}$ of transport maps $T_{n}\colon\mathcal{M}\to\mathcal{M}_{n}$ for which (3.1) holds and
$\varepsilon_{n}\lesssim_{\mathcal{M}}(\log n)^{c_{m}}n^{-1/m}\to 0\quad\text{as }n\to\infty,$ (3.8)
where $c_{m}=\frac{3}{4}$ if $m=2$ and $c_{m}=\frac{1}{m}$ otherwise.
We will now introduce the linear operators which we consider on the domains $\mathcal{M}$ and $(\mathcal{M}_{n})_{n\in\mathbb{N}}$ as in Settings 3.2 and 3.3. Given the coefficient functions $\tau\colon\mathcal{M}\to[0,\infty)$ and $\kappa\colon\mathcal{M}\to[0,\infty)$, respectively assumed to be Lipschitz and continuously differentiable, we consider the nonnegative and symmetric second-order linear differential operator $\mathcal{L}^{\tau,\kappa}_{\infty}$ formally defined by
$\mathcal{L}^{\tau,\kappa}_{\infty}u\coloneqq\tau u-\nabla\cdot(\kappa\nabla u),$ (3.9)
for $u$ belonging to some appropriate domain $\mathsf{D}(\mathcal{L}^{\tau,\kappa}_{\infty})\subseteq L^{2}(\mathcal{M})$.
For each $n\in\mathbb{N}$, we endow $\mathcal{M}_{n}$ with a (weighted, undirected) graph structure by viewing its points as vertices and defining the weight matrix $\mathbf{W}_{n}\in\mathbb{R}^{n\times n}$ by
$(\mathbf{W}_{n})_{ij}\coloneqq\frac{2(m+2)}{\nu_{m}}\frac{1}{nh_{n}^{m+2}}\mathbf{1}_{[0,h_{n}]}\bigl{(}\|x_{n}^{(i)}-x_{n}^{(j)}\|_{\mathbb{R}^{d}}\bigr{)},$ (3.10)
where $\nu_{m}$ denotes the volume of the unit sphere in $\mathbb{R}^{m}$ and $h_{n}\in(0,\infty)$ is a given graph connectivity length scale. With these weights, the resulting graph is an example of a geometric graph (or in fact a random geometric graph if the nodes are sampled randomly as in Setting 3.3). The results in this section are likely to remain valid if the indicator function $\mathbf{1}_{[0,h_{n}]}$ in (3.10) is replaced by a more general (e.g., Gaussian) cut-off kernel (such as in [12]), but we only consider $\eta=\mathbf{1}_{[0,h_{n}]}$ in order to also cite sources which are not formulated in this generality.
The graph-discretized counterpart $\mathcal{L}_{n}^{\tau,\kappa}$ of (3.9) is then the operator which acts on a given function $u\colon\mathcal{M}_{n}\to\mathbb{R}$ as
$\mathcal{L}_{n}^{\tau,\kappa}u(x_{n}^{(i)})\coloneqq\tau(x_{n}^{(i)})u(x_{n}^{(i)})+\sum_{j=1}^{n}(\mathbf{W}_{n})_{ij}\sqrt{\kappa(x_{n}^{(i)})\kappa(x_{n}^{(j)})}\bigl{(}u(x_{n}^{(i)})-u(x_{n}^{(j)})\bigr{)}.$ (3.11)
This can be seen as a generalized version of the (unnormalized) graph Laplacian $\Delta_{\mathcal{M}_{n}}$, and in fact reduces to it if $\tau\equiv 0$ and $\kappa\equiv 1$.
Assumption 3.4 (Coefficients of $\mathcal{L}_{n}^{\tau,\kappa}$).
Let $\tau\colon\mathcal{M}\to[0,\infty)$ and $\kappa\colon\mathcal{M}\to[0,\infty)$ be the coefficient functions used to define the base operators $(\mathcal{L}_{n}^{\tau,\kappa})_{n\in\overline{\mathbb{N}}}$ in (3.9) and (3.11). We shall suppose that
  1. (i)
    $\tau$ is Lipschitz, whereas $\kappa$ is continuously differentiable and bounded below away from zero.
For some results, we specialize to the case that
  1. (ii)
    $\tau\equiv 0$ and $\kappa\equiv 1$, i.e., $\mathcal{L}^{\tau,\kappa}_{n}=\Delta_{\mathcal{M}_{n}}$ and $\mathcal{L}^{\tau,\kappa}_{\infty}$ and reduces to the Laplace–Beltrami operator on $\mathcal{M}$.
Assumption 3.5 (Connectivity length scale of random graph).
Let the manifold $\mathcal{M}$ and the random point clouds $(\mathcal{M}_{n})_{n\in\mathbb{N}}$ on the probability space $(\widetilde{\Omega},\widetilde{\mathcal{F}},\widetilde{\mathbb{P}})$ be as in Setting 3.3. Let $(h_{n})_{n\in\mathbb{N}}\subseteq(0,\infty)$ determine the connectivity length scales of the graphs associated to $(\mathcal{M}_{n})_{n\in\mathbb{N}}$ via the weights (3.10), and suppose that $s\in(0,\infty)$. We will assume one of the following:
  1. (i)
    There exists a $\beta>\frac{m}{4s}$ such that $(\log n)^{c_{m}}n^{-\frac{1}{m}}\ll h_{n}\ll n^{-\frac{1}{4s\beta}}.$
  2. (ii)
    There exists a $\delta>0$, so small that $\frac{m}{1-\delta}<m+4+\delta$, and a $\beta>\frac{m+4+\delta}{2s}$ such that $n^{-\frac{1}{m+4+\delta}}\lesssim h_{n}\ll n^{-\frac{1}{2s\beta}}.$
Given two sequences $(a_{n})_{n\in\mathbb{N}}$ and $(b_{n})_{n\in\mathbb{N}}$ of positive real numbers, the notation $a_{n}\ll b_{n}$ means $a_{n}/b_{n}\to 0$ as $n\to\infty$.
Since $\operatorname{Id}_{n}+\mathcal{L}_{n}^{\tau,\kappa}$ is self-adjoint, positive definite and has a compact inverse for all $n\in\overline{\mathbb{N}}$ (cf. [75, Chapter XII] in the case $n=\infty$), there exists an orthonormal basis $(\psi^{(j)}_{n})_{j=1}^{n}$ of $L^{2}(\mathcal{M}_{n})$ and a non-decreasing sequence $(\lambda^{(j)}_{n})_{j=1}^{n}\subseteq[0,\infty)$, accumulating only at infinity for $n=\infty$, such that $\mathcal{L}_{n}^{\tau,\kappa}\psi^{(j)}_{n}=\lambda^{(j)}_{n}\psi^{(j)}_{n}$ for all $j\in\{1,\dots,n\}$. We summarize this state of affairs by saying that $(\psi^{(j)}_{n},\lambda^{(j)}_{n})_{j=1}^{n}$ is an orthonormal eigenbasis of $L^{2}(\mathcal{M}_{n})$ associated to $\mathcal{L}_{n}^{\tau,\kappa}$. The asymptotic behavior of the eigenvalues $(\lambda_{\infty}^{(j)})_{j\in\mathbb{N}}$ is described by Weyl’s law, cf. [75, Theorem XII.2.1]:
$\lambda^{(j)}_{\infty}\eqsim_{(\mathcal{M},\tau,\kappa)}j^{2/m}\quad\text{for all }j\in\mathbb{N}.$ (3.12)
Given any of the above settings and $n\in\overline{\mathbb{N}}$, we define the generalized Whittle–Matérn operator $A_{n}$ on $L^{2}(\mathcal{M}_{n})$ as a fractional power of the symmetric elliptic operator $\mathcal{L}_{n}^{\tau,\kappa}$ given by (3.9) and (3.11). That is, we set $A_{n}\coloneqq(\mathcal{L}^{\tau,\kappa}_{n})^{s}$ for some $s\in[0,\infty)$, where we use the spectral definition of fractional powers:
$A_{n}u=(\mathcal{L}^{\tau,\kappa}_{n})^{s}u\coloneqq\sum_{j=1}^{n}[\lambda^{(j)}_{n}]^{s}\langle u,\psi_{n}^{(j)}\rangle_{L^{2}(\mathcal{M}_{n})}\psi_{n}^{(j)},\quad u\in\mathsf{D}(A_{n})\subseteq L^{2}(\mathcal{M}_{n}).$ (3.13)
These will be used as the linear operators $(A_{n})_{n\in\overline{\mathbb{N}}}$ in the stochastic partial differential equations in the next subsection.
Since $A_{n}$ is a nonnegative definite and self-adjoint operator on $L^{2}(\mathcal{M}_{n})$ for any $n\in\overline{\mathbb{N}}$, the Lumer–Phillips theorem [81, Theorem 13.35] implies that $-A_{n}$ generates a contractive analytic $C_{0}$-semigroup $(S_{n}(z))_{z\in\Sigma_{\eta}}\subseteq\mathscr{L}(L^{2}(\mathcal{M}_{n}))$ on the sector
$\Sigma_{\eta}\coloneqq\{\lambda\in\mathbb{C}\setminus\{0\}:\arg\lambda\in(-\eta,\eta)\}$ (3.14)
for every $\eta\in(0,\frac{1}{2}\pi)$. Thus, the operators $(A_{n})_{n\in\overline{\mathbb{N}}}$ on $(L^{2}(\mathcal{M}_{n}))_{n\in\overline{\mathbb{N}}}$ are uniformly sectorial of angle 00, see Appendix C.
The following additional assumption(s) on the $L^{\infty}(\mathcal{M}_{n})$-boundedness of the semigroups will be needed for some of the results in Section 3.2:
Assumption 3.6 (Uniform $L^{\infty}$-boundedness of semigroups).
Suppose that
  1. (i)
    there exists a constant $M_{S,\infty}\in[1,\infty)$ such that
    $\|S_{n}(t)\|_{\mathscr{L}(L^{\infty}(\mathcal{M}_{n}))}\leq M_{S,\infty}\quad\text{for all }n\in\overline{\mathbb{N}}\text{ and }t\geq 0.$ (3.15)
We may sometimes additionally assume that
  1. (ii)
    $(S_{n}(t))_{t\geq 0}$ is $L^{\infty}(\mathcal{M}_{n})$-contractive for all $n\in\overline{\mathbb{N}}$, i.e., $M_{S,\infty}=1$ in (i).
Under this assumption, it follows from [68, Proposition 3.12] that $(S_{n}(z))_{z\in\Sigma_{\eta_{q}}}$ is bounded analytic on $L^{q}(\mathcal{M}_{n})$ with $\eta_{q}=\frac{2}{q}\eta$ for all $n\in\overline{\mathbb{N}}$ and $q\in(2,\infty)$, and its uniform norm bound on the sector $\Sigma_{\eta_{q}}$ only depends on $q$ and $M_{S,\infty}$. Therefore, the sequence of operators $(A_{n})_{n\in\overline{\mathbb{N}}}$ on $(L^{q}(\mathcal{M}_{n}))_{n\in\overline{\mathbb{N}}}$ is uniformly sectorial of angle at most $(\frac{1}{2}-\frac{1}{q})\pi$.
3.2. Convergence of graph-discretized semilinear SPDEs
Let $(W(t))_{t\geq 0}$ be an $L^{2}(\mathcal{M})$-valued cylindrical Wiener process with respect to a filtered probability space $(\Omega,\mathcal{F},(\mathcal{F}_{t})_{t\in[0,T]},\mathbb{P})$. The spaces $\mathcal{M}$ and $(\mathcal{M}_{n})_{n\in\mathbb{N}}$ are as in Setting 3.2 or 3.3 above; in the latter case, note that the space $(\Omega,\mathcal{F},(\mathcal{F}_{t})_{t\in[0,T]},\mathbb{P})$ associated to the Wiener noise is independent of the probability space $(\widetilde{\Omega},\widetilde{\mathcal{F}},\widetilde{\mathbb{P}})$ describing the randomness of the point cloud. For every $n\in\overline{\mathbb{N}}$, we set $W_{n}\coloneqq\Pi_{n}W$ and consider the following semilinear stochastic partial differential equation (SPDE):
[…]
where $s\in(0,\infty)$, $T\in(0,\infty)$ is a finite time horizon, $f_{n}\colon\Omega\times[0,T]\times\mathbb{R}\to\mathbb{R}$ is the nonlinearity, and $\xi_{n}\colon\Omega\times\mathcal{M}_{n}\to\mathbb{R}$ is the initial datum. Note that $(W_{n}(t))_{t\geq 0}$ is a $Q_{n}$-cylindrical Wiener process with $Q_{n}=\Pi_{n}^{*}\Pi_{n}=\Lambda_{n}\Pi_{n}=\operatorname{Id}_{n}$, where we recall (3.7) for the second identity. Therefore, $(W_{n}(t))_{t\geq 0}$ is a cylindrical Wiener process on $L^{2}(\mathcal{M}_{n})$ for all $n\in\overline{\mathbb{N}}$, and its formal time derivative $\mathop{}\!\mathrm{d}W_{n}$ represents spatiotemporal Gaussian white noise on $[0,T]\times\mathcal{M}_{n}$.
Solutions to (3.16)—and all the other (semi)linear SPDEs that we consider in this work—are always interpreted in the mild sense. This notion of solutions is defined using the semigroup $(S_{n}(t))_{t\geq 0}$ generated by $-[\mathcal{L}^{\tau,\kappa}_{n}]^{s}$. We say that $u_{n}$ is a global mild solution to (3.16) if it satisfies the following relation for all $t\in[0,T]$:
[…]
Here, we interpret $u_{n}=(u_{n}(t))_{t\in[0,T]}$ as a process taking its values in an infinite-dimensional Banach space of functions on $\mathcal{M}_{n}$ (such as $L^{q}(\mathcal{M}_{n})$ or $C(\mathcal{M}_{n})$), and we define for every $(\omega,t)\in\Omega\times[0,T]$ the Nemytskii operator $u_{n}\mapsto F_{n}(\omega,t,u_{n})$ on this function space by setting $[F_{n}(\omega,t,u_{n})](\xi)\coloneqq f_{n}(\omega,t,u_{n}(\xi))$ for all $\xi\in\mathcal{M}_{n}$.
This notion of solution is called “global” because it exists on the whole of $[0,T]$, in contrast with “local” solutions, which may blow up before time $T$. However, we note that global solutions generally grow unbounded as $T\to\infty$. We will not consider local solutions in this section, but we do work with them in Section 5.2.
In this section, the real-valued functions $f_{n}$ are supposed to satisfy the following:
Assumption 3.7 (Nonlinearities).
We will assume one of the following conditions:
  1. (i)
    The nonlinearities $(f_{n})_{n\in\overline{\mathbb{N}}}$ are globally Lipschitz continuous and grow linearly, both uniformly in $n$. I.e., there exist $\widetilde{L}_{f},\widetilde{C}_{f}\in[0,\infty)$ such that, for all $n\in\overline{\mathbb{N}}$ and $x,y\in\mathbb{R}$,
    $\lvert f_{n}(\omega,t,x)-f_{n}(\omega,t,y)\rvert\leq\widetilde{L}_{f}\lvert x-y\rvert\quad\text{and}\quad\lvert f_{n}(\omega,t,x)\rvert\leq\widetilde{C}_{f}(1+\lvert x\rvert).$ (3.18)
  2. (ii)
    The nonlinearities $(f_{n})_{n\in\overline{\mathbb{N}}}$ are of the polynomial form
    $f_{n}(\omega,t,x)\coloneqq-a_{2k+1,n}(\omega,t)x^{2k+1}+\sum_{j=0}^{2k}a_{j,n}(\omega,t)x^{j},$ (3.19)
    where $k\in\mathbb{N}_{0}$ and $a_{j,n}\colon\Omega\times[0,T]\to\mathbb{R}$ for each $j\in\{0,\dots,2k+1\}$, and there exist constants $c,C\in(0,\infty)$ such that
    $c\leq a_{2k+1,n}(\omega,t)\leq C\quad\text{and}\quad\lvert a_{j,n}(\omega,t)\rvert\leq C$ (3.20)
    for all $j\in\{0,\dots,2k\}$, $n\in\overline{\mathbb{N}}$ and $(\omega,t)\in\Omega\times[0,T]$.
In either case, we suppose moreover that $f_{n}\to f$ uniformly on compact intervals; i.e., for all $r\in[0,\infty)$ and $(\omega,t)\in\Omega\times[0,T]$,
$\sup_{x\in[-r,r]}\lvert f_{n}(\omega,t,x)-f_{\infty}(\omega,t,x)\rvert\to 0\quad\text{as }n\to\infty.$ (3.21)
Example 3.8.
The cubic polynomial $f_{n}(\omega,t,x)\coloneqq-x^{3}+x$, which turns (3.16) into the (fractional) stochastic Allen–Cahn equation in case $\tau\equiv 0$, $\kappa\equiv 1$ and $s\leq 1$, is of the form asserted in Assumption 3.7(ii). Note that this is also an example of the important situation where (3.21) is trivially satisfied by taking the same function $f_{n}\coloneqq f$ for all $n\in\overline{\mathbb{N}}$.
The final technical assumption that we record before moving on to the main theorem of this section is the following:
Assumption 3.9 (Uniform $L^{\infty}$-boundedness of eigenfunctions).
There exists a constant $M_{\psi,\infty}\in(0,\infty)$ such that
$\|\psi_{n}^{(j)}\|_{L^{\infty}(\mathcal{M}_{n})}\leq M_{\psi,\infty}\quad\text{for all }n\in\overline{\mathbb{N}}\text{ and }j\in\{1,\dots,n\}.$ (3.22)
The interplay of the various choices of spatial domains $\mathcal{M}_{n}$, linear operators $A_{n}$ and nonlinearity functions $f_{n}$ determines the class of SPDEs to which (3.16) belongs. Rigorous definitions of the corresponding mild solution concepts, as well as well-posedness and discrete-to-continuum convergence results can be found in Sections 46, respectively. Applying these results in their respective regimes of applicability yields the following discrete-to-continuum convergence theorem for the solutions to (3.16); note that the setting of part c covers the stochastic (fractional) Allen–Cahn equation on the one-dimensional torus, see Example 3.8.
Theorem 3.10.
Let $\mathcal{M}$ and $(\mathcal{M}_{n})_{n\in\overline{\mathbb{N}}}$ be as in Setting 3.2 or 3.3.
  1. (a)
    Consider Setting 3.3. Let $s>\frac{1}{2}m$ and suppose that Assumption 3.5(i) holds with $\beta\in(\frac{m}{4s},\frac{1}{2})$. If Assumptions 3.4(i) and 3.7(i) are satisfied, and $p\in[1,\infty)$ is such that $\Lambda_{n}\xi_{n}\to\xi_{\infty}$ in $L^{p}(\Omega,\mathcal{F}_{0},\mathbb{P};L^{2}(\mathcal{M}))$, then there exists a unique global mild solution $u_{n}$ in $L^{p}(\Omega;C([0,T];L^{2}(\mathcal{M}_{n})))$ to (3.16) for every $n\in\overline{\mathbb{N}}$, and as $n\to\infty$ we have
    $\Lambda_{n}u_{n}\to u_{\infty}\quad\widetilde{\mathbb{P}}\text{-a.s.\ in }L^{p}(\Omega;C([0,T];L^{2}(\mathcal{M}))).$ (3.23)
  2. (b)
    In Setting 3.3, let $\delta>0$ be such that $\frac{m}{1-\delta}<m+4+\delta$, suppose $s>m+4+\delta$ and Assumption 3.5(ii) holds with $\beta\in(\frac{m+4+\delta}{2s},\frac{1}{2})$. Let Assumptions 3.4(ii), 3.6(i), 3.7(i) and 3.9 be satisfied. If $p\in[1,\infty)$ is such that $\Lambda_{n}\xi_{n}\to\xi_{\infty}$ in $L^{p}(\Omega,\mathcal{F}_{0},\mathbb{P};L^{\infty}(\mathcal{M}))$, then there exists a unique global mild solution $u_{n}$ in $L^{p}(\Omega;C([0,T];L^{\infty}(\mathcal{M}_{n})))$ to (3.16) for every $n\in\mathbb{N}$, as well as $u_{\infty}$ in $L^{p}(\Omega;C([0,T];C(\mathcal{M})))$ for $n=\infty$, and as $n\to\infty$ we have
    $\Lambda_{n}u_{n}\to u_{\infty}\quad\text{in }L^{0}(\widetilde{\Omega},L^{p}(\Omega;C([0,T];L^{\infty}(\mathcal{M})))).$ (3.24)
  3. (c)
    Consider Setting 3.2 with $\mathcal{M}\coloneqq\mathbb{T}$. Let $s\in(\frac{1}{2},1]$ and suppose that Assumptions 3.4(ii) and 3.7(ii) are satisfied. If $p\in(1,\infty)$ is such that $\Lambda_{n}\xi_{n}\to\xi_{\infty}$ in $L^{p}(\Omega,\mathcal{F}_{0},\mathbb{P};L^{\infty}(\mathbb{T}))$, then there exists a unique global mild solution $u_{n}$ in $L^{p}(\Omega;C([0,T];L^{\infty}(\mathcal{M}_{n})))$ to (3.16) for every $n\in\mathbb{N}$, as well as $u_{\infty}$ in $L^{p}(\Omega;C([0,T];C(\mathbb{T})))$ for $n=\infty$ and for all $p^{-}\in[1,p)$ we have, as $n\to\infty$,
    $\Lambda_{n}u_{n}\to u_{\infty}\quad\text{in }L^{p^{-}}\!(\Omega;C([0,T];L^{\infty}(\mathbb{T}))).$ (3.25)
The proof is presented in Section 3.4. In the next section, we list the intermediate results on which it relies. The motivations behind the various assumptions listed above, as well as their role in Theorem 3.10, are discussed in Section 3.5.
3.3. Intermediate results
In this subsection, we collect a number of intermediate results which imply that the conditions imposed in Theorem 3.10 are sufficient to fit into the setting of the various convergence theorems in Sections 46. More precisely, depending on the setting, we wish to verify a subset of the following: Conditions (A1)(A3) from Section 4 on the linear operators, conditions (F1)(F2) and (IC) from Section 5 on the nonlinearities and initial conditions, respectively, as well as their extended counterparts (A1-B)(A4-B), (IC-B), (F1-B)(F2-B) and (F1′′-B) from Section 6. The proofs of the results in this section are deferred to Appendix A for ease of exposition.
The necessary convergence of the linear operators, given by (A3) and (A3-B), will ultimately be derived from the spectral convergence of $(\mathcal{L}_{n}^{\tau,\kappa})_{n\in\mathbb{N}}$ to $\mathcal{L}_{\infty}^{\tau,\kappa}$, i.e., the convergence of the respective eigenvalues and (lifted) eigenfunctions. In the square grid Setting 3.2, we can argue directly using closed-form expressions of all the eigenvalues and eigenfunctions involved, see Lemma 3.11 below. A subtlety arising in the random graph Setting 3.3 is that, for any $n\in\mathbb{N}$, we cannot in general control the errors $\lvert\lambda_{n}^{(j)}-\lambda_{\infty}^{(j)}\rvert$ and $\|\psi_{n}^{(j)}\circ T_{n}-\psi_{\infty}^{(j)}\|_{L^{q}(\mathcal{M})}$ for all $j\in\{1,\dots,n\}$, but only for indices $j$ up to a sufficiently small integer $k_{n}$. We present the precise statements below: Theorems 3.12 and 3.13a, which cover eigenvalue convergence and $L^{2}(\mathcal{M})$-convergence of eigenfunctions, are respectively taken from [71, Theorems 4.6 and 4.7]. Theorem 3.13b, concerning the $L^{\infty}(\mathcal{M})$-convergence of Laplacian eigenvalues, is a consequence of the main results from [12], as shown in [72, Lemma 15] and the discussion preceding it.
Lemma 3.11 (Spectral convergence—square grid).
Let $\mathcal{M}=\mathbb{T}^{m}$ be discretized by the sequence of square grids described in Setting 3.2. If $\tau\equiv 0$ and $\kappa\equiv 1$, then for all $n\in\mathbb{N}$ such that $n^{1/m}\in\mathbb{N}$, the eigenfunction–eigenvalue pairs $(\psi^{(j)}_{n},\lambda^{(j)}_{n})_{j=1}^{n}$ and $(\psi^{(j)}_{\infty},\lambda^{(j)}_{\infty})_{j\in\mathbb{N}}$ corresponding to the graph Laplacian $\mathcal{L}_{n}^{\tau,\kappa}=\Delta_{n}$ and the Laplace–Beltrami operator $\mathcal{L}_{\infty}^{\tau,\kappa}=-\Delta_{\mathcal{M}}$, respectively, satisfy
$0\leq\lambda_{\infty}^{(j)}-\lambda_{n}^{(j)}$ $\leq\tfrac{1}{12}j^{4}\pi^{4}n^{-\frac{2}{m}}\quad\text{for all }j\in\{1,\dots,n\};$ (3.26)
$\|\psi_{\infty}^{(j)}-\psi_{n}^{(j)}\circ T_{n}\|_{L^{\infty}(\mathcal{M})}$ $\leq\tfrac{1}{2}\sqrt{2}j\pi n^{-\frac{1}{m}}\quad\text{for all }j\in\{1,\dots,n-1\}.$ (3.27)
Theorem 3.12 (Eigenvalue convergence—random graphs).
Let the manifold $\mathcal{M}$ and the random point clouds $(\mathcal{M}_{n})_{n\in\mathbb{N}}\subseteq\mathcal{M}$ on the probability space $(\widetilde{\Omega},\widetilde{\mathcal{F}},\widetilde{\mathbb{P}})$ be as in Setting 3.3. Suppose that $\tau\colon\mathcal{M}\to[0,\infty)$ is Lipschitz, and that $\kappa\colon\mathcal{M}\to[0,\infty)$ is continuously differentiable and bounded below away from zero.
If the graph connectivity length scales $(h_{n})_{n\in\mathbb{N}}$ (see (3.10)) are chosen in such a way that there exist positive integers $(k_{n})_{n\in\mathbb{N}}$ satisfying
$\varepsilon_{n}\ll h_{n}\ll[\lambda_{\infty}^{(k_{n})}]^{-\frac{1}{2}},$ (3.28)
then there exists a constant $C_{(\mathcal{M},\tau,\kappa)}>0$ such that
~$\widetilde{\mathbb{P}}\biggl{(}\frac{\lvert\lambda_{n}^{(j)}-\lambda_{\infty}^{(j)}\rvert}{\lambda^{(j)}_{\infty}+1}\leq C_{(\mathcal{M},\tau,\kappa)}\;\varepsilon_{n}h_{n}^{-1}+h_{n}[\lambda_{\infty}^{(j)}]^{\frac{1}{2}}\text{ for all }n\in\mathbb{N},\;j\in\{1,\dots,k_{n}\}\biggr{)}=1.$ (3.29)
Theorem 3.13 (Eigenfunction convergence—random graphs).
Let the manifold $\mathcal{M}$ and the random point clouds $(\mathcal{M}_{n})_{n\in\mathbb{N}}\subseteq\mathcal{M}$ on the probability space $(\widetilde{\Omega},\widetilde{\mathcal{F}},\widetilde{\mathbb{P}})$ be as in Setting 3.3. Let $(h_{n})_{n\in\mathbb{N}}\subseteq(0,\infty)$ be the connectivity length scales of the graphs associated to $(\mathcal{M}_{n})_{n\in\mathbb{N}}$ via the weights (3.10), and consider the (graph-discretized) differential operators $\mathcal{L}_{n}^{\tau,\kappa}$ with coefficients $\tau\colon\mathcal{M}\to[0,\infty)$ and $\kappa\colon\mathcal{M}\to[0,\infty)$.
  1. (a)
    If Assumption 3.4(i) holds, and there exist integers $(k_{n})_{n\in\mathbb{N}}$ such that (3.28) is satisfied, then there exists a constant $C_{(\mathcal{M},\tau,\kappa)}>0$ such that, for all $n\in\mathbb{N}$,
    ~$\widetilde{\mathbb{P}}\Bigl{(}\|\psi_{n}^{(j)}\circ T_{n}-\psi_{\infty}^{(j)}\|_{L^{2}(\mathcal{M})}\leq C_{(\mathcal{M},\tau,\kappa)}$ $j^{\frac{3}{2}}\bigl{(}\varepsilon_{n}h_{n}^{-1}+h_{n}[\lambda_{\infty}^{(j)}]^{\frac{1}{2}}\bigr{)}^{\frac{1}{2}}$ (3.30)
[…]
  2. (b)
    Let Assumption 3.4(ii) be satisfied. If there exist $(k_{n})_{n\in\mathbb{N}}\subseteq\mathbb{N}$ and $\delta>0$ such that
    $n^{-\frac{1}{m+4+\delta}}\lesssim_{\mathcal{M}}h_{n}\lesssim_{\mathcal{M}}[\lambda_{\infty}^{(k_{n})}]^{-1}\quad\text{and}\quad\lambda_{\infty}^{(k_{n})}\lesssim_{\mathcal{M}}n^{\frac{1-\delta}{m}},$ (3.31)
    then there exists a constant $C_{\mathcal{M}}>0$ such that, as $n\to\infty$,
    ~$\widetilde{\mathbb{P}}\Bigl{(}\|\psi_{n}^{(j)}\circ T_{n}-\psi_{\infty}^{(j)}\|_{L^{\infty}(\mathcal{M})}\leq C_{\mathcal{M}}[\lambda_{\infty}^{(j)}]^{m+1}j^{\frac{3}{2}}$ $\bigl{(}\varepsilon_{n}h_{n}^{-1}+h_{n}[\lambda_{\infty}^{(j)}]^{\frac{1}{2}}\bigr{)}^{\frac{1}{2}}$ (3.32)
[…]
From the above results, we can derive the following convergence of the sequence $(A_{n})_{n\in\overline{\mathbb{N}}}$. Its proof is analogous to that of [71, Theorem 4.2], see Appendix A.
Theorem 3.14.
Given $\tau\colon\mathcal{M}\to[0,\infty)$, $\kappa\colon\mathcal{M}\to[0,\infty)$ and $s\in[0,\infty)$, consider the generalized Whittle–Matérn operators $A_{n}\coloneqq(\mathcal{L}_{n}^{\kappa,\tau})^{s}$ defined by (3.13), and set $\widetilde{R}_{n}^{\alpha}\coloneqq\Lambda_{n}(\operatorname{Id}_{n}+A_{n})^{-\alpha}\Pi_{n}$, for all $\alpha\in[0,\infty)$ and $n\in\overline{\mathbb{N}}$.
  1. (a)
    Suppose that $\mathcal{M}$ and its discretizations $(\mathcal{M}_{n})_{n\in\mathbb{N}}$ are as in Setting 3.3, Assumption 3.5(i) holds with $\beta\in(\frac{m}{4s},\infty)$ and Assumption 3.4(i) is satisfied. Then we have for all $\beta^{\prime}\in[\beta,\infty)$:
    $\widetilde{R}_{n}^{\beta^{\prime}}\to R_{\infty}^{\beta^{\prime}}\quad\widetilde{\mathbb{P}}\text{-a.s.}\text{ in }\mathscr{L}_{2}(L^{2}(\mathcal{M}))\quad\text{as }n\to\infty.$ (3.33)
  2. (b)
    Suppose that $\mathcal{M}$ and its discretizations $(\mathcal{M}_{n})_{n\in\mathbb{N}}$ are as in Setting 3.3, Assumption 3.5(ii) holds with $\beta\in(\frac{m+4+\delta}{2s},\infty)$, and Assumptions 3.4(ii) and 3.9 are satisfied. Then we have for all $\beta^{\prime}\in[\beta,\infty)$:
    $\widetilde{R}_{n}^{\beta^{\prime}}\to R_{\infty}^{\beta^{\prime}}\quad\text{in}L^{0}(\widetilde{\Omega};\mathscr{L}(L^{2}(\mathcal{M});L^{\infty}(\mathcal{M})))\quad\text{as }n\to\infty.$ (3.34)
    Here, $L^{0}(\widetilde{\Omega})$ denotes convergence in probability with respect to $(\widetilde{\Omega},\widetilde{\mathcal{F}},\widetilde{\mathbb{P}})$.
  3. (c)
    Suppose that $\mathcal{M}\coloneqq\mathbb{T}^{m}$ is discretized using the square grids $(\mathcal{M}_{n})_{n\in\mathbb{N}}$ from Setting 3.2, and that Assumption 3.4(ii) holds. For all $\beta\in(\frac{m}{4s},\infty)$,
    $\widetilde{R}_{n}^{\beta}\to R_{\infty}^{\beta}\quad\text{in }\mathscr{L}(L^{2}(\mathcal{M});L^{\infty}(\mathcal{M}))\quad\text{as }n\to\infty.$ (3.35)
The following property, which we call the uniform ultracontractivity of the semigroups $(S_{n})_{n\in\overline{\mathbb{N}}}$, will be needed in order to obtain the $L^{\infty}(\mathcal{M})$-convergence in Theorem 3.10b and c. Its proof relies on Riesz–Thorin interpolation, Assumption 3.6, and some arguments from Theorem 3.14.
Lemma 3.15 (Uniform ultracontractivity).
Let $s\in(0,\infty)$ and consider the generalized Whittle–Matérn operators $A_{n}\coloneqq(\mathcal{L}_{n}^{\kappa,\tau})^{s}$ defined by (3.13) for all $n\in\overline{\mathbb{N}}$. Assume either of the following statements:
  1. (a)
    In Setting 3.3, Assumption 3.5(i) or (ii) holds with corresponding $\beta$, as well as Assumptions 3.4(i), 3.6(i) and 3.9.
  2. (b)
    In Setting 3.2, $\beta\in(\frac{m}{4s},\infty)$ is arbitrary, and Assumptions 3.4(ii) and 3.6(i) hold.
Then, for every $q\in[2,\infty]$, there exists $M_{S,q}\in[1,\infty)$ such that
$\|S_{n}(t)\|_{\mathscr{L}(L^{q}(\mathcal{M}_{n});L^{\infty}(\mathcal{M}_{n}))}\leq M_{S,q}t^{-\frac{2}{q}\beta}\quad\text{for all }n\in\overline{\mathbb{N}}\text{ and }t>0.$ (3.36)
In case of a, (3.36) holds $\widetilde{\mathbb{P}}$-a.s.
3.4. Proof of convergence
Using the intermediate results from Subsection 3.3, we can now prove Theorem 3.10:
Proof of Theorem 3.10.
To prove parts ac, we will apply Theorems 5.46.6 and Corollary 6.10, respectively, which are the rigorous counterparts of the discrete-to-continuum Theorem 1.1 in the respective settings.
The argument preceding Setting 3.2 shows that (A1) and (A1-B) hold in any of the given situations, with $H_{n}\coloneqq L^{2}(\mathcal{M}_{n})$ and $E_{n}\coloneqq L^{q}(\mathcal{M}_{n})$ for $n\in\overline{\mathbb{N}}$ and $q\in[2,\infty)$, as well as $B_{n}\coloneqq L^{\infty}(\mathcal{M}_{n})$ for all $n\in\mathbb{N}$, $B_{\infty}\coloneqq C(\mathcal{M})$ and $\widetilde{B}\coloneqq L^{\infty}(\mathcal{M})$. Moreover, note that (IC) (or (IC-B)) is explicitly assumed in each case.
a Here, we take $q=2$, i.e., $E_{n}=H_{n}=L^{2}(\mathcal{M}_{n})$ for all $n\in\overline{\mathbb{N}}$. As discussed at the end of Subsection 3.1, the operators $(A_{n})_{n\in\overline{\mathbb{N}}}\coloneqq([\mathcal{L}_{n}^{\tau,\kappa}]^{s})_{n\in\overline{\mathbb{N}}}$ are uniformly sectorial of angle 00 on $(L^{2}(\mathcal{M}_{n}))_{n\in\overline{\mathbb{N}}}$. Letting $\beta\in(\frac{m}{4s},\frac{1}{2})$ be as in Assumption 3.5(i), it follows from Theorem 3.14a that $\widetilde{R}_{n}^{\beta^{\prime}}\to R_{\infty}^{\beta^{\prime}}$, $\widetilde{\mathbb{P}}$-a.s., in $\mathscr{L}_{2}(L^{2}(\mathcal{M}))$ as $n\to\infty$, for all $\beta^{\prime}\geq\beta$. Applying this with $\beta^{\prime}\coloneqq\beta\in(0,\frac{1}{2})$ and $\beta^{\prime}\coloneqq 1$ yields (A2) and (A3). Setting, for all $(\omega,t)\in\Omega\times[0,T]$, $u\in L^{2}(\mathcal{M}_{n})$ and $x\in\mathcal{M}_{n}$,
$[F_{n}(\omega,t,u)](x)\coloneqq f_{n}(\omega,t,u(x)),$ (3.37)
it is immediate from Assumption 3.7(i) that (F1) is satisfied. Moreover, combining the definition of $\widetilde{F}_{n}$ from (5.8) with (3.37) yields
$[\widetilde{F}_{n}(\omega,t,u)](x)=[\Lambda_{n}F_{n}(\omega,t,\Pi_{n}u)](x)=f_{n}(t,\omega,\Lambda_{n}\Pi_{n}u(x)),$ (3.38)
so that
$\|\widetilde{F}_{n}(\omega,t,u)-F_{\infty}(\omega,t,u)\|_{L^{2}(\mathcal{M})}=\|f_{n}(\omega,t,\Lambda_{n}\Pi_{n}u(\,\cdot\,))-f_{\infty}(\omega,t,u(\,\cdot\,))\|_{L^{2}(\mathcal{M})}$ (3.39)
$\qquad\leq\widetilde{L}_{f}\|\Lambda_{n}\Pi_{n}u-u\|_{L^{2}(\mathcal{M})}+\|f_{n}(\omega,t,u(\,\cdot\,))-f_{\infty}(\omega,t,u(\,\cdot\,))\|_{L^{2}(\mathcal{M})}.$
As $n\to\infty$, the first term vanishes by Assumption (A1), and the second term by dominated convergence using (3.21) and the uniform linear growth condition in Assumption 3.7(i). Therefore, condition (F2) is also satisfied.
b Now we need Assumption 3.6(i) in order for $([\mathcal{L}_{n}^{\tau,\kappa}]^{s})_{n\in\overline{\mathbb{N}}}$ to be uniformly sectorial of angle less than $\frac{1}{2}\pi$ on $(L^{q}(\mathcal{M}_{n}))_{n\in\overline{\mathbb{N}}}$ for all $q\in[2,\infty)$. Letting $\delta>0$ and $\beta\in(\frac{m+4+\delta}{2s},\frac{1}{2})$ be as in Assumption 3.5(ii), it follows from Theorem 3.14b that $\widetilde{R}_{n}^{\beta^{\prime}}\to R_{\infty}^{\beta^{\prime}}$ in $L^{0}(\widetilde{\Omega};\mathscr{L}(L^{2}(\mathcal{M});L^{\infty}(\mathcal{M})))$ as $n\to\infty$, for all $\beta^{\prime}\geq\beta$, under Assumptions 3.4(ii) and 3.9. In particular, we have $\widetilde{R}_{n}^{\beta}\to R_{\infty}^{\beta}$ in $L^{0}(\widetilde{\Omega};\gamma(L^{2}(\mathcal{M});L^{q}(\mathcal{M})))$ for all $q\in[1,\infty)$ by [49, Corollary 9.3.3], and $\widetilde{R}_{n}\to R_{\infty}$ in $L^{0}(\widetilde{\Omega};\mathscr{L}(L^{\infty}(\mathcal{M})))$. This shows (A2-B) and (A3-B). By Lemma 3.15a, we have (A4-B) with $\theta=\frac{4}{q}\beta$. Thus, choosing $q>\frac{4\beta}{1-2\beta}$ yields $\theta+2\beta<1$. Conditions (F1-B) and (F2-B) follow similarly to part a.
c As in part b, we need to verify conditions (A1-B)(A4-B), now with contractive semigroups $(S_{n}(t))_{t\geq 0}$, i.e., Assumption 3.6(ii). For $s=1$, $(S_{\infty}(t))_{t\geq 0}$ is $L^{\infty}(\mathbb{T})$-contractive since the $L^{1}(\mathbb{T})$-norm of its heat kernel coincides with the $L^{1}(\mathbb{R})$-norm of the Gauss–Weierstrass kernel, which is equal to $1$. For finite $n$, the $L^{\infty}(\mathcal{M}_{n})$-contractivity of $S_{n}(t)=e^{-tA_{n}}$ is equivalent to $A_{n}$ being diagonally dominant with positive diagonal by [64, Lemma 6.1], which holds for Laplacian matrices. Since these assertions can be extended to all $s\in(0,1]$ by subordination, see for instance [50, Theorem 15.2.17], we indeed find that Assumption 3.6(ii) holds. Thus, we can proceed to argue as in b, using Theorem 3.14c and Lemma 3.15b for an arbitrary $\beta\in(\frac{1}{4s},\frac{1}{2})$, to obtain (A1-B)(A4-B) with $\widetilde{M}_{S}=1$ and $\theta+2\beta<1$ for $q\in[2,\infty)$ large enough.
It remains to establish that the nonlinearities from Assumption 3.7(ii) are such that (F1′′-B) holds. This is done in Example 6.8, noting that the space $L^{\infty}(\mathcal{M}_{n})$ coincides with $C(\mathcal{M}_{n})$ if we equip $\mathcal{M}_{n}$ with the discrete topology. ∎
3.5. Discussion of the assumptions
In this subsection, we comment on the various assumptions made in Theorem 3.10, the extent to which they are necessary, and how one might check them in practice.
The distinction between parts (i) and (ii) of Assumption 3.4, i.e., whether to allow for spatially varying coefficient functions $\tau$ and $\kappa$ in the second-order symmetric base operators $(\mathcal{L}_{n}^{\tau,\kappa})_{n\in\overline{\mathbb{N}}}$ instead of merely considering Laplacians, is mainly due to the availability of spectral convergence theorems in the respective situations. Most of the literature on eigenfunction convergence of graph-discretized second-order operators is focused on the Laplacian case, see for instance [11, 32] for $L^{2}$-convergence and [12, 25, 86] for $L^{\infty}$-convergence. However, the authors of [71] show how the $L^{2}$-convergence results can be extended to coefficient functions satisfying Assumption 3.4(i). We expect that most spectral convergence results for graph Laplacians can be extended to allow for varying coefficients, but doing so requires significant effort, hence we sometimes make Assumption 3.4(ii) for the sake of convenience.
Similarly, the difference between the bounds on the graph connectivity length scales in the two parts of Assumption 3.5 is a result of the current availability of spectral convergence literature. Eigenfunction convergence of $(\mathcal{L}_{n}^{\tau,\kappa})_{n\in\overline{\mathbb{N}}}$ in $L^{2}$ has for instance been proved in [71] under Assumption 3.5(i), but for graphs and manifolds as in our setting, the optimal available $L^{\infty}$-convergence results (for graph Laplacians) seem to be those of [12], which require Assumption 3.5(ii). However, according to [12, Remark 2.7], it is plausible that the $L^{\infty}(\mathcal{M})$-convergence of Laplacian eigenfunctions can be proved under the same assumptions as the $L^{2}(\mathcal{M})$-convergence, with the same rate. Some recent results in this direction can be found in [3], where the authors show $L^{\infty}$-convergence of Laplacian eigenvectors with optimal rates and loose lower bounds on the connectivity lengths, using homogenization theory, for point clouds on less general spatial domains.
Assumption 3.6 is natural in the sense that the results regarding $L^{\infty}$-convergence in space (for instance Theorem 3.10b and c) rely on uniform $L^{\infty}$-convergence of semigroup orbits on compact time intervals. The latter necessitates that Assumption 3.6(i) is satisfied, at least for $t\in[0,T]$ with arbitrarily large $T\in(0,\infty)$.
Moreover, for typical choices of differential operators, one can often check that Assumption 3.6(ii) holds, meaning that the semigroups are in fact $L^{\infty}$-contractive. One such example is outlined in the proof of Theorem 3.10c: Matrix exponentials $(e^{-tL_{n}})_{t\geq 0}$ are $L^{\infty}$-contractive if and only if $L_{n}\in\mathbb{R}^{n\times n}$ is diagonally dominant with nonnegative diagonal entries [64, Lemma 6.1]. Sufficient conditions for the $L^{\infty}$-contractivity of the semigroup $(S_{\infty}(t))_{t\geq 0}$ generated by the negative of a uniformly elliptic second-order differential operator on a Euclidean domain $\mathcal{D}\subsetneq\mathbb{R}^{d}$, subject to appropriate boundary conditions, can be found in [68, Section 4.3]. Likewise, the heat semigroup associated to the Laplace–Beltrami operator on a compact Riemannian manifold $\mathcal{M}$ is $L^{\infty}$-contractive, cf. [23, p. 148].
As mentioned in the proof of Theorem 3.10c, all of the above $L^{\infty}$-contractivity results for second-order differential operators can be extended to fractional powers $s\in(0,1)$ by using a subordination formula such as [50, Theorem 15.2.17], noting that the definition of fractional power operators in this reference coincides with ours (more details are given in the first half of the proof of Lemma C.1 below). The semigroups generated by higher-order differential operators, however, are in general not contractive on $L^{\infty}$ (or any $L^{q}$ for $q\neq 2$, see for instance [57]); this is closely related to their lack of positivity preservation. As an example, the fractional heat kernel associated to $(-\Delta)^{s}$ on $\mathbb{R}^{d}$ with $s\in(0,\infty)$ at time $t\in(0,\infty)$ is given by the inverse Fourier transform of $\xi\mapsto\exp(-t\|\xi\|_{\mathbb{R}^{d}}^{2s})$, which is positive for $s\leq 1$ but fails to be sign-definite if $s>1$, see [28, p. 626 and pp. 632–633], respectively.
Thus, for operators $([\mathcal{L}_{n}^{\tau,\kappa}]^{s})_{n\in\overline{\mathbb{N}}}$ with $s>1$, we have to content ourselves with uniform $L^{\infty}$-boundedness of the semigroups $(S_{n}(t))_{t\geq 0}$ in $n$ and $t$, i.e., Assumption 3.6(i). In the absence of positivity preservation, one route to verifying such uniformity is through Gaussian upper bounds on the integral kernels corresponding to the semigroups, cf. [68, Proposition 7.1]. Such bounds have been established for higher-order differential operators on Euclidean domains, as well as Laplacian operators on more general domains such as manifolds, graphs and fractals (see [68, pp. 194–196] and the references therein). While it may be possible to unify these results in the setting of graph-discretized higher-order differential operators on manifolds, and thus obtain the uniform $L^{\infty}$-bounds required by Assumption 3.6(i), this appears to be highly nontrivial and outside the scope of this work.
We also remark that certain higher-order operators have been shown to exhibit (local) eventual positivity, meaning that for every nonnegative initial datum $u_{0}\geq 0$ and subset $\mathcal{D}^{*}$ of the spatial domain $\mathcal{D}$, there exists $t^{*}>0$ such that $S(t)u_{0}\geq 0$ on $\mathcal{D}^{*}$ for all $t\geq t^{*}$. For instance, in [34], this was shown for the bi-Laplacian $\Delta^{2}$ on $\mathcal{D}=\mathbb{R}^{d}$. In [38], the authors apply the theory of [21, 22] to treat the squared graph Laplacian $\Delta_{n}^{2}$, deduce that $(e^{-\Delta_{n}^{2}t})_{t\geq 0}$ is eventually $L^{\infty}$-contractive [38, Proposition 6.7], and note that this implies $L^{\infty}$-boundedness uniformly in $t\geq 0$ [38, Remark 6.8]. However, as $n\to\infty$, their upper bound $\|e^{-\Delta_{n}^{2}t}\|_{\infty}\leq\exp(\|\Delta_{n}\|_{\infty}^{2}t^{*})$ blows up in our setting. Hence, these results do not appear to be directly useful for our purpose of verifying Assumption 3.6(i).
If we restrict ourselves to nonlinearities of the form $[F_{n}(\omega,t,u)](x)\coloneqq f_{n}(\omega,t,u(x))$ (see (5.1) and (6.1)), then the conditions in Assumption 3.7 are the natural ones to ensure the global (in time) convergence results formulated in Theorem 3.10. Convergence results for more general nonlinearities, possibly formulated only in terms of local-in-time convergence and in weaker norms, can be found in Sections 5 and 6.
[…]
4. Infinite-dimensional Ornstein–Uhlenbeck process
This section and the subsequent Sections 5 and 6 are devoted to proving the abstract discrete-to-continuum approximation results which are applied to the Whittle–Matérn graph discretization setting in Section 3. Thus, from this point onwards we no longer necessarily work with graphs or Whittle–Matérn operators. Instead, we have the following abstract setting.
Let the filtered probability space $(\Omega,\mathcal{F},(\mathcal{F}_{t})_{t\in[0,T]},\mathbb{P})$ be given. For any $n\in\overline{\mathbb{N}}$, we consider the following linear stochastic evolution equation, whose state space is a real and separable UMD-type-2 Banach space $(E_{n},\|\,\cdot\,\|_{E_{n}})$:
[…]
Here, $A_{n}\colon\mathsf{D}(A_{n})\subseteq E_{n}\to E_{n}$ is a linear operator and $T\in(0,\infty)$ is a time horizon. Moreover, we take $W_{n}\coloneqq\Pi_{n}W_{\infty}$, where $(W_{\infty}(t))_{t\geq 0}$ denotes a cylindrical Wiener process on $(\Omega,\mathcal{F},(\mathcal{F}_{t})_{t\in[0,T]},\mathbb{P})$ taking values in a separable Hilbert space $(H_{\infty},\langle\,\cdot\,,\,\cdot\,\rangle_{H_{\infty}})$ and the operator $\Pi_{n}\in\mathscr{L}(H_{\infty};H_{n})$ is as in assumption (A1) below. Thus, the formal time derivative $\dot{W}_{\infty}$ of $W_{\infty}$ represents space–time Gaussian white noise and $(W_{n})_{t\geq 0}$ is an $H_{n}$-valued $Q_{n}$-cylindrical Wiener process colored in space by the covariance operator $Q_{n}\coloneqq\Pi_{n}\Pi_{n}^{*}\in\mathscr{L}(H_{n})$.
We impose the following uniformity assumptions on the spaces $(E_{n})_{n\in\overline{\mathbb{N}}}$, $(H_{n})_{n\in\overline{\mathbb{N}}}$ and the operators $(A_{n})_{n\in\overline{\mathbb{N}}}$:
  1. (A1)
    Assumption 2.1 holds for the UMD-type-2 Banach spaces $(E_{n})_{n\in\overline{\mathbb{N}}}$ and the Hilbert spaces $(H_{n})_{n\in\overline{\mathbb{N}}}$, with $\widetilde{E}\coloneqq E_{\infty}$ and $\widetilde{H}\coloneqq H_{\infty}$, both with the same sequence $(\Lambda_{n},\Pi_{n})_{n\in\overline{\mathbb{N}}}$ of lifting and projection operators.
  2. (A2)
    The operators $(A_{n})_{n\in\overline{\mathbb{N}}}$ on $(E_{n})_{n\in\overline{\mathbb{N}}}$ are uniformly sectorial of angle less than $\frac{1}{2}\pi$ (see Appendix C for the definition of this concept). In particular, their negatives generate bounded analytic $C_{0}$-semigroups $(S_{n}(t))_{t\geq 0}\subseteq\mathscr{L}(E_{n})$ which satisfy Assumption 2.2 with $w=0$. Moreover, there exists a $\beta\in[0,\frac{1}{2})$ such that $R_{n}^{\beta}\coloneqq(\operatorname{Id}_{n}+A_{n})^{-\beta}\in\gamma(H_{n};E_{n})$ for every $n\in\overline{\mathbb{N}}$.
In general, the fractional powers of the sectorial operators $A_{n}$ appearing in (A2) can be defined using any of the equivalent definitions in [41, Chapter 3]. If, as in Section 3, the operator $A_{n}$ given as the restriction of an operator whose eigenvalues form an orthonormal basis on some Hilbert space containing $E_{n}$, then one can use the spectral definition (3.13) of fractional powers of $A_{n}$.
The solution concept which we consider for all of the equations in this work is that of a mild solution. For the linear equation (4.1), it is given by the following stochastic convolution:
Proposition 4.1.
Let $n\in\overline{\mathbb{N}}$ and $T\in(0,\infty)$. Under Assumptions (A1)(A2), the stochastic convolution
$W_{A_{n}}(t)\coloneqq\int_{0}^{t}S_{n}(t-s)\mathop{}\!\mathrm{d}W_{n}(s),\quad t\in[0,T],$ (4.2)
is a well-defined process in $C([0,T];L^{p}(\Omega;E_{n}))$ for every $p\in[1,\infty)$.
Proof.
For every $p\in[1,\infty)$ and $t\in[0,T]$, we have by the Itô inequality (2.7):
$\|W_{A_{n}}(t)\|_{L^{p}(\Omega;E_{n})}^{2}\lesssim_{(p,E_{n})}\int_{0}^{t}\|S_{n}(t-s)\|_{\gamma(H_{n};E_{n})}^{2}\mathop{}\!\mathrm{d}s\leq\|S_{n}\|_{L^{2}(0,T;\gamma(H_{n};E_{n}))}^{2}$ (4.3)
To thow that the right-hand side is bounded, we use the ideal property (2.4) of $\gamma(H_{n};E_{n})$ and the estimate (C.3) for analytic semigroups (in conjuction with Assumptions (A1) and (A2)) to see that
$\|S_{n}\|_{L^{2}(0,T;\gamma(H_{n};E_{n}))}^{2}$ $=\int_{0}^{T}\|(\operatorname{Id}_{n}+A_{n})^{\beta}S_{n}(t)R_{n}^{\beta}\|_{\gamma(H_{n};E_{n})}^{2}\mathop{}\!\mathrm{d}t$ (4.4)
$\leq\|R_{n}^{\beta}\|_{\gamma(H_{n};E_{n})}^{2}\int_{0}^{T}\|(\operatorname{Id}_{n}+A_{n})^{\beta}S_{n}(t)\|_{\mathscr{L}(E_{n})}^{2}\mathop{}\!\mathrm{d}t$
$\lesssim_{\beta}\|R_{n}^{\beta}\|_{\gamma(H_{n};E_{n})}^{2}\int_{0}^{T}t^{-2\beta}\mathop{}\!\mathrm{d}t=\|R_{n}^{\beta}\|_{\gamma(H_{n};E_{n})}^{2}\frac{T^{1-2\beta}}{1-2\beta}<\infty.$
Note that $\|R_{n}^{\beta}\|_{\gamma(H_{n};E_{n})}^{2}$ is finite by (A2). Next, applying the Itô inequality (2.7) to the difference $W_{A_{n}}(t+h)-W_{A_{n}}(t)$ for small enough $h\in\mathbb{R}$ yields
$\|W_{A_{n}}(t+h)-W_{A_{n}}(t)\|_{L^{p}(\Omega;E_{n})}\lesssim_{(p,E_{n})}\|S_{n}(\,\cdot\,+h)-S_{n}\|_{L^{2}(0,T;\gamma(H_{n};E_{n}))}\to 0,$ (4.5)
as $h\to 0$, by the strong continuity of translation operators on the Bochner space $L^{2}(0,T;\gamma(H_{n};E_{n}))$. This shows $W_{A_{n}}\in L^{p}(\Omega;C([0,T];E_{n}))$. ∎
Two stochastic processes $(X(t))_{t\in[0,T]}$ and $(Y(t))_{t\in[0,T]}$ are said to be modifications of each other if $\mathbb{P}(X(t)=Y(t))=1$ for all $t\in[0,T]$.
Definition 4.2.
An $E_{n}$-valued stochastic process $X_{n}=(X_{n}(t))_{t\in[0,T]}$ belonging to $C([0,T];L^{p}(\Omega;E_{n}))$ for some $p\in[1,\infty)$ is said to be a mild solution to (4.1) if it is a modification of the process $W_{A_{n}}$ defined in (4.2).
The existence and uniqueness (up to modification) of the mild solution to (4.1) in $C([0,T];L^{p}(\Omega;E_{n}))$ is then immediate from Definition 4.2.
As remarked in Section 2.2, we mainly have applications in mind where the problem corresponding to $n=\infty$ is interpreted as a spatiotemporal stochastic partial differential equation. In the present linear setting, its solution is also known as an infinite-dimensional Ornstein–Uhlenbeck process, and solutions to (4.1) for $n\in\mathbb{N}$ are spatially discretized approximations. Therefore, it is natural to ask whether we can identify the right mode of convergence of the operators $(A_{n})_{n\in\mathbb{N}}$ to $A_{\infty}$ as $n\to\infty$ to ensure the convergence of the processes $(W_{A_{n}})_{n\in\mathbb{N}}$ to $W_{A_{\infty}}$.
The answer is provided by Proposition 4.4 below, which is a stochastic counterpart of the discrete-to-continuum Trotter–Kato approximation theorem for strongly continuous semigroups recalled in Theorem 2.3. In fact, with an eye towards the proof of Proposition 4.5 below, we consider a more general class of auxiliary processes, see equation (4.8).
Before stating any discrete-to-continuum results, let us introduce some convenient notation for this goal. Using the operators $(\Lambda_{n})_{n\in\mathbb{N}}$ and $(\Pi_{n})_{n\in\mathbb{N}}$, we can take a mapping which has $E_{n}$ as its domain or state space, and turn it into an analogous mapping from or to $E_{\infty}$. For instance, we define the $E_{\infty}$-valued processes $\widetilde{W}_{A_{n}}\coloneqq\Lambda_{n}W_{A_{n}}$, as well as the operators $\widetilde{R}_{n}^{\alpha}\coloneqq\Lambda_{n}R_{n}^{\alpha}\Pi_{n}$ and $\widetilde{S}_{n}(t)\coloneqq\Lambda_{n}S_{n}(t)\Pi_{n}$ in $\mathscr{L}(E_{\infty})$ for $\alpha,t\in[0,\infty)$. Now we can formulate our notion of the convergence $A_{n}\to A_{\infty}$ as $n\to\infty$ as follows:
  1. (A3)
    For every $x\in E_{\infty}$, we have $\widetilde{R}_{n}^{1}x\to R_{\infty}^{1}x$ in $E_{\infty}$ as $n\to\infty$. Moreover, given $\beta\in[0,\frac{1}{2})$ as in (A2), we have $\widetilde{R}_{n}^{\beta}\to R_{\infty}^{\beta}$ in $\gamma(H_{\infty};E_{\infty})$.
Proposition 4.4 states that this type of convergence of the operators is sufficient to ensure convergence of the solutions to the linear stochastic evolution equation (4.1). Its proof is based on Lemma C.1 in Appendix C, as well as the following general approximation lemma for square-integrable functions with values in the space of $\gamma$-radonifying operators. It is a simpler analog to [55, Lemma 2.6], which was only necessary to allow for stochastic equations in UMD Banach spaces without type $2$.
Lemma 4.3.
Let $(E,\|\,\cdot\,\|_{E})$ and $(F,\|\,\cdot\,\|_{F})$ be Banach spaces, and let $(H,\langle\,\cdot\,,\,\cdot\,\rangle_{H})$ be a Hilbert space. Given $a,b\in\mathbb{R}$ with $a<b$, let $M_{n}\colon(a,b)\to\mathscr{L}(E;F)$ for all $n\in\overline{\mathbb{N}}$ and suppose that
  1. (i)
    $M_{n}\otimes x\to M\otimes x$ uniformly on compact subsets of $(a,b)$ for all $x\in E$, and
  2. (ii)
    $\sup_{n\in\overline{\mathbb{N}}}\sup_{t\in(a,b)}\|M_{n}(t)\|_{\mathscr{L}(E;F)}<\infty$.
For all $R\in L^{2}(a,b;\gamma(H;E))$ and $n\in\overline{\mathbb{N}}$, we have $M_{n}\otimes R\in L^{2}(a,b;\gamma(H;F))$ and
$M_{n}R\to MR\quad\text{in }L^{2}(a,b;\gamma(H;F))\quad\text{as }n\to\infty.$ (4.6)
Proof.
Arguing as in the proof of Proposition B.3 and using (ii), it follows that we only need to prove the claim for all $R$ belonging to some dense subset $D$ of $L^{2}(a,b;\gamma(H;E))$. Note that every $R\in L^{2}(a,b;\gamma(H;E))$ can be approximated by a step function $\sum_{j=1}^{N}\mathbf{1}_{(a_{j},b_{j}^{\prime})}\otimes T_{j}$ with $a<a_{j}^{\prime}<b_{j}^{\prime}<b$ and $T_{j}\in\gamma(H;E)$, and by definition of $\gamma(H;E)$ the latter can be approximated by finite-rank operators. By linearity, it thus suffices to prove the statement for $R$ of the form
$R(t)=\mathbf{1}_{(a^{\prime},b^{\prime})}(t)\,h\otimes x,\quad\text{where }a<a^{\prime}<b^{\prime}<b\text{ and }(h,x)\in H\times E.$
Substituting this representation, using (2.5) and (i), we find as $n\to\infty$:
$\|M_{n}R-MR\|_{L^{2}(a,b;\gamma(H;F))}^{2}$ $=\int_{a^{\prime}}^{b^{\prime}}\|h\otimes[M_{n}(t)x-M(t)x]\|_{\gamma(H;F)}^{2}\mathop{}\!\mathrm{d}t$ (4.7)
$=\|h\|_{H}^{2}\int_{a^{\prime}}^{b^{\prime}}\|M_{n}(t)x-M(t)x\|_{F}^{2}\mathop{}\!\mathrm{d}t$
$\leq\|h\|_{H}^{2}(b^{\prime}-a^{\prime})\sup_{t\in(a^{\prime},b^{\prime})}\|M_{n}(t)x-M(t)x\|_{F}^{2}\to 0.\qed$
Proposition 4.4.
Suppose that Assumptions (A1) and (A2) hold. Let us define the auxiliary processes
$W^{\delta}_{A_{n}}(t)\coloneqq\frac{1}{\Gamma(\delta)}\int_{0}^{t}(t-s)^{\delta-1}S_{n}(t-s)\mathop{}\!\mathrm{d}W_{n}(s),\quad\delta\in(\nicefrac{{1}}{{2}},\infty),\;t\in[0,\infty),$ (4.8)
where $\Gamma$ denotes the Gamma function [67, Section 5.2]. Then, for every $\beta^{\prime}\in(\beta,\infty)$, $T\in(0,\infty)$ and $p\in[1,\infty)$, we have $W_{A_{n}}^{\beta^{\prime}+\frac{1}{2}}\in C([0,T];L^{p}(\Omega;E_{n}))$ for all $n\in\overline{\mathbb{N}}$. If we suppose in addition that Assumption (A3) is satisfied, then
$\widetilde{W}_{A_{n}}^{\beta^{\prime}+\frac{1}{2}}\to W_{A_{\infty}}^{\beta^{\prime}+\frac{1}{2}}\quad\text{in }C([0,T];L^{p}(\Omega;E_{\infty}))\quad\text{as }n\to\infty.$ (4.9)
Proof.
The fact that $W_{A_{n}}^{\beta^{\prime}+\frac{1}{2}}\in C([0,T];L^{p}(\Omega;E_{n}))$ for all $n\in\overline{\mathbb{N}}$ can be established by arguing as in Proposition 4.1, thus using Assumptions (A1) and (A2). For all $t\in[0,T]$, the Itô inequality (2.7) yields
$\|\widetilde{W}_{A_{n}}^{\beta^{\prime}+\frac{1}{2}}(t)-W_{A_{\infty}}^{\beta^{\prime}+\frac{1}{2}}(t)\|_{L^{p}(\Omega;E_{\infty})}$ (4.10)
$\quad\lesssim_{(p,E_{\infty})}\frac{1}{\Gamma(\beta^{\prime}+\frac{1}{2})}\biggl{(}\int_{0}^{t}(t-s)^{2\beta^{\prime}-1}\|\widetilde{S}_{n}(t-s)-S_{\infty}(t-s)\|_{\gamma(H_{\infty};E_{\infty})}^{2}\mathop{}\!\mathrm{d}s\biggr{)}^{\frac{1}{2}}$ (4.11)
$\quad\leq\frac{1}{\Gamma(\beta^{\prime}+\frac{1}{2})}\biggl{(}\int_{0}^{T}s^{2\beta^{\prime}-1}\|\widetilde{S}_{n}(s)-S_{\infty}(s)\|_{\gamma(H_{\infty};E_{\infty})}^{2}\mathop{}\!\mathrm{d}s\biggr{)}^{\frac{1}{2}}.$ (4.12)
Since semigroups commute with fractional powers of their infinitesimal generators, we can write the difference between the semigroups as follows:
$\widetilde{S}_{n}(s)-S_{\infty}(s)$ $=\Lambda_{n}(\operatorname{Id}_{n}+A_{n})^{\beta}S_{n}(s)\Pi_{n}\widetilde{R}_{n}^{\beta}-(\operatorname{Id}_{\infty}+A_{\infty})^{\beta}S_{\infty}(s)R_{\infty}^{\beta}$ (4.13)
[…]
Thus, by the triangle inequality, it suffices to show that
(I) $\coloneqq\int_{0}^{T}s^{2\beta^{\prime}-1}\|\Lambda_{n}(\operatorname{Id}_{n}+A_{n})^{\beta}S_{n}(s)\Pi_{n}(\widetilde{R}_{n}^{\beta}-R_{\infty}^{\beta})\|_{\gamma(H_{\infty};E_{\infty})}^{2}\mathop{}\!\mathrm{d}s\quad\text{ and }$ (4.15)
(II) $\coloneqq\int_{0}^{T}s^{2\beta^{\prime}-1}\|[\Lambda_{n}(\operatorname{Id}_{n}+A_{n})^{\beta}S_{n}(s)\Pi_{n}-(\operatorname{Id}_{\infty}+R_{\infty})^{\beta}S_{\infty}]R_{\infty}^{\beta}\|_{\gamma(H_{\infty};E_{\infty})}^{2}\mathop{}\!\mathrm{d}s$
tend to zero as $n\to\infty$. Applying the ideal property (2.4) of $\gamma(H_{\infty};E_{\infty})$, followed by the analytic semigroup estimate (C.3) in conjunction with Assumptions (A1) and (A2), we find
(I) $\leq\|\widetilde{R}_{n}^{\beta}-R_{\infty}^{\beta}\|_{\gamma(H_{\infty};E_{\infty})}^{2}\int_{0}^{T}s^{2\beta^{\prime}-1}\|\Lambda_{n}(\operatorname{Id}_{n}+A_{n})^{\beta}S_{n}(s)\Pi_{n}\|_{\mathscr{L}(E_{\infty})}^{2}\mathop{}\!\mathrm{d}s$ (4.16)
$\lesssim_{(\beta,M_{\Lambda},M_{\Pi})}\|\widetilde{R}_{n}^{\beta}-R_{\infty}^{\beta}\|_{\gamma(H_{\infty};E_{\infty})}^{2}\int_{0}^{T}s^{2(\beta^{\prime}-\beta)-1}\mathop{}\!\mathrm{d}s$
$=\frac{T^{2(\beta^{\prime}-\beta)}}{2(\beta^{\prime}-\beta)}\|\widetilde{R}_{n}^{\beta}-R_{\infty}^{\beta}\|_{\gamma(H_{\infty};E_{\infty})}^{2}\to 0,$
where the convergence on the last line follows from the second part of (A3). The convergence $\mathrm{(II)}\to 0$ follows by applying Lemma 4.3 with
$M_{n}(t)\coloneqq t^{\beta}\Lambda_{n}(\operatorname{Id}_{n}+A_{n})^{\beta}S_{n}(t)\Pi_{n}\quad\text{and}\quad R(t)\coloneqq t^{\beta^{\prime}-\beta-\frac{1}{2}}R_{\infty}^{\beta},$ (4.17)
Indeed, this is justified since $R\in L^{2}(0,T;\gamma(H_{\infty};E_{\infty}))$ with
$\|R\|_{L^{2}(0,T;\gamma(H_{\infty};E_{\infty}))}^{2}=\frac{T^{2(\beta^{\prime}-\beta)}}{2(\beta^{\prime}-\beta)}\|R^{\beta}_{\infty}\|_{\gamma(H_{\infty};E_{\infty})}^{2}<\infty,$ (4.18)
condition (ii) is verified by applying (C.3) to $\|M_{n}(t)\|_{\mathscr{L}(E_{\infty})}$ combined with Assumptions (A1) and (A2) as in (4.16), and hypothesis (i) holds by Lemma C.1. ∎
We will show that there exist modifications of $W_{A_{\infty}}$ and $(\widetilde{W}_{A_{n}})_{n\in\mathbb{N}}$ which, for all $p\in[1,\infty)$ and $T\in(0,\infty)$, belong to $L^{p}(\Omega;C([0,T];E_{\infty}))$ and converge in this norm. In particular, as $n\to\infty$, their trajectories converge uniformly on bounded time intervals, $\mathbb{P}$-a.s.
The proof is based on the Da Prato–Kwapień–Zabczyk factorization method, first formulated in [19] for Hilbert spaces (see also [20, Section 5.3]), and later extended to UMD-type-2 Banach spaces in [9, Theorem 3.2]. The general idea is to express the process $W_{A_{n}}$ as the ‘product’ $\mathfrak{I}_{A_{n}}^{\frac{1}{2}-\beta^{\prime}}W^{\frac{1}{2}+\beta^{\prime}}_{A_{n}}$ of a fractional parabolic integral operator $\mathfrak{I}_{A_{n}}^{\frac{1}{2}-\beta^{\prime}}$ as in Appendix B and auxiliary process $W^{\frac{1}{2}+\beta^{\prime}}_{A_{n}}$ as in (4.8), and using the smoothing properties of the former.
Proposition 4.5.
Let $p\in[1,\infty)$ and $T\in(0,\infty)$. If Assumptions (A1)(A2) hold, then for every $n\in\overline{\mathbb{N}}$ there exists a modification of $W_{A_{n}}$ which belongs to $L^{p}(\Omega;C([0,T];E_{n})$, which we will identify with $W_{A_{n}}$ itself.
If, in addition, Assumption (A3) holds, then the sequence $(\widetilde{W}_{A_{n}})_{n\in\overline{\mathbb{N}}}$ satisfies
$\widetilde{W}_{A_{n}}\to W_{A_{\infty}}\quad\text{in }L^{p}(\Omega;C([0,T];E_{\infty}))\quad\text{as }n\to\infty.$ (4.19)
Proof.
Let $\beta^{\prime}\in(\beta,\frac{1}{2})$, where $\beta\in[0,\frac{1}{2})$ is as in (A2). Since $L^{p}(\Omega)$-spaces with higher exponents are embedded in those with lower ones contractively (because $\mathbb{P}(\Omega)=1$), we assume without loss of generality that $p\in((\frac{1}{2}-\beta)^{-1},\infty)$. By the first part of Proposition 4.4 (thus using Assumptions (A1) and (A2)) and the Fubini theorem, we have for all $n\in\overline{\mathbb{N}}$:
$W^{\frac{1}{2}+\beta^{\prime}}_{A_{n}}\in C([0,T];L^{p}(\Omega;E_{n}))\hookrightarrow L^{p}(0,T;L^{p}(\Omega;E_{n}))\cong L^{p}(\Omega;L^{p}(0,T;E_{n})),$ (4.20)
where the constants for the first embedding depend only on $p$ and $T$. In particular, there exists an event $\Omega_{1}\subseteq\Omega$ with $\mathbb{P}(\Omega_{1})=1$ such that $W^{\frac{1}{2}+\beta^{\prime}}_{A_{n}}(\omega,\,\cdot\,)$ belongs to $L^{p}(0,T;E_{n})$ for all $\omega\in\Omega_{1}$. It then follows from Proposition B.1b that $\mathfrak{I}_{A_{n}}^{\frac{1}{2}-\beta^{\prime}}W^{\frac{1}{2}+\beta^{\prime}}_{A_{n}}(\omega,\,\cdot\,)$ belongs to $C([0,T];E_{n})$, where $(\mathfrak{I}_{A_{n}}^{s})_{s\in[0,\infty)}$ are the fractional parabolic integral operators defined by (B.2) in Appendix B. In this case, the process $\mathfrak{I}_{A_{n}}^{\frac{1}{2}-\beta^{\prime}}W^{\frac{1}{2}+\beta^{\prime}}_{A_{n}}$ (set to zero outside of $\Omega_{1}$) belongs to $L^{p}(\Omega;C([0,T];E_{n}))$, and by the factorization theorem [9, Theorem 3.2] it is a modification of $W_{A_{n}}$.
For the lifted processes, the properties of the embeddings and projections from assumption (A1) imply that ~$\widetilde{\mathfrak{I}}_{A_{n}}^{\frac{1}{2}-\beta^{\prime}}\widetilde{W}_{A_{n}}^{\frac{1}{2}+\beta^{\prime}}\in L^{p}(\Omega;C([0,T];E_{\infty}))$ is a continuous modification of $\widetilde{W}_{A_{n}}$, where ~$\widetilde{\mathfrak{I}}_{A_{n}}^{s}\coloneqq\Lambda_{n}\mathfrak{I}_{A_{n}}^{s}\Pi_{n}$. Identifying $\widetilde{W}_{A_{n}}$ with its factorized continuous modification for every $n\in\overline{\mathbb{N}}$, we can estimate as follows:
$\|\widetilde{W}_{A_{n}}-W_{A_{\infty}}\|_{L^{p}(\Omega;C([0,T];E_{\infty}))}$ =~$=\|\widetilde{\mathfrak{I}}_{A_{n}}^{\frac{1}{2}-\beta^{\prime}}\widetilde{W}_{A_{n}}^{\frac{1}{2}+\beta^{\prime}}-{\mathfrak{I}}_{A_{\infty}}^{\frac{1}{2}-\beta^{\prime}}W_{A_{\infty}}^{\frac{1}{2}+\beta^{\prime}}\|_{L^{p}(\Omega;C([0,T];E_{\infty}))}$ (4.21)
~$\leq\|\widetilde{\mathfrak{I}}_{A_{n}}^{\frac{1}{2}-\beta^{\prime}}(\widetilde{W}_{A_{n}}^{\frac{1}{2}+\beta^{\prime}}-W_{A_{\infty}}^{\frac{1}{2}+\beta^{\prime}})\|_{L^{p}(\Omega;C([0,T];E_{\infty}))}$
+(~$\quad+\|(\widetilde{\mathfrak{I}}_{A_{n}}^{\frac{1}{2}-\beta^{\prime}}-\mathfrak{I}_{A_{\infty}}^{\frac{1}{2}-\beta^{\prime}})W_{A_{\infty}}^{\frac{1}{2}+\beta^{\prime}}\|_{L^{p}(\Omega;C([0,T];E_{\infty}))}.$
Since $\frac{1}{2}-\beta^{\prime}>\frac{1}{p}$, we can apply Corollary B.2b to find that ~$\widetilde{\mathfrak{I}}_{A_{n}}^{\frac{1}{2}-\beta^{\prime}}$ is a bounded linear operator from $L^{p}(0,T;E_{\infty})$ to $C([0,T];E_{\infty})$ whose norm can be bounded independently of $n$. Thus, by the above discussion and the second part of Proposition 4.4 (which uses Assumption (A3)), we find
~$\|\widetilde{\mathfrak{I}}_{A_{n}}^{\frac{1}{2}-\beta^{\prime}}(\widetilde{W}_{A_{n}}^{\frac{1}{2}+\beta^{\prime}}-W_{A_{\infty}}^{\frac{1}{2}+\beta^{\prime}})\|_{L^{p}(\Omega;C([0,T];E_{\infty}))}$ (4.22)
$\lesssim\|\widetilde{W}_{A_{n}}^{\frac{1}{2}+\beta^{\prime}}-W_{A_{\infty}}^{\frac{1}{2}+\beta^{\prime}}\|_{L^{p}(\Omega\times(0,T);E_{\infty})}\leq T^{\frac{1}{p}}\|\widetilde{W}_{A_{n}}^{\frac{1}{2}+\beta^{\prime}}-W_{A_{\infty}}^{\frac{1}{2}+\beta^{\prime}}\|_{C([0,T];L^{p}(\Omega;E_{\infty}))}\to 0.$
Now we note that, for all $\omega\in\Omega$, Proposition B.3a implies that
~$\|\widetilde{\mathfrak{I}}_{A_{n}}^{\frac{1}{2}-\beta^{\prime}}W_{A_{\infty}}^{\frac{1}{2}+\beta^{\prime}}(\omega,\,\cdot\,)-\mathfrak{I}_{A_{\infty}}^{\frac{1}{2}-\kappa}W_{A_{\infty}}^{\frac{1}{2}+\beta^{\prime}}(\omega,\,\cdot\,)\|_{C([0,T];E_{\infty})}\to 0.$ (4.23)
Again by Corollary B.2b, we moreover have
~$\|\widetilde{\mathfrak{I}}_{A_{n}}^{\frac{1}{2}-\beta^{\prime}}W_{A_{\infty}}^{\frac{1}{2}+\beta^{\prime}}(\omega,\,\cdot\,)-\mathfrak{I}_{A_{\infty}}^{\frac{1}{2}-\beta^{\prime}}W_{A_{\infty}}^{\frac{1}{2}+\beta^{\prime}}(\omega,\,\cdot\,)\|_{C([0,T];E_{\infty})}\lesssim 2\|W_{A_{\infty}}^{\frac{1}{2}+\beta^{\prime}}\|_{L^{p}(0,T;E_{\infty})}$ (4.24)
with constant independent of $n\in\mathbb{N}$, and since $W_{A_{\infty}}^{\frac{1}{2}+\beta^{\prime}}\in L^{p}((0,T)\times\Omega;E_{\infty})$, the dominated convergence theorem yields
(~$\|(\widetilde{\mathfrak{I}}_{A_{n}}^{\frac{1}{2}-\beta^{\prime}}-\mathfrak{I}_{A_{\infty}}^{\frac{1}{2}-\beta^{\prime}})W_{A_{\infty}}^{\frac{1}{2}+\beta^{\prime}}\|_{L^{p}(\Omega;C([0,T];E_{\infty}))}\to 0.\qed$
5. Approximation of semilinear stochastic evolution equations with additive cylindrical Wiener noise
In this section, we shall extend the results from Section 4 regarding the linear $E_{n}$-valued equation (4.1) to the semilinear case. As before, let the spaces $(E_{n})_{n\in\overline{\mathbb{N}}}$, $(H_{n})_{n\in\overline{\mathbb{N}}}$ and the operators $(A_{n})_{n\in\overline{\mathbb{N}}}$ satisfy assumptions (A1) and (A2), respectively, and suppose that $W_{n}\coloneqq\Pi_{n}W_{\infty}$ is an $H_{n}$-valued $Q_{n}$-cylindrical Wiener process (with $Q_{n}\coloneqq\Pi_{n}\Pi_{n}^{*}$), supported on $(\Omega,\mathcal{F},(\mathcal{F}_{t})_{t\in[0,T]},\mathbb{P})$. Let $T\in(0,\infty)$ be a finite time horizon. In this section, we suppose moreover that we are given a drift coefficient function $F_{n}\colon\Omega\times[0,T]\times E_{n}\to E_{n}$ and initial datum $\xi_{n}\colon\Omega\to E_{n}$. We will consider the following semilinear stochastic evolution equation:
[…]
Note that $F_{n}(\omega,t,\,\cdot\,)$ is a (nonlinear) operator on $E_{n}$ for all $(\omega,t)\in\Omega\times[0,T]$; in Section 3, we considered the specific case $[F_{n}(\omega,t,u_{n})](x)\coloneqq f_{n}(\omega,t,u_{n}(x))$ for some real-valued nonlinearity $f_{n}$.
In what follows, we shall impose more precise conditions on the $F_{n}$ and $\xi_{n}$ to ensure the well-posedness of (5.1) for every fixed $n\in\overline{\mathbb{N}}$ and to obtain discrete-to-continuum convergence of the respective solutions as $n\to\infty$.
In Section 5.1 we will assume in particular that the drifts $(F_{n})_{n\in\overline{\mathbb{N}}}$ are uniformly globally Lipschitz and of linear growth to obtain the existence of unique global solutions $(X_{n}(t))_{t\in[0,T]}$, whose lifted counterparts $\widetilde{X}_{n}$ converge to $X_{\infty}$ in $L^{p}(\Omega;C([0,T];E_{\infty}))$ as $n\to\infty$, where $p\in[1,\infty)$ is the stochastic integrability of the initial data. These assumptions are relaxed in Section 5.2, where we suppose that the drifts are uniformly locally Lipschitz and uniformly bounded near zero. In general, this comes at the cost of obtaining merely local solutions, converging in a weaker norm. However, if one can show independently that the solutions are global and the $L^{p}(\Omega;C([0,T];E_{n}))$-norms of $X_{n}$ are uniformly bounded in $n\in\mathbb{N}$, then we recover the stronger sense of convergence.
5.1. Globally Lipschitz drifts of linear growth
In this section we suppose that the drift coefficients $F_{n}\colon\Omega\times[0,T]\times E_{n}\to E_{n}$ in (5.1) for $n\in\overline{\mathbb{N}}$ are uniformly globally Lipschitz continuous and of linear growth. More precisely:
  1. (F1)
    There exist $L_{F},C_{F}\in(0,\infty)$ such that, for all $t\in[0,T]$, $\omega\in\Omega$, $n\in\overline{\mathbb{N}}$ and $x_{n},y_{n}\in E_{n}$,
    $\|F_{n}(\omega,t,x_{n})-F_{n}(\omega,t,y_{n})\|_{E_{n}}$ $\leq L_{F}\|x_{n}-y_{n}\|_{E_{n}};$ (5.2)
    $\|F_{n}(\omega,t,x_{n})\|_{E_{n}}$ $\leq C_{F}(1+\|x_{n}\|_{E_{n}}).$ (5.3)
    Moreover, the process $(\omega,t)\mapsto F_{n}(\omega,t,x_{n})$ is strongly measurable and adapted to the filtration $(\mathcal{F}_{t})_{t\in[0,T]}$.
Now let us comment on the existence and uniqueness of solutions to (5.1) for fixed $n\in\overline{\mathbb{N}}$. We will use the following concept of global mild solutions, see [83, pp. 969–970]. In Subsection 5.2, we also introduce the concept of local solutions, which may blow up in finite time. In particular, a local solution which exists $\mathbb{P}$-a.s. on the whole of $[0,T]$ is in fact global.
Recall that $(S_{n}(t))_{t\geq 0}$ denotes the $C_{0}$-semigroup on $E_{n}$ generated by $-A_{n}$.
Definition 5.1.
An $E_{n}$-valued stochastic process $X_{n}=(X_{n}(t))_{t\in[0,T]}$ is a global mild solution to (5.1) with coefficients $(A_{n},F_{n},\xi_{n})$ if
  1. (i)
    $X_{n}\colon\Omega\times[0,T]\to E_{n}$ is strongly measurable and $(\mathcal{F}_{t})_{t\in[0,T]}$-adapted;
  2. (ii)
    $s\mapsto S_{n}(t-s)F_{n}(s,X_{n}(s))\in L^{0}(\Omega;L^{1}(0,t;E_{n}))$ for every $t\in[0,T]$;
  3. (iii)
    $s\mapsto S_{n}(t-s)\in L^{2}(0,t;\gamma(H_{n};E_{n}))$ for every $t\in[0,T]$;
  4. (iv)
    for all $t\in[0,T]$, we have
[…]
In the present framework, existence and uniqueness can be proven by showing that the operator $\Phi_{n,T}$ given by
$[\Phi_{n,T}(u_{n})](t)\coloneqq S_{n}(t)\xi_{n}+\int_{0}^{t}S_{n}(t-s)F_{n}(s,u_{n}(s))\mathop{}\!\mathrm{d}s+W_{A_{n}}(t)$ (5.5)
is a well-defined and Lipschitz-continuous mapping on $L^{p}(\Omega;C([0,T];E_{n}))$, whose Lipschitz constant tends to zero as $T\downarrow 0$ (see [83, Proposition 6.1] or [55, Theorem 3.7] for more general results):
Proposition 5.2.
Suppose that Assumptions (A1)(A2) and (F1) are satisfied. Let $n\in\overline{\mathbb{N}}$, $p\in[1,\infty)$, $\xi_{n}\in L^{p}(\Omega,\mathcal{F}_{0},\mathbb{P};E_{n})$ and $T\in(0,\infty)$. The operator $\Phi_{n,T}$ given by (5.5) is well defined and Lipschitz continuous on $L^{p}(\Omega;C([0,T];E_{n}))$. Its Lipschitz constant is independent of $\xi_{n}$, depends on $A_{n}$ and $F_{n}$ only through $M_{S}$ and $L_{F}$, and tends to zero as $T\downarrow 0$.
Proof.
The fact that $S_{n}\otimes\xi_{n}\in L^{p}(\Omega;C([0,T];E_{n}))$ is immediate from (A1)(A2) and $\xi_{n}\in L^{p}(\Omega,\mathcal{F}_{0},\mathbb{P};E_{n})$. We also have $W_{A_{n}}\in L^{p}(\Omega;C([0,T];E_{n}))$ by the first part of Proposition 4.5. Given $u_{n}\in L^{p}(\Omega;C([0,T];E_{n}))$, it follows from (F1) that $\|s\mapsto F_{n}(s,u_{n}(s))\|_{L^{\infty}(0,T;E_{n})}\leq C_{F}(1+\|u_{n}\|_{C([0,T];E_{n})})$, so that $S_{n}*F_{n}(\,\cdot\,,u_{n})$ belongs to $L^{p}(\Omega;C([0,T];E_{n}))$ with
$\|S_{n}*F_{n}(\,\cdot\,,u_{n})\|_{L^{p}(\Omega;C([0,T];E_{n}))}\leq C_{F}(1+\|u_{n}\|_{L^{p}(\Omega;C([0,T];E_{n}))})$ (5.6)
by Proposition B.1b with $E=F=E_{n}$, $\alpha=0$ and $s=1$ (noting that $S_{n}*f=\mathfrak{I}_{A_{n}}^{1}f$, see Appendix B). This shows that $\Phi_{n,T}$ is well-defined.
Now let $u_{n},v_{n}\in L^{p}(\Omega;C([0,T];E_{n}))$ and observe that
$\Phi_{n,T}(u_{n})-\Phi_{n,T}(v_{n})=\int_{0}^{t}S_{n}(t-s)[F_{n}(s,u_{n}(s))-F_{n}(s,v_{n}(s))]\mathop{}\!\mathrm{d}s.$ (5.7)
A straightforward estimate involving Assumptions (A1)(A2) and (F1) then yields
$\|\Phi_{n,T}(u_{n})-\Phi_{n,T}(v_{n})\|_{L^{p}(\Omega;C([0,T];E_{n}))}\leq M_{S}L_{F}T\,\|u_{n}-v_{n}\|_{L^{p}(\Omega;C([0,T];E_{n}))}.\qed$
Under the conditions of Proposition 5.2, it follows from the Banach fixed-point theorem that (5.1) has a unique solution on a small enough time interval $[0,T_{0}]$, which can be extended to a unique global mild solution on any $[0,T]$ by “patching together” solutions on small time intervals:
Proposition 5.3.
Suppose that Assumptions (A1)(A2) and (F1) are satisfied, and let $n\in\overline{\mathbb{N}}$, $p\in[1,\infty)$, $\xi_{n}\in L^{p}(\Omega,\mathcal{F}_{0},\mathbb{P};E_{n})$ and $T\in(0,\infty)$. Then (5.1) has a unique global mild solution $X_{n}\in L^{p}(\Omega;C([0,T];E_{n}))$.
Proof.
By Proposition 5.2, there exists $T_{0}\in(0,\infty)$ such that $\Phi_{n,T_{0}}$ is a strict contraction on $L^{p}(\Omega;C([0,T_{0}];E_{n}))$, and thus has a unique fixed point $X_{n}$. Since the bound on the Lipschitz constant of $\Phi_{n,T}$ only depended on $M_{S}$, $L_{F}$ and $T$, we can repeat the previous argument to obtain a unique solution $Y\in L^{p}(\Omega;C([0,T_{0}];E_{n}))$ for (5.1) with initial datum $\eta_{n}\coloneqq X_{n}(\frac{1}{2}T_{0})$, drift $G_{n}(\,\cdot\,,u_{n})\coloneqq F_{n}(\,\cdot\,+\frac{1}{2}T_{0},u_{n})$ and noise $\widehat{W}_{n}\coloneqq W_{n}(\,\cdot\,+\frac{1}{2}T_{0})$. It can then be argued directly using Definition 5.1 that the concatenation of the processes $X_{n}$ and $Y_{n}(\,\cdot\,+\frac{1}{2}T_{0})$ is the unique mild solution to (5.1) with the original data on $[0,\frac{3}{2}T_{0}]$. Proceeding inductively, we find the same conclusion for all intervals $[0,(k+\frac{1}{2})T_{0}]$ with $k\in\mathbb{N}$ and thus for $[0,T]$. ∎
For every $n\in\mathbb{N}$, we analogously define the lifted initial datum $\widetilde{\xi}_{n}\colon\Omega\to E_{\infty}$ by $\widetilde{\xi}_{n}\coloneqq\Lambda_{n}\xi_{n}$ and the lifted drift coefficient $\widetilde{F}_{n}\colon\Omega\times[0,T]\times E_{\infty}\to E_{\infty}$ by
$\widetilde{F}_{n}(\omega,t,x)\coloneqq\Lambda_{n}F_{n}(\omega,t,\Pi_{n}x),\quad(\omega,t,x)\in\Omega\times[0,T]\times E_{\infty}.$ (5.8)
We will assume that the initial data and drift coefficients are approximated in the following way:
  1. (IC)
    There exists $p\in[1,\infty)$ such that $(\xi_{n})_{n\in\overline{\mathbb{N}}}\in\prod_{n\in\overline{\mathbb{N}}}L^{p}(\Omega,\mathcal{F}_{0},\mathbb{P};E_{n})$ and
    $\widetilde{\xi}_{n}\to\xi_{\infty}\quad\text{in }L^{p}(\Omega;E_{\infty})\text{ as }n\to\infty.$ (5.9)
  1. (F2)
    For a.e. $(\omega,t)\in\Omega\times[0,T]$ and every $x\in E_{\infty}$, we have
    $\widetilde{F}_{n}(\omega,t,x)\to F_{\infty}(\omega,t,x)\quad\text{ in }E_{\infty}\text{ as }n\to\infty.$ (5.10)
Under these assumptions, we obtain the main result of this section, namely the following discrete-to-continuum convergence theorem in the context of uniformly globally Lipschitz nonlinearities of linear growth. It is analogous to [55, Theorem 4.3].
Theorem 5.4.
Suppose that (A1)(A3)(F1)(F2) and (IC) are satisfied, with $p\in[1,\infty)$. For all $n\in\overline{\mathbb{N}}$ and $T\in(0,\infty)$, let $X_{n}=(X_{n}(t))_{t\in[0,T]}$ denote the unique global solution to (5.1), and let $\widetilde{X}_{n}\coloneqq\Lambda_{n}X_{n}$. Then we have
$\widetilde{X}_{n}\to X_{\infty}\quad\text{in }L^{p}(\Omega;C([0,T];E_{\infty}))\quad\text{as }n\to\infty.$ (5.11)
Its proof involves the lifted counterparts of $\Phi_{n,T}$, defined by
$\widetilde{\Phi}_{n,T}\coloneqq\Lambda_{n}\Phi_{n,T}\circ\Pi_{n}\colon L^{p}(\Omega;C([0,T];E_{\infty}))\to L^{p}(\Omega;C([0,T];E_{\infty})),$ (5.12)
i.e., for all $u\in L^{p}(\Omega;C([0,T];E_{\infty}))$ and $t\in[0,T]$, we have
$[\widetilde{\Phi}_{n,T}(u)](t)\coloneqq{}$ $\Lambda_{n}S_{n}(t)\xi_{n}+\int_{0}^{t}\Lambda_{n}S_{n}(t-s)F_{n}(s,\Pi_{n}u(s))\mathop{}\!\mathrm{d}s+\Lambda_{n}W_{A_{n}}(t)$ (5.13)
[…]
where the second identity is due to Assumption 2.1(iii). Using the tensor and convolution notations from Section 2.1, it can be expressed even more concisely as
$\widetilde{\Phi}_{n,T}(u)=\widetilde{S}_{n}\otimes\widetilde{\xi}_{n}+\widetilde{S}_{n}*\widetilde{F}_{n}(\,\cdot\,,u)+\widetilde{W}_{A_{n}}$ (5.15)
In particular, we will show that all three terms of (5.15) converge to their “continuum” counterparts; they are addressed by Lemmas 5.55.6 below (which are analogous to [55, Lemma 4.4, 4.5(1) and 4.5(3)]), as well as Proposition 4.5.
Lemma 5.5.
If (A1)(A3) and (IC) hold with $p\in[1,\infty)$, then we have
$\widetilde{S}_{n}\otimes\xi_{n}\to S_{\infty}\otimes\xi_{\infty}\quad\text{in }L^{p}(\Omega;C([0,T];E_{\infty}))\quad\text{as }n\to\infty.$ (5.16)
Proof.
As in the beginning of the proof of Proposition 5.2, it follows from (A1)(A2) and (IC) that $S_{n}\otimes\xi_{n}\in L^{p}(\Omega;C([0,T];E_{n}))$ for all $n\in\overline{\mathbb{N}}$. Applying the projection and lifting operators from (A1), we thus find $\widetilde{S}_{n}\otimes\widetilde{\xi}_{n}\in L^{p}(\Omega;C([0,T];E_{\infty}))$.
The triangle inequality implies
$\|\widetilde{S}_{n}\otimes\widetilde{\xi}_{n}-S_{\infty}\otimes\xi_{\infty}\|_{L^{p}(\Omega;C([0,T];E_{\infty}))}$ (5.17)
$\quad\leq\|\widetilde{S}_{n}\otimes(\widetilde{\xi}_{n}-\xi_{\infty})\|_{L^{p}(\Omega;C([0,T];E_{\infty}))}+\|(\widetilde{S}_{n}-S_{\infty})\otimes\xi_{\infty}\|_{L^{p}(\Omega;C([0,T];E_{\infty}))}.$ (5.18)
By (A1)(A2) and (IC), for the first term we have, as $n\to\infty$:
$\|\widetilde{S}_{n}\otimes(\widetilde{\xi}_{n}-\xi_{\infty})\|_{L^{p}(\Omega;C([0,T];E_{\infty}))}\leq M_{\Lambda}M_{S}M_{\Pi}\|\widetilde{\xi}_{n}-\xi_{\infty}\|_{L^{p}(\Omega;E_{\infty})}\to 0$ (5.19)
For the second term, first note that $\widetilde{S}_{n}\otimes\xi_{\infty}(\omega)\to S_{\infty}\otimes\xi_{\infty}(\omega)$ in $C([0,T];E_{\infty})$, $\mathbb{P}$-a.s., by Theorem 2.3, where we now also use (A3). Since we moreover have
$\|\widetilde{S}_{n}\otimes\xi_{\infty}(\omega)-S_{\infty}\otimes\xi_{\infty}(\omega)\|_{C([0,T];E_{\infty})}\leq M_{S}(M_{\Lambda}M_{\Pi}+1)\|\xi_{\infty}(\omega)\|_{E_{\infty}},$ (5.20)
and the right-hand side belongs to $L^{p}(\Omega)$ by assumption, we deduce that also
$\|(\widetilde{S}_{n}-S_{\infty})\otimes\xi_{\infty}\|_{L^{p}(\Omega;C([0,T];E_{\infty}))}\to 0\quad\text{as }n\to\infty.\qed$
Lemma 5.6.
Suppose that Assumptions (A1)(A2)(F1) and (F2) are satisfied. Let $p\in[1,\infty)$ and $u\in L^{p}(\Omega;C([0,T];E_{\infty}))$ be given. Then we have
$\widetilde{S}_{n}*\widetilde{F}_{n}(\,\cdot\,,u)\to S_{\infty}*F_{\infty}(\,\cdot\,,u)\quad\text{in }L^{p}(\Omega;C([0,T];E_{\infty}))\text{ as }n\to\infty.$ (5.21)
Proof.
Similarly to the proof of Proposition 5.2, it follows from (A1)(A2) and (F1) that $\widetilde{S}_{n}*\widetilde{F}_{n}(\,\cdot\,,u)\in L^{p}(\Omega;C([0,T];E_{\infty}))$ for all $n\in\overline{\mathbb{N}}$. By the triangle inequality, we can split up the statement into the following two assertions:
  1. (i)
    $\widetilde{S}_{n}*\widetilde{F}_{n}(\,\cdot\,,u)-\widetilde{S}_{n}*F_{\infty}(\,\cdot\,,u)\to 0$ in $L^{p}(\Omega;C([0,T];E_{\infty}))$ as $n\to\infty$;
  2. (ii)
    $\widetilde{S}_{n}*F_{\infty}(\,\cdot\,,u)\to S_{\infty}*F_{\infty}(\,\cdot\,,u)$ in $L^{p}(\Omega;C([0,T];E_{\infty}))$ as $n\to\infty$.
For almost every $(\omega,t)\in\Omega\times[0,T]$, we have by (F1) and (A1):
$\|\widetilde{F}_{n}(\omega,t,u(\omega,t))-F_{\infty}(\omega,t,u(\omega,t))\|_{E_{\infty}}$ (5.22)
$\quad\leq C_{F}(M_{\Lambda}+1+(M_{\Pi}M_{\Lambda}+1)\|u(\omega,t)\|_{E_{\infty}}).$
It follows that
$\|\widetilde{F}_{n}(\,\cdot\,,u)-F_{\infty}(\,\cdot\,,u)\|_{L^{p}(\Omega\times(0,T);E_{\infty})}$ $\lesssim_{(C_{F},M_{\Lambda},M_{\Pi})}\|u\|_{L^{p}(\Omega\times(0,T);E_{\infty})}$ (5.23)
$\lesssim_{(p,T)}\|u\|_{L^{p}(\Omega;C([0,T];E_{\infty}))}<\infty.$
Since $\widetilde{S}_{n}*f=\widetilde{\mathfrak{I}}_{A_{n}}^{1}f$ for all $f\in L^{p}(0,T;E_{\infty})$, we can apply Proposition B.1b with $E=F=E_{\infty}$, $\alpha=0$ and $s=1$ to find that
$\|\widetilde{S}_{n}*(\widetilde{F}_{n}(\,\cdot\,,u)-F_{\infty}(\,\cdot\,,u))\|_{C([0,T];E_{\infty})}$ (5.24)
$\quad\lesssim_{(s,r,T,M_{S})}\|\widetilde{F}_{n}(\,\cdot\,,u)-F_{\infty}(\,\cdot\,,u)\|_{L^{p}(\Omega\times(0,T);E_{\infty})}.$
The latter tends to zero as $n\to\infty$ by the dominated convergence theorem, which applies in view of (F2) and (5.22)–(5.23). This shows (i).
For (ii), we derive in the same way that, for almost every $\omega\in\Omega$,
$t\mapsto F_{\infty}(\omega,t,u(\omega,t))\in L^{p}(0,T;E_{\infty}),$ (5.25)
which implies, cf. Proposition B.3a with $\widetilde{E}\coloneqq E_{\infty}$, that
$\widetilde{S}_{n}*F_{\infty}(\omega,\,\cdot\,,u(\omega,\,\cdot\,))\to S_{\infty}*F_{\infty}(\omega,\,\cdot\,,u(\omega,\,\cdot\,))\quad\text{in }C([0,T];E_{\infty})\quad\text{as }n\to\infty.$ (5.26)
The conclusion follows by using the uniform boundedness of the operators (~$(\widetilde{\mathfrak{I}}_{A_{n}}^{1})_{n\in\overline{\mathbb{N}}}$ in $\mathscr{L}(L^{p}(0,T;E_{\infty});C([0,T];E_{\infty}))$, asserted in Corollary B.2b (with $\widetilde{E}\coloneqq E_{\infty}$ once more), and finishing the dominated convergence argument as in part (i). ∎
Proof of Theorem 5.4.
By Proposition 5.2 and (A1), for small enough $T_{0}\in(0,\infty)$ there exists a constant $c\in[0,1)$, depending only on $L_{F}$, $M_{S}$, $M_{\Lambda}$ and $M_{\Pi}$, such that, for all $u,v\in L^{p}(\Omega;C([0,T_{0}];E_{\infty}))$,
$\sup\nolimits_{n\in\mathbb{N}}\,\|\widetilde{\Phi}_{n,T_{0}}(u)-\widetilde{\Phi}_{n,T_{0}}(v)\|_{L^{p}(\Omega;C([0,T_{0}];E_{\infty}))}\leq c\|u-v\|_{L^{p}(\Omega;C([0,T_{0}];E_{\infty}))}.$ (5.27)
Moreover, by Proposition 5.3, for every $n\in\overline{\mathbb{N}}$, there exists a unique global solution $X_{n}\in L^{p}(\Omega;C([0,T];E_{n}))$ to (5.1), which in particular satisfies $X_{n}=\Phi_{n,T_{0}}(X_{n})$ when restricted to $[0,T_{0}]$. By (A1), this implies $\widetilde{X}_{n}=\widetilde{\Phi}_{n,T_{0}}(\widetilde{X}_{n})$. Hence,
$\|X_{\infty}-\widetilde{X}_{n}\|_{L^{p}(\Omega;C([0,T_{0}];E_{\infty}))}=\|\Phi_{\infty,T_{0}}(X_{\infty})-\widetilde{\Phi}_{n,T_{0}}(\widetilde{X}_{n})\|_{L^{p}(\Omega;C([0,T_{0}];E_{\infty}))}$ (5.28)
$\quad\leq\|\Phi_{\infty,T_{0}}(X_{\infty})-\widetilde{\Phi}_{n,T_{0}}(X_{\infty})\|_{L^{p}(\Omega;C([0,T_{0}];E_{\infty}))}$ (5.29)
$\qquad\quad+\|\widetilde{\Phi}_{n,T_{0}}(X_{\infty})-\widetilde{\Phi}_{n,T_{0}}(\widetilde{X}_{n})\|_{L^{p}(\Omega;C([0,T_{0}];E_{\infty}))}$ (5.30)
$\quad\leq\|\Phi_{\infty,T_{0}}(X_{\infty})-\widetilde{\Phi}_{n,T_{0}}(X_{\infty})\|_{L^{p}(\Omega;C([0,T_{0}];E_{\infty}))}$ (5.31)
$\qquad\quad+c\|X_{\infty}-\widetilde{X}_{n}\|_{L^{p}(\Omega;C([0,T_{0}];E_{\infty}))},$ (5.32)
so that Lemmas 5.55.6 and Proposition 4.5 yield (using all of the Assumptions (A1)(A3)(F1)(F2) and (IC)):
$\|X_{\infty}-\widetilde{X}_{n}\|_{L^{p}(\Omega;C([0,T_{0}];E_{\infty}))}$ (5.33)
$\quad\leq\frac{1}{1-c}\,\|\Phi_{\infty,T_{0}}(X_{\infty})-\widetilde{\Phi}_{n,T_{0}}(X_{\infty})\|_{L^{p}(\Omega;C([0,T_{0}];E_{\infty}))}\to 0$
as $n\to\infty$. In order to extend the convergence to arbitrary time horizons, we write
$\|X_{\infty}-\widetilde{X}_{n}\|_{L^{p}(\Omega;C([0,\frac{3}{2}T_{0}];E_{\infty}))}$ $\leq\|X_{\infty}-\widetilde{X}_{n}\|_{L^{p}(\Omega;C([0,\frac{1}{2}T_{0}];E_{\infty}))}$ (5.34)
$\quad+\|X_{\infty}-\widetilde{X}_{n}\|_{L^{p}(\Omega;C([\frac{1}{2}T_{0},\frac{3}{2}T_{0}];E_{\infty}))}.$ (5.35)
The first term tends to zero as $n\to\infty$ by (5.33). As for the second term, we note that $X_{n}|_{[\frac{1}{2}T_{0},\frac{3}{2}T_{0}]}$ are the respective solutions to (5.1) with shifted drift functions $(F_{n}(\,\cdot\,+\frac{1}{2}T_{0},\,\cdot\,))_{n\in\overline{\mathbb{N}}}$ and initial values $(X_{n}(\frac{1}{2}T_{0}))_{n\in\overline{\mathbb{N}}}$. Since the Lipschitz constants of the fixed point operators defined above did not depend on the initial datum and only depended on $F$ through its (time-independent) Lipschitz constant $L_{F}$, we can repeat the same argument to find that the second term tends to zero. Proceeding by induction, we obtain the convergence $\widetilde{X}_{n}\to X_{\infty}$ in $L^{p}(\Omega;C([0,(1+\frac{k}{2})T_{0}];E_{\infty}))$ for any $k\in\mathbb{N}$, and thus in $L^{p}(\Omega;C([0,T];E_{\infty}))$ for any $T\in(0,\infty)$. ∎
5.2. Locally Lipschitz nonlinearities
In this section, we work under the weaker assumption that the drift coefficients $(F_{n})_{n\in\overline{\mathbb{N}}}$ are locally Lipschitz continuous, with local-Lipschitz constants uniformly bounded in $t\in[0,T]$, $\omega\in\Omega$ and $n\in\overline{\mathbb{N}}$. Moreover, we replace the uniform linear growth condition from (F1) by the assumption that $F_{n}(t,\omega,0)$ is bounded, again uniformly in $t$, $\omega$ and $n$; we will also call this notion of boundedness local since it only involves $u=0$. Thus, we assume that the $(F_{n})_{n\in\overline{\mathbb{N}}}$ are locally uniformly Lipschitz and locally uniformly bounded:
  1. (F1)
    For every $r\in(0,\infty)$ there exists a constant $L_{F}^{(r)}\in(0,\infty)$ such that for almost every $(\omega,t)\in\Omega\times[0,T]$, all $n\in\overline{\mathbb{N}}$ and every $x_{n},y_{n}\in E_{n}$ such that $\|x_{n}\|_{E_{n}},\|y_{n}\|_{E_{n}}\leq r$, we have
    $\|F_{n}(\omega,t,x_{n})-F_{n}(\omega,t,y_{n})\|_{E_{n}}\leq L_{F}^{(r)}\|x_{n}-y_{n}\|_{E_{n}}.$ (5.36)
    Moreover, for every $x_{n}\in E_{n}$, $n\in\overline{\mathbb{N}}$ the process $(\omega,t)\mapsto F_{n}(\omega,t,x_{n})$ is strongly measurable and adapted, and there exists a constant $C_{F,0}$ such that
    $\|F_{n}(\omega,t,0)\|_{E_{n}}\leq C_{F,0}\quad\text{for all }n\in\overline{\mathbb{N}}.$ (5.37)
Under these conditions, we can in general not expect to obtain global solutions of (5.1) in the sense of Definition 5.1. Instead, we need to work with locally defined $E_{n}$-valued stochastic processes, i.e., with mappings of the form
$Y\colon\{(\omega,t)\in\Omega\times[0,T]:t\in[0,\tau(\omega))\}\to E_{n}$ (5.38)
for some stopping time $\tau\colon\Omega\to[0,T]$. We denote such a process by $Y=(Y(t))_{t\in[0,\tau)}$. If the half-open interval $[0,\tau(\omega))$ in (5.38) is replaced by $[0,\tau(\omega)]$, then we write $Y=(Y(t))_{t\in[0,\tau]}$ instead. We say that $(Y(t))_{t\in[0,\tau)}$ is admissible if
  • for all $t\in[0,T]$, the mapping $\{\omega\in\Omega:t<\tau(\omega)\}\ni\omega\mapsto Y(\omega,t)\in E_{n}$ is $\mathcal{F}_{t}$-measurable;
  • the mapping $[0,\tau(\omega))\ni t\mapsto Y(\omega,t)\in E_{n}$ is continuous, $\mathbb{P}$-a.s.
We denote by $V^{\mathrm{loc}}([0,\tau)\times\Omega;E_{n})$ the space of admissible $E_{n}$-valued processes $(Y(t))_{t\in[0,\tau)}$ for which there exists a sequence $(\tau_{m})_{m\in\mathbb{N}}$ of stopping times such that, for $\mathbb{P}$-a.e. $\omega\in\Omega$, we have $\tau_{m}(\omega)\uparrow\tau(\omega)$ as $m\to\infty$ and $\|Y\|_{C([0,\tau_{m}(\omega)];E_{n})}<\infty$ for all $m\in\mathbb{N}$. As in [83, Section 8], we define local solutions to (5.1) as follows:
Definition 5.7.
An admissible $E_{n}$-valued stochastic process $X_{n}=(X_{n}(t))_{t\in[0,\tau)}$ is said to be a local solution to (5.1) with coefficients $(A_{n},F_{n},\xi_{n})$ if there exists a sequence $(\tau_{m})_{m\in\mathbb{N}}$ of stopping times such that $\tau_{m}\uparrow\tau$ as $m\to\infty$, $\mathbb{P}$-a.s., and for all $m\in\mathbb{N}$ we have
  1. (i)
    for every $t\in[0,T]$, the process $(\omega,s)\mapsto S_{n}(t-s)F_{n}(\omega,s,X_{n}(\omega,s))\mathbf{1}_{[0,\tau_{m}]}(s)$ belongs to $L^{0}(\Omega;L^{1}(0,t;E_{n}))$;
  2. (ii)
    for every $t\in[0,T]$, $s\mapsto S_{n}(t-s)\mathbf{1}_{[0,\tau_{m}]}(s)\in L^{2}(0,t;\gamma(H_{n};E_{n}))$ ;
  3. (iii)
    it holds $\mathbb{P}$-a.s. that for all $t\in[0,\tau_{m}]$, we have
    $X_{n}(t)=S_{n}(t)\xi_{n}$ $+\int_{0}^{t}S_{n}(t-s)F_{n}(s,X_{n}(s))\mathbf{1}_{[0,\tau_{m}]}(s)\mathop{}\!\mathrm{d}s$ (5.39)
    $+\int_{0}^{t}S_{n}(t-s)\mathbf{1}_{[0,\tau_{m}]}(s)\mathop{}\!\mathrm{d}W_{n}(s).$ (5.40)
We say that a local solution $(X_{n}(t))_{t\in[0,\tau)}$ to (5.1) is maximal if for any other local solution $(\overline{X}_{n}(t))_{t\in[0,\overline{\tau})}$ it holds $\mathbb{P}$-a.s. that $\overline{\tau}\leq\tau$ and $X_{n}|_{[0,\overline{\tau})}\equiv\overline{X}_{n}$. It is called global if $\tau=T$ holds $\mathbb{P}$-a.s. and there exists a solution $(\widehat{X}_{n}(t))_{t\in[0,T]}$ to (5.1) in the sense of Definition 5.1 such that $\widehat{X}_{n}|_{[0,\tau)}\equiv X_{n}$, $\mathbb{P}$-a.s. The stopping time $\tau$ is called an explosion time if
$\limsup\nolimits_{t\uparrow\tau(\omega)}\|X_{n}(\omega,t)\|_{E_{n}}=\infty\quad\text{for a.e.\ }\omega\in\Omega\text{ such that }\tau(\omega)<T.$ (5.41)
The following local well-posedness result then follows from [83, Theorem 8.1]:
Theorem 5.8 ([83, Theorem 8.1]).
Suppose that Assumptions (A1)(A2) and (F1) are satisfied, and let $n\in\overline{\mathbb{N}}$, $p\in[1,\infty)$, $\xi_{n}\in L^{p}(\Omega,\mathcal{F}_{0},\mathbb{P};E_{n})$ and $T\in(0,\infty)$ be given. Then (5.1) has a unique maximal local mild solution $(X_{n}(t))_{t\in[0,\sigma_{n})}$ in $V^{\mathrm{loc}}([0,\sigma_{n})\times\Omega;E_{n})$, where $\sigma_{n}\colon\Omega\to[0,T]$ is an explosion time.
Combined with the convergence assumptions (IC) and (F2), we can argue analogously to [56, Theorem 3.3 and Corollary 3.4] to derive the following extension of Theorem 5.4 to the present setting.
Theorem 5.9.
Suppose that Assumptions (A1)(A2)(IC)(F1) and (F2) are satisfied. For $n\in\overline{\mathbb{N}}$, let $(X_{n}(t))_{t\in[0,\sigma_{n})}$ be the maximal local solution to (5.1) with explosion time $\sigma_{n}\colon\Omega\to[0,T]$, and set $\widetilde{X}_{n}\coloneqq\Lambda_{n}X_{n}$. Then the following is true:
  1. (i)
    We have $\widetilde{X}_{n}\mathbf{1}_{[0,\sigma_{\infty}\wedge\sigma_{n})}\to X_{\infty}\mathbf{1}_{[0,\sigma_{\infty})}$ in $L^{0}(\Omega\times[0,T];E_{\infty})$ as $n\to\infty$.
If, moreover, $\sigma_{n}=T$ holds $\mathbb{P}$-a.s. for all $n\in\mathbb{N}$ and $p\in[1,\infty)$ is such that
$\sup_{n\in\mathbb{N}}\|X_{n}\|_{L^{p}(\Omega;C([0,T];E_{n}))}<\infty,$ (5.42)
then the following assertions also hold:
  1. (ii)
    We have $\sigma_{\infty}=T$, $\mathbb{P}$-a.s.
  2. (iii)
    If $p\in(1,\infty)$, then for all $p^{-}\in[1,p)$ we have
    $\widetilde{X}_{n}\to X_{\infty}\quad\text{ in }L^{p^{-}}(\Omega;C([0,T];E_{\infty}))\text{ as }n\to\infty.$ (5.43)
Similarly to [56, Theorem 3.3 and Corollary 3.4], the proof of Theorem 5.9 relies on the following general approximation results for locally defined processes:
Theorem 5.10 ([56, Theorem 2.1 and Corollary 2.5]).
Let $(E,\|\,\cdot\,\|_{E})$ be a real and separable Banach space and $T\in(0,\infty)$. For every $n\in\overline{\mathbb{N}}$, suppose that $(Y_{n}(t))_{t\in[0,\sigma_{n})}$ is a continuous and adapted $E$-valued locally defined process with explosion time $\sigma_{n}\colon\Omega\to(0,T]$, and define the stopping times $\rho_{n}^{(r)}\colon\Omega\to[0,T]$ by
$\rho_{n}^{(r)}\coloneqq\inf\{t\in(0,\sigma_{n}):\|Y_{n}(t)\|_{E}>r\},\quad r\in(0,\infty),$ (5.44)
with the convention that $\inf\emptyset\coloneqq T$. Moreover, suppose that for each $r\in(0,\infty)$ there exists a (globally defined) continuous and adapted $E$-valued process $(Y_{n}^{(r)}(t))_{t\in[0,T]}$ which satisfies the following two conditions:
  1. (a)
    For all $n\in\overline{\mathbb{N}}$ and $r\in(0,\infty)$, it holds $\mathbb{P}$-a.s. that
    $Y_{n}^{(r)}\mathbf{1}_{[0,\rho_{n}^{(r)}]}\equiv Y_{n}\mathbf{1}_{[0,\rho_{n}^{(r)}]}\quad\text{ on }[0,T];$ (5.45)
  2. (b)
    For all $r\in(0,\infty)$ we have
    $Y_{n}^{(r)}\to Y_{\infty}^{(r)}\quad\text{ in }L^{0}(\Omega;C([0,T];E))\text{ as }n\to\infty.$ (5.46)
Then the following assertions hold:
  1. (i)
    For all $r\in(0,\infty)$ and $\varepsilon>0$ it holds $\mathbb{P}$-a.s. that
    $\liminf_{n\to\infty}\rho_{n}^{(r)}\leq\rho_{\infty}^{(r)}\leq\limsup_{n\to\infty}\rho_{n}^{(r+\varepsilon)}.$ (5.47)
  2. (ii)
    For all $r\in(0,\infty)$ and $\varepsilon>0$, we have
    $Y_{n}\mathbf{1}_{[0,\rho_{\infty}^{(r)}\wedge\rho_{n}^{(r+\varepsilon)})}\to Y_{\infty}\mathbf{1}_{[0,\rho_{\infty}^{(r)})}\quad\text{ in }L^{0}(\Omega;B_{\mathrm{b}}([0,T];E))\text{ as }n\to\infty,$ (5.48)
    where $B_{\mathrm{b}}([0,T];E)$ denotes the space of bounded and strongly measurable functions from $[0,T]$ to $E$.
  3. (iii)
    We have
    $Y_{n}\mathbf{1}_{[0,\sigma_{\infty}\wedge\sigma_{n})}\to Y_{\infty}\mathbf{1}_{[0,\sigma_{\infty})}\quad\text{ in }L^{0}(\Omega\times[0,T];E)\text{ as }n\to\infty.$ (5.49)
If, in addition, we have $\mathbb{P}(\sigma_{n}=T)=1$ for all $n\in\mathbb{N}$, and $p\in[1,\infty)$ is such that
$\sup_{n\in\mathbb{N}}\|Y_{n}\|_{L^{p}(\Omega;C([0,T];E))}<\infty,$ (5.50)
then the following assertions also hold:
  1. (iv)
    We have $\mathbb{P}(\sigma_{\infty}=T)=1$ and $X_{\infty}\in L^{q}(\Omega;C([0,T];E))$.
  2. (v)
    If $p\in(1,\infty)$, then for all $p^{-}\in[1,p)$ we have
    $Y_{n}\to Y_{\infty}\quad\text{in }L^{p^{-}}(\Omega;C([0,T];E))\text{ as }n\to\infty.$ (5.51)
Proof of Theorem 5.9.
For every $n\in\overline{\mathbb{N}}$ and $r\in(0,\infty)$, let us define the mapping $F_{n}^{(r)}\colon\Omega\times[0,T]\times E_{n}\to E_{n}$ by
$F_{n}^{(r)}(\omega,t,x_{n})\coloneqq\begin{cases}F_{n}(\omega,t,x_{n}),&\text{if }\|\Lambda_{n}x_{n}\|_{E_{\infty}}\leq r,\\ F_{n}\Bigl{(}\omega,t,\frac{rx_{n}}{\|\Lambda_{n}x_{n}\|_{E_{\infty}}}\Bigr{)},&\text{otherwise}.\end{cases}$ (5.52)
For any fixed $r>0$, the sequence $(F_{n}^{(r)})_{n\in\overline{\mathbb{N}}}$ satisfies conditions (F1) and (F2). Indeed, to establish the former, we first note that $F_{n}^{(r)}$ can be written as
$F_{n}^{(r)}(\omega,t,x_{n})=F_{n}(\omega,t,\Pi_{n}R_{r}(\Lambda_{n}x_{n})),$ (5.53)
where $R_{r}\colon E_{\infty}\to E_{\infty}$ denotes the canonical retraction of $E_{\infty}$ onto the closed ball around $0\in E_{\infty}$ with radius $r$:
$R_{r}(x)\coloneqq\begin{cases}x,&\text{if }\|x\|_{E_{\infty}}\leq r;\\ \frac{rx}{\|x\|_{E_{\infty}}},&\text{otherwise}.\end{cases}$ (5.54)
An elementary estimate shows that $R_{r}$ is Lipschitz with constant $2$. It follows that $(F_{n}^{(r)})_{n\in\overline{\mathbb{N}}}$ is uniformly globally Lipschitz with constant $L_{F^{(r)}}\leq L_{F}^{(rM_{\Pi})}M_{\Pi}M_{\Lambda}$, and thus of linear growth with uniform constant $C_{F^{(r)}}\leq\max\{L_{F^{(r)}},C_{F,0}\}$, so that it satisfies (F1).
In order to show that $(F_{n}^{(r)})_{n\in\overline{\mathbb{N}}}$ satisfies (F2), first note that for every sequence $(y_{n})_{n\in\mathbb{N}}\subseteq E_{\infty}$ converging to some $y$ in $E_{\infty}$, we have $\widetilde{F}_{n}(\omega,t,y_{n})\to F_{\infty}(\omega,t,y)$. Indeed, by the triangle inequality it suffices to note that
$\|\widetilde{F}_{n}(\omega,t,y_{n})-\widetilde{F}_{n}(\omega,t,y)\|_{E_{\infty}}\to 0\quad\text{and}\quad\|\widetilde{F}_{n}(\omega,t,y)-F_{\infty}(\omega,t,y)\|_{E_{\infty}}\to 0$ (5.55)
as $n\to\infty$, respectively because $(\widetilde{F}_{n})_{n\in\overline{\mathbb{N}}}$ is uniformly locally Lipschitz (with constants $L_{\widetilde{F}}^{(r)}\leq M_{\Pi}M_{\Lambda}L_{F}^{(r)}$) and since (F2) was assumed for $(F_{n})_{n\in\mathbb{N}}$. Writing
$\widetilde{F}_{n}^{(r)}(\omega,t,x)=\widetilde{F}_{n}(\omega,t,R_{r}(\Lambda_{n}\Pi_{n}x)),$ (5.56)
see (5.53), we can apply the above observation to the sequence $y_{n}\coloneqq R_{r}(\Lambda_{n}\Pi_{n}x)$, which converges to $y\coloneqq R_{r}(x)$ in $E_{\infty}$ as $n\to\infty$ in view of Assumption 2.1(ii) and the (Lipschitz) continuity of $R_{r}$. Therefore, we find $\widetilde{F}_{n}^{(r)}(\omega,t,x)\to F^{(r)}_{\infty}(\omega,t,x)$, thus proving the claim that (F2) holds for $(F_{n}^{(r)})_{n\in\overline{\mathbb{N}}}$ as well.
For each $r>0$, condition (F1) for $(F_{n}^{(r)})_{n\in\overline{\mathbb{N}}}$ yields the existence of a unique global solution $(X_{n}^{(r)}(t))_{t\in[0,T]}$ to (5.1) with coefficients $(A_{n},F_{n}^{(r)},\xi_{n})$. In order to establish statement (i) of the present theorem, we will apply the corresponding parts (i)(iii) of Theorem 5.10 to the processes $Y_{n}\coloneqq\widetilde{X}_{n}$; hence we need to verify its conditions (a) and (b) for $(\widetilde{X}_{n})_{n\in\overline{\mathbb{N}}}$. First note that we have $\rho_{n}^{(r)}\leq\sigma_{n}$, where $\rho_{n}^{(r)}$ is defined by (5.44), and that the restrictions of $X_{n}^{(r)}$ and $X_{n}$ to $[0,\rho_{n}^{(r)})$ are local solutions to (5.1) with coefficients $(A_{n},F_{n}^{(r)},\xi_{n})$ and $(A_{n},F_{n},\xi_{n})$, respectively. Since it holds $\mathbb{P}$-a.s. that $\|\Lambda_{n}X_{n}(t)\|_{E_{\infty}}\leq r$ for $t\in[0,\rho_{n}^{(r)})$, we find
[…]
hence $(X_{n}(t))_{t\in[0,\rho_{n}^{(r)})}$ is in fact also a local solution of the equation with coefficients $(A_{n},F_{n}^{(r)},\xi_{n})$. Therefore, the local uniqueness of (5.1) (cf. [83, Lemma 8.2]) implies that $X_{n}^{(r)}\equiv X_{n}$ on $[0,\rho_{n}^{(r)}]$ holds $\mathbb{P}$-a.s., and applying $\Lambda_{n}$ on both sides verifies (a). Condition (b) follows by applying Theorem 5.4 to $(F_{n}^{(r)})_{n\in\overline{\mathbb{N}}}$, proving (i).
Finally, since $(\Lambda_{n})_{n\in\mathbb{N}}$ is uniformly bounded in view of Assumption 2.1(i), we see that (5.42) implies that the conditions of Theorem 5.10(iv) and (v) are satisfied, which directly yields the remaining assertions (ii) and (iii). ∎
6. Reaction–diffusion type equations
In this section, we introduce another family of Banach spaces $(B_{n})_{n\overline{\mathbb{N}}}$ such that each $B_{n}$ embeds into $E_{n}$ and $\widetilde{B}\supseteq B_{\infty}$ (along with other assumptions, given in Subsection 6.1), and consider the $B_{n}$-valued counterparts of (5.1). The main purpose of this setting is to eventually specialize to the class of stochastic reaction–diffusion type equations which are formally given, for any $n\in\overline{\mathbb{N}}$, by
[…]
where $t\in(0,T]$, $T\in(0,\infty)$, $x\in\mathcal{D}_{n}\subseteq\mathbb{R}^{d}$ and $f_{n}\colon\Omega\times[0,T]\times\mathbb{R}\to\mathbb{R}$ is a locally Lipschitz (real-valued) function. This problem amounts to letting the drift $F_{n}$ in (5.1) be a Nemytskii operator (also called a superposition operator), i.e., defining it by
$[F_{n}(\omega,t,u)](x)\coloneqq f_{n}(\omega,t,u(x)),\quad(\omega,t,x)\in\Omega\times[0,T]\times\mathcal{D}_{n},$ (6.2)
for a given function $u\colon\mathcal{D}_{n}\to\mathbb{R}$. However, in order for a Nemytskii operator $F_{n}$ to inherit the local Lipschitz continuity from $f_{n}$, we cannot view it as acting on the UMD-type-2 Banach space $E_{n}=L^{q}(\mathcal{D}_{n})$ with $q\in[1,\infty)$. In fact, by [1, Theorem 3.9], the operator $u\mapsto F_{n}(\omega,t,u)$ defined by (6.2) is weakly continuous on $L^{q}(\mathcal{D}_{n})$ (meaning that it maps weakly convergent sequences to weakly convergent sequences) if and only if it is affine in $u$, i.e., there exist coefficients $a_{n}(\omega,t),b_{n}(\omega,t)\in\mathbb{R}$ such that $[F_{n}(\omega,t,u)](x)=a_{n}(\omega,t)+b_{n}(\omega,t)u(x)$ for all $x\in\mathcal{D}_{n}$. In particular, if $(F_{n})_{n\in\overline{\mathbb{N}}}$ is a family of Nemytskii operators which is uniformly locally Lipschitz in the sense of (F1) on the spaces $E_{n}=L^{q}(\mathcal{D}_{n})$, then it is in fact globally Lipschitz and of linear growth in the sense of (F1).
Thus, we will instead view $F_{n}$ as an operator on $B_{n}\coloneqq C(\overline{\mathcal{D}_{n}})$ (which coincides with $B_{n}=L^{\infty}(\mathcal{D}_{n})$ if $\mathcal{D}_{n}$ is discrete). This, in turn, poses a difficulty for stochastic evolution equations, since there is no theory for the stochastic integration of integrands taking their values in a space of continuous functions. This is due to the poor geometric properties of $(C(\overline{\mathcal{D}_{n}}),\|\,\cdot\,\|_{\infty})$ as a Banach space: The most general notion of stochastic integration in Banach spaces (see [82]) requires at least the UMD property; such spaces are, in particular, reflexive [48, Theorem 4.3.3], which $C(\overline{\mathcal{D}_{n}})$ fails to be.
One way to circumvent this issue is to proceed as in [56, Section 3.2]; namely, defining the fractional domain spaces
$\dot{E}_{n}^{\alpha}\coloneqq\mathsf{D}(A_{n}^{\alpha/2}),\quad\|x_{n}\|_{\dot{E}_{n}^{\alpha}}\coloneqq\|(\operatorname{Id}_{n}+A_{n})^{\alpha/2}x_{n}\|_{E_{n}},$ (6.3)
and supposing that $\dot{E}_{n}^{\theta}\hookrightarrow C(\overline{\mathcal{D}_{n}})\hookrightarrow L^{q}(\mathcal{D}_{n})$ continuously and densely for some $\theta\in[0,1)$, one can carry out the stochastic integration in the space $\dot{E}_{n}^{\theta}$, while working with $C(\overline{\mathcal{D}_{n}})$-valued processes for the fixed-point arguments.
In applications, we typically assume that $A_{n}$ is an (unbounded) linear differential operator on $L^{q}(\mathcal{D}_{n})$, where $q\in[2,\infty)$, augmented with some boundary conditions (“b.c.”) such that $\dot{E}^{\alpha}_{n}$ is the fractional Sobolev space $W^{\alpha,q}_{\text{b.c.}}(\mathcal{D}_{n})$ of order $\alpha$. We then suppose that $\theta$ is chosen large enough in relation to the dimension $d$ that we have the continuous and dense Sobolev embedding $W^{\alpha,q}_{\text{b.c.}}(\mathcal{D}_{n})\hookrightarrow C_{\text{b.c.}}(\overline{\mathcal{D}_{n}})$.
In Section 6.1 we will specify the abstract formulation of the setting outlined above, as well as some additional uniformity conditions with respect to $n\in\overline{\mathbb{N}}$. These will be used to, as a first step, derive $B_{n}$-valued counterparts to the $E_{n}$-valued discrete-to-continuum approximation results for globally Lipschitz drifts of linear growth from Subsection 5.1; in Subsection 6.2 we do the same for the $B_{n}$-valued setting with locally Lipschitz and locally bounded drifts. In the latter case, the solution are local in general, but in Section 6.3 we state an extra dissipativity assumption on $F_{n}$ under which the existence of global solutions to (6.1) has been proven in [56, Section 4]. These processes then also converge in an improved sense, and we can apply this to Section 3.
6.1. Setting and convergence for globally Lipschitz drifts
We start by specifying the abstract setting for the treatment of reaction–diffusion type equations which was outlined at the beginning of this section. That is, we complement the Hilbert spaces $(H_{n})_{n\in\overline{\mathbb{N}}}$ and UMD-type-2 Banach spaces $(E_{n})_{n\in\overline{\mathbb{N}}}$ from Section 5 with a sequence of real separable Banach spaces $(B_{n},\|\,\cdot\,\|_{B_{n}})_{n\in\overline{\mathbb{N}}}$, embedded continuously and densely into $E_{n}$ for each $n\in\overline{\mathbb{N}}$. Moreover, we introduce the real Banach space $(\widetilde{B},\|\,\cdot\,\|_{\widetilde{B}})$, containing $B_{\infty}$ as a closed subspace, and we suppose that all the $B_{n}$ are embedded by into $\widetilde{B}$ by the lifting operators $(\Lambda_{n})_{n\in\mathbb{N}}$ from (A1). The spaces $B_{\infty}\subseteq\widetilde{B}$ should respectively be thought of as $C(\overline{\mathcal{D}})\subseteq L^{\infty}(\mathcal{D})$. More precisely, we will work with the following extensions of assumptions (A1)(A3):
  1. (A1-B)
    Assumption (A1) holds, and Assumption 2.1 is satisfied for $(B_{n},\|\,\cdot\,\|_{B_{n}})_{n\in\overline{\mathbb{N}}}$ and $(\widetilde{B},\|\,\cdot\,\|_{B_{\infty}})$, with the same projection and lifting operators $(\Pi_{n})_{n\in\overline{\mathbb{N}}}$, $(\Lambda_{n})_{n\in\overline{\mathbb{N}}}$ from (A1), for which we set $\widetilde{M}_{\Pi}\coloneqq\sup_{n\in\mathbb{N}}\|\Pi_{n}\|_{\mathscr{L}(\widetilde{B};B_{n})}$ and $\widetilde{M}_{\Lambda}\coloneqq\sup_{n\in\mathbb{N}}\|\Lambda_{n}\|_{\mathscr{L}(B_{n};\widetilde{B})}$.
  2. (A2-B)
    Assumption (A2) holds, the semigroup $(S_{n}(t))_{t\geq 0}\subseteq\mathscr{L}(E_{n})$ leaves $B_{n}$ invariant for all $n\in\overline{\mathbb{N}}$, and its restriction $(S_{n}(t)|_{B_{n}})_{t\geq 0}$ to $B_{n}$ is a strongly continuous semigroup in $\mathscr{L}(B_{n})$. Moreover, there exists an $\widetilde{M}_{S}\in[1,\infty)$ such that
    $\|S_{n}(t)\|_{\mathscr{L}(B_{n})}\leq\widetilde{M}_{S}\quad\text{for all }n\in\overline{\mathbb{N}}\text{ and }t\in[0,\infty).$ (6.4)
  3. (A3-B)
    Assumption (A3) holds, and $\widetilde{R}_{n}x\to R_{\infty}x$ in $\widetilde{B}$ as $n\to\infty$ for all $x\in B_{\infty}$.
By [29, Chapter II, Proposition 2.3], assumptions (A1-B) and (A2-B) imply that the generator of $(S_{n}(t)|_{B_{n}})_{t\geq 0}$ is the operator $-A_{n}|_{B_{n}}\colon\mathsf{D}(A_{n}|_{B_{n}})\subseteq B_{n}\to B_{n}$ defined by
$-A_{n}|_{B_{n}}x_{n}\coloneqq-A_{n}x_{n}\quad\text{on }\mathsf{D}(A_{n}|_{B_{n}})\coloneqq\{x_{n}\in B_{n}\cap\mathsf{D}(A_{n}):A_{n}x_{n}\in B_{n}\},$
which is known as the part of $-A_{n}$ in $B_{n}$. Therefore, by Theorem 2.3, assumption (A3-B) implies $\widetilde{S}_{n}|_{B_{n}}\otimes x\to S|_{B_{\infty}}\otimes x$ in $C([0,T];\widetilde{B})$, as $n\to\infty$, for all $x\in B_{\infty}$ and $T\in(0,\infty)$.
The following uniform ultracontractivity assumption is necessary in order to circumvent the aforementioned problem regarding stochastic integration in arbitrary separable Banach spaces $B_{n}$. It replaces assumption (A3) in [56], which forces all the spaces $(\dot{E}_{n}^{\alpha})_{n\in\overline{\mathbb{N}}}$ to essentially be the same, which is not satisfied in applications such as discrete-to-continuum approximation, where each $\dot{E}_{n}^{\alpha}$ consists of functions defined on a different domain.
  1. (A4-B)
    There exist $\theta\in[0,1)$ and $M_{\theta}\in[0,\infty)$ such that, for all $n\in\overline{\mathbb{N}}$, we have $\dot{E}^{\theta}_{n}\hookrightarrow B_{n}\hookrightarrow E_{n}$ continuously and densely, with
    $\|S_{n}(t)x_{n}\|_{B_{n}}\leq M_{\theta}t^{-\theta/2}\|x_{n}\|_{E_{n}}\quad\text{for all }x_{n}\in E_{n}\text{ and }t\in[0,\infty).$ (6.5)
The uniformity in $n$ of the constant $M_{\theta}$ enables us to prove the following discrete-to-continuum approximation result for the stochastic convolutions $(W_{A_{n}})_{n\in\overline{\mathbb{N}}}$ as $B_{n}$-valued processes (i.e., a $B_{n}$-valued counterpart to Proposition 4.5):
Proposition 6.1.
Let $p\in[1,\infty)$ and $T\in(0,\infty)$ be given, and suppose that Assumptions (A1-B), (A2-B) and (A4-B) hold, with $\theta+2\beta<1$. For every $n\in\overline{\mathbb{N}}$, there exists a modification of $W_{A_{n}}$ which belongs to $L^{p}(\Omega;C([0,T];B_{n}))$, and we identify these modifications with the processes $(W_{A_{n}})_{n\in\overline{\mathbb{N}}}$ themselves.
Under the additional Assumption (A3-B), we have
$\widetilde{W}_{A_{n}}\to W_{A_{\infty}}\quad\text{in }L^{p}(\Omega;C([0,T];\widetilde{B}))\text{ as }n\to\infty.$ (6.6)
Proof.
Fix a $\beta^{\prime}\in(\beta,\frac{1}{2})$ such that $\theta+2\beta^{\prime}<1$, where $\beta\in[0,\frac{1}{2})$ is as in (A2). Without loss of generality, we may assume that $p\in(2,\infty)$ is so large that, in fact, $\theta+2\beta^{\prime}<1-\frac{2}{p}$. As in the proof of Proposition 4.5, we see that $W_{A_{n}}^{\frac{1}{2}+\beta^{\prime}}(\omega,\,\cdot\,)$, where $W_{A_{n}}^{\frac{1}{2}+\beta^{\prime}}$ is the auxiliary process defined by (4.8), belongs to $L^{p}(0,T;E_{n})$ for $\mathbb{P}$-a.e. $\omega\in\Omega$, hence $\mathfrak{I}_{A_{n}}^{\frac{1}{2}-\beta^{\prime}}W_{A_{n}}^{\frac{1}{2}+\beta^{\prime}}(\omega,\,\cdot\,)\in C([0,T];B_{n})$ by applying Proposition B.1b with $E\coloneqq E_{n}$, $F\coloneqq B_{n}$ and $\alpha\coloneqq\theta$ from assumption (A4-B). Moreover, one finds that the process $\mathfrak{I}_{A_{n}}^{\frac{1}{2}-\beta^{\prime}}W_{A_{n}}^{\frac{1}{2}+\beta^{\prime}}$ is a continuous modification of $W_{A_{n}}$, belonging to $L^{p}(\Omega;C([0,T];B_{n}))$. It now suffices to establish the following, as $n\to\infty$:
~$\|\widetilde{\mathfrak{I}}_{A_{n}}^{\frac{1}{2}-\kappa}(\widetilde{W}_{A_{n}}^{\frac{1}{2}+\kappa}-W_{A_{\infty}}^{\frac{1}{2}+\kappa})\|_{L^{p}(\Omega;C([0,T];\widetilde{B}))}$ $\to 0,$ (6.7)
~$\|\widetilde{\mathfrak{I}}_{A_{n}}^{\frac{1}{2}-\kappa}W_{A_{\infty}}^{\frac{1}{2}+\kappa}-\mathfrak{I}_{A_{\infty}}^{\frac{1}{2}-\kappa}W_{A_{\infty}}^{\frac{1}{2}+\kappa}\|_{L^{p}(\Omega;C([0,T];\widetilde{B}))}$ $\to 0.$ (6.8)
These convergences follow by arguing as in the proof of Proposition 4.5, where we now need Corollary B.2c and Proposition B.3b instead of Corollary B.2b and Proposition B.3a, respectively. ∎
For the initial data, we consider the following analog to (IC):
  1. (IC-B)
    There exists $p\in[1,\infty)$ such that $(\xi_{n})_{n\in\overline{\mathbb{N}}}\in\prod_{n\in\overline{\mathbb{N}}}L^{p}(\Omega;B_{n})$ and
    $\widetilde{\xi}_{n}\to\xi_{\infty}\quad\text{in }L^{p}(\Omega;\widetilde{B})\text{ as }n\to\infty.$ (6.9)
Finally, regarding the drift coefficients $(F_{n})_{n\in\overline{\mathbb{N}}}$, we now suppose that
  1. (F1-B)
    Assumption (F1) holds with $(B_{n})_{n\in\overline{\mathbb{N}}}$ in place of $(E_{n})_{n\in\overline{\mathbb{N}}}$, and $B_{n}$-valued Lipschitz and growth constants respectively denoted by $\widetilde{L}_{F}$ and $\widetilde{C}_{F}$.
  2. (F2-B)
    For almost every $(\omega,t)\in\Omega\times[0,T]$ and every $x\in B_{\infty}$ we have
    $\widetilde{F}_{n}(\omega,t,x)\to F_{\infty}(\omega,t,x)\quad\text{ in }\widetilde{B}\text{ as }n\to\infty.$ (6.10)
Note that in the approximation assumptions (A3-B) and (F2-B), we only impose convergence for $x\in B_{\infty}\subseteq\widetilde{B}$, and similarly we only consider $\xi_{\infty}\in L^{p}(\Omega;B_{\infty})$ in (IC-B). Recall that this is sufficient since Theorem 2.3, on which the approximation results ultimately rely, is formulated in this setting.
Under (A1-B), (A2-B), (A4-B) (with $\theta+2\beta<1$), (IC-B) and (F1-B), we can derive well-posedness of $B_{n}$-valued global solutions to (5.1); these are defined by simply replacing the space $E_{n}$ by $B_{n}$ in Definition 5.1. To this end, we again investigate the fixed-point operators $\Phi_{n,T}$ defined by (5.5), viewing them now as acting on $L^{p}(\Omega;C([0,T];B_{n}))$. We have the following analog to Proposition 5.2:
Proposition 6.2.
Suppose that Assumptions (A1-B), (A2-B), (A4-B) hold with $\theta+2\beta<1$, and (F1-B) is satisfied. Let $n\in\overline{\mathbb{N}}$, $p\in[1,\infty)$, $\xi_{n}\in L^{p}(\Omega,\mathcal{F}_{0},\mathbb{P};B_{n})$ and $T\in(0,\infty)$ be given. The operator $\Phi_{n,T}$ given by (5.5) is well defined and Lipschitz continuous on $L^{p}(\Omega;C([0,T];B_{n}))$. Its Lipschitz constant is independent of $\xi_{n}$, depends on $A_{n}$ and $F_{n}$ only through $\widetilde{M}_{S}$ and $\widetilde{L}_{F}$, and tends to zero as $T\downarrow 0$.
Proof.
The fact that $S_{n}\otimes\xi_{n}\in L^{p}(\Omega;C([0,T];B_{n}))$ is immediate from Assumptions (A1-B), (A2-B) and $\xi_{n}\in L^{p}(\Omega,\mathcal{F}_{0},\mathbb{P};E_{n})$. By the first part of Proposition 6.1 (which also uses (A4-B)), we find $W_{A_{n}}\in L^{p}(\Omega;C([0,T];B_{n}))$. By (F1-B) and Proposition B.1b (with $E=F=B_{n}$, $\alpha=0$ and $s=1$), we have $S_{n}*F_{n}(\,\cdot\,,u_{n})\in L^{p}(\Omega;C([0,T];B_{n}))$ for all $u_{n}\in L^{p}(\Omega;C([0,T];B_{n}))$. This shows that $\Phi_{n,T}$ is well-defined. A straightforward estimate involving Assumptions (A1-B)(A2-B) and (F1-B) yields, for all $u_{n},v_{n}\in L^{p}(\Omega;C([0,T];B_{n}))$,
$\|\Phi_{n,T}(u_{n})-\Phi_{n,T}(v_{n})\|_{L^{p}(\Omega;C([0,T];B_{n}))}\leq\widetilde{M}_{S}\widetilde{L}_{F}T\,\|u_{n}-v_{n}\|_{L^{p}(\Omega;C([0,T];B_{n}))}.\qed$
Consequently, the proof of the following global well-posedness result is entirely analogous to that of Proposition 5.3:
Proposition 6.3.
Suppose that Assumptions (A1-B), (A2-B), (A4-B) hold with $\theta+2\beta<1$, and (F1-B) is satisfied. Let $n\in\overline{\mathbb{N}}$, $p\in[1,\infty)$, $\xi_{n}\in L^{p}(\Omega,\mathcal{F}_{0},\mathbb{P};E_{n})$ and $T\in(0,\infty)$ be given. Then (5.1) has a unique global mild solution $X_{n}$ in $L^{p}(\Omega;C([0,T];B_{n}))$.
Under the additional convergence assumptions (A3-B), (IC-B) and (F2-B), we can again set out to prove discrete-to-continuum convergence of $B_{n}$-valued global mild solutions to (5.1) by showing that all the expressions appearing in the fixed point maps $\widetilde{\Phi}_{n,T}$ from (5.5) are continuous as mappings on $L^{p}(\Omega;C([0,T];\widetilde{B}))$. For the first term, the following can be proven exactly in the same way as Lemma 5.5:
Lemma 6.4.
If Assumptions (A1-B)(A3-B) and (IC-B) are satisfied, then we have $\widetilde{S}_{n}\otimes\widetilde{\xi}_{n}\to S_{\infty}\otimes\xi_{\infty}$ in $L^{p}(\Omega;C([0,T];\widetilde{B}))$ as $n\to\infty.$
The following is an analog to Lemma 5.6:
Lemma 6.5.
Suppose that Assumptions (A1-B)(A3-B)(F1-B) and (F2-B) are satisfied. Let $p\in[1,\infty)$ and $u\in L^{p}(\Omega;C([0,T];B_{\infty}))$ be given. Then we have
$\widetilde{S}_{n}*\widetilde{F}_{n}(\,\cdot\,,u)\to S_{\infty}*F_{\infty}(\,\cdot\,,u)\quad\text{in }L^{p}(\Omega;C([0,T];\widetilde{B}))\text{ as }n\to\infty.$ (6.11)
Proof.
As in Lemma 5.6, we split up the statement into the following two assertions:
  1. (i)
    $\widetilde{S}_{n}*\widetilde{F}_{n}(\,\cdot\,,u)-\widetilde{S}_{n}*F_{\infty}(\,\cdot\,,u)\to 0$ in $L^{p}(\Omega;C([0,T];\widetilde{B}))$ as $n\to\infty$;
  2. (ii)
    $\widetilde{S}_{n}*F_{\infty}(\,\cdot\,,u)\to S_{\infty}*F_{\infty}(\,\cdot\,,u)$ in $L^{p}(\Omega;C([0,T];\widetilde{B}))$ as $n\to\infty$.
Part (i) is shown exactly as Lemma 5.6(i), up to replacing $E_{\infty}$ by $B_{\infty}$ (or $\widetilde{B}$).
For (ii), we instead note that (F1-B) implies, for almost every $\omega\in\Omega$,
$t\mapsto F_{\infty}(\omega,t,u(\omega,t))\in L^{p}(0,T;B_{\infty}).$ (6.12)
Hence, in order to argue as in Lemma 5.6(ii), one needs to apply Proposition B.3a and Corollary B.2b with $E_{n}\coloneqq B_{n}$ for all $n\in\overline{\mathbb{N}}$ and $\widetilde{E}\coloneqq\widetilde{B}$. ∎
With these auxiliary results in place, we can prove the first main discrete-to-continuum approximation result for solutions to (6.1) with globally Lipschitz drift coefficients of linear growth, analogously to Theorem 5.4:
Theorem 6.6.
Let Assumptions (A1-B)(A4-B)(F1-B), (F2-B) and (IC-B) be satisfied, with $\theta+2\beta<1$ and $p\in[1,\infty)$. Denoting by $X_{n}$ the unique $B_{n}$-valued global mild solution to (5.1), we have
$\widetilde{X}_{n}\to X_{\infty}\quad\text{in }L^{p}(\Omega;C([0,T];\widetilde{B}))\quad\text{as }n\to\infty.$ (6.13)
Proof.
By Proposition 6.2 and (A1-B), for small enough $T_{0}\in(0,\infty)$ there exists a constant $c\in[0,1)$, depending only on $\widetilde{L}_{F}$, $\widetilde{M}_{S}$, $\widetilde{M}_{\Lambda}$ and $\widetilde{M}_{\Pi}$, such that, for all $u,v\in L^{p}(\Omega;C([0,T_{0}];B_{\infty}))$,
$\sup\nolimits_{n\in\mathbb{N}}\,\|\widetilde{\Phi}_{n,T_{0}}(u)-\widetilde{\Phi}_{n,T_{0}}(v)\|_{L^{p}(\Omega;C([0,T_{0}];\widetilde{B}))}\leq c\|u-v\|_{L^{p}(\Omega;C([0,T_{0}];\widetilde{B}))}.$ (6.14)
By Proposition 6.3, there exists a unique global solution $X_{n}\in L^{p}(\Omega;C([0,T];B_{n}))$ to (5.1) for every $n\in\overline{\mathbb{N}}$. In particular, note that $X_{\infty}$ takes its values in $B_{\infty}\subseteq\widetilde{B}$. Thus, in order to finish the argument analogously to the proof of Theorem 5.4, it suffices to establish that $\widetilde{\Phi}_{n,T}(\phi)\to\Phi_{\infty,T}(\phi)$ in $L^{p}(\Omega;C([0,T];\widetilde{B}))$ as $n\to\infty$ for all $\phi\in L^{p}(\Omega;C([0,T];B_{\infty}))$. This is precisely the combined content of (the second part of) Proposition 6.1, along with Lemmas 6.4 and 6.5. ∎
6.2. Locally Lipschitz drifts
As in Section 5.2, we can extend Theorem 6.6 to the locally Lipschitz setting. Namely, we assume that
  1. (F1-B)
    Assumption (F1) holds with $(B_{n})_{n\in\overline{\mathbb{N}}}$ in place of $(E_{n})_{n\in\overline{\mathbb{N}}}$. For every $r>0$, the $B_{n}$-valued local Lipschitz and local boundedness constants are respectively denoted by $\widetilde{L}_{F}^{(r)}$ and $\widetilde{C}_{F,0}^{(r)}$.
Then, arguing in the same way as Theorem 5.9, we obtain the following result.
Theorem 6.7.
Suppose that Assumptions (A1-B)(A4-B)(F1-B), (F2-B) and (IC-B) are satisfied, with $\theta+2\beta<1.$ For $n\in\overline{\mathbb{N}}$, let $(X_{n}(t))_{t\in[0,\sigma_{n})}$ be the maximal local solution to (6.1) with explosion time $\sigma_{n}\colon\Omega\to[0,T]$. Then we have
  1. (i)
    $\widetilde{X}_{n}\mathbf{1}_{[0,\sigma_{\infty}\wedge\sigma_{n})}\to X_{\infty}\mathbf{1}_{[0,\sigma_{\infty})}$ in $L^{0}(\Omega\times[0,T];\widetilde{B})$ as $n\to\infty$.
If, moreover, $\sigma_{n}=T$ holds $\mathbb{P}$-a.s. for all $n\in\mathbb{N}$ and $p\in[1,\infty)$ is such that
$\sup_{n\in\mathbb{N}}\|X_{n}\|_{L^{p}(\Omega;C([0,T];B_{n}))}<\infty,$ (6.15)
then the following assertions also hold:
  1. (ii)
    We have $\sigma_{\infty}=T$, $\mathbb{P}$-a.s.
  2. (iii)
    If $p\in(1,\infty)$, then for all $p^{-}\in[1,p)$ we have
    $\widetilde{X}_{n}\to X_{\infty}\quad\text{ in }L^{p^{-}}(\Omega;C([0,T];\widetilde{B}))\text{ as }n\to\infty.$ (6.16)
6.3. Global well-posedness and convergence for dissipative drifts
In this section, we consider a class of equations whose drift coefficients satisfy not only (F1-B) (which would only guarantee local well-posedness), but additionally a type of dissipativity condition, also used in [56], which allows us to deduce global existence.
Let the subdifferential $\partial\|x_{n}\|_{B_{n}}$ of the norm $\|\cdot\|_{B_{n}}$ at $x_{n}\in B_{n}$ be given by
$\partial\|x_{n}\|_{B_{n}}=\{x^{*}_{n}\in B^{*}_{n}:\|x^{*}_{n}\|_{B^{*}_{n}}\leq 1\text{ and }{}_{B_{n}}\langle x_{n},x^{*}_{n}\rangle_{B^{*}_{n}}=\|x_{n}\|_{B_{n}}\}.$ (6.17)
The assumptions on $(F_{n})_{n\in\overline{\mathbb{N}}}$ under which we can derive global well-posedness, see Lemma 6.9 below (which is a simplified version of [56, Theorem 4.3] for equations driven by additive noise), are as follows:
  1. (F1′′-B)
    Let the conditions of (F1-B) hold. Suppose that there exist $a^{\prime},b^{\prime}\in[0,\infty)$ and $N\in\mathbb{N}$ such that for all $n\in\overline{\mathbb{N}}$, $(\omega,t)\in\Omega\times[0,T]$, $x_{n}\in\mathsf{D}(A_{n}|_{B_{n}})$, $x^{*}_{n}\in\partial\|x_{n}\|_{B_{n}}$ and $y_{n}\in B_{n}$ we have
    ${}_{B_{n}}\langle-A_{n}x_{n}+F_{n}(\omega,t,x_{n}+y_{n}),x_{n}^{*}\rangle_{B^{*}_{n}}\leq a^{\prime}(1+\|y_{n}\|_{B_{n}})^{N}+b^{\prime}\|x_{n}\|_{B_{n}}.$ (6.18)
If the semigroups $(S_{n}(t)|_{B_{n}})_{t\geq 0}$ are contractive on $B_{n}$, i.e., if $\widetilde{M}_{S}=1$ in (A2-B), then we know that $A_{n}|_{B_{n}}$ is accretive, i.e.,
${}_{B_{n}}\langle A_{n}x_{n},x^{*}_{n}\rangle_{B^{*}_{n}}\geq 0\quad\text{for all }x_{n}\in\mathsf{D}(A_{n}|_{B_{n}}),\,x^{*}_{n}\in\partial\|x\|_{B_{n}}.$ (6.19)
Thus, in this case, it suffices to check that
${}_{B_{n}}\langle F_{n}(\omega,t,x_{n}+y_{n}),x^{*}_{n}\rangle_{B^{*}_{n}}\leq a^{\prime}(1+\|y_{n}\|_{B_{n}})^{N},$ (6.20)
in order to establish that (F1′′-B) holds for $b^{\prime}=0$. The next example shows how the relation (6.20) can be verified in our situation of main interest. It is an elaborated version of [56, Example 4.2].
Example 6.8.
Given $n\in\overline{\mathbb{N}}$, let $B_{n}\coloneqq C(\mathcal{M}_{n})$ be the space of continuous real-valued functions on a compact Hausdorff space $\mathcal{M}_{n}$ equipped with the supremum norm $\|u_{n}\|_{B_{n}}\coloneqq\sup_{\xi\in\mathcal{M}_{n}}\lvert u_{n}(\xi)\rvert$. In this case, for all $u_{n}\in B_{n}$, the subdifferential $\partial\|u_{n}\|_{B_{n}}$ is the weak-closed convex hull of
$\{r\delta_{\hat{\xi}}:r\in\operatorname{sgn}u_{n}(\hat{\xi})\text{ for }\hat{\xi}\in\mathcal{M}_{n}\text{ such that }\|u_{n}\|_{B_{n}}=\lvert u_{n}(\hat{\xi})\rvert\},$ (6.21)
where $\delta_{\xi}\in C(\mathcal{M}_{n})^{*}$ denotes the evaluation functional at $\xi\in\mathcal{M}_{n}$ and, for $y\in\mathbb{R}$,
$\operatorname{sgn}y\coloneqq\begin{cases}\{-1\},&\text{if }y<0;\\ \{-1,1\},&\text{if }y=0;\\ \{1\},&\text{if }y>0.\end{cases}$ (6.22)
Indeed, since subdifferential sets are convex by definition and $\partial\|u_{n}\|_{B_{n}}$, being contained in the closed unit ball in $B_{n}^{*}$, is weak compact by the Banach–Alaoglu theorem, the Krein–Milman theorem implies that it suffices to argue that the extreme points of $\partial\|u_{n}\|_{B_{n}}$ are precisely given by (6.21). This, in turn, follows from a characterization of the extreme points of the closed unit ball in $C(\mathcal{M}_{n})^{*}$ due to Arens and Kelley [2].
Moreover, suppose that $(F_{n})_{n\in\overline{\mathbb{N}}}$ is a family of Nemytskii operators on $(B_{n})_{n\in\overline{\mathbb{N}}}$ (see equation (6.2)), generated by a family $(f_{n})_{n\in\overline{\mathbb{N}}}$ of functions satisfying the polynomial form introduced in (3.19)–(3.20). Fixing $n\in\overline{\mathbb{N}}$, $(\omega,t)\in\Omega\times[0,T]$ and $u_{n},v_{n}\in B_{n}$, inequality (6.20) becomes
${}_{B_{n}}\langle F_{n}(\omega,t,u_{n}+v_{n}),x^{*}_{n}\rangle_{B^{*}_{n}}\leq a^{\prime}(1+\|v_{n}\|_{B_{n}})^{N}\quad\text{for all }x_{n}^{*}\in\partial\|u_{n}\|_{B_{n}}.$ (6.23)
Since the inequality is preserved under convex combinations and weak limits of $x_{n}^{*}$, the above characterization of $\partial\|u_{n}\|_{B_{n}}$ shows that it only needs to be checked for $x_{n}^{*}=r\delta_{\hat{\xi}}$, where $r\in\operatorname{sgn}u_{n}(\hat{\xi})$ for $\hat{\xi}\in\mathcal{M}_{n}$ such that $\|u_{n}\|_{B_{n}}=\lvert u_{n}(\hat{\xi})\rvert$. That is, it suffices to verify that
$rf_{n}(\omega,t,u_{n}(\hat{\xi})+v_{n}(\hat{\xi}))\leq a^{\prime}(1+\|v_{n}\|_{B_{n}})^{N}.$ (6.24)
Indeed, for $(f_{n})_{n\in\overline{\mathbb{N}}}$ satisfying (3.19)–(3.20), we can establish the estimate
$rf_{n}(\omega,t,y+z)\lesssim_{(c,C,k)}(1+\lvert z\rvert)^{2k+1}$ (6.25)
for all $(\omega,t)\in\Omega\times[0,T]$, $y,z\in\mathbb{R}$ and $r\in\operatorname{sgn}y$. This implies the existence of a constant $a^{\prime}\in[0,\infty)$, depending only on $c,C\in(0,\infty)$ from (3.20) and $k\in\mathbb{N}_{0}$ from (3.19), such that (6.23) holds with $N=2k+1$.
Lemma 6.9.
Let Assumptions (A1-B)(A4-B) and (F1′′-B) hold with $\widetilde{M}_{S}=1$, let $n\in\overline{\mathbb{N}}$ and suppose that $\theta+2\beta<1$. If $\xi_{n}\in L^{p}(\Omega,\mathcal{F}_{0},\mathbb{P};B_{n})$ for some $p\in[1,\infty)$, then the maximal solution $(X_{n}(t))_{t\in[0,\sigma_{n})}$ to (6.1) is global (i.e., it holds $\mathbb{P}$-a.s. that $\sigma_{n}=T$), and we have
$\|X_{n}\|_{L^{p}(\Omega;C([0,T];B_{n}))}\lesssim_{(a^{\prime},b^{\prime},T,N)}1+\|\xi_{n}\|_{L^{p}(\Omega;B_{n})}+\|W_{A_{n}}\|_{L^{Np}(\Omega;C([0,T];B_{n}))}^{N}.$ (6.26)
Proof.
Fix $n\in\overline{\mathbb{N}}$. For each $m\in\mathbb{N}$, let $F_{n,m}$ denote the globally Lipschitz retraction of $F_{n}$ onto the closed ball of radius $m$ around $0\in B_{n}$, cf. (5.52) (replacing $E_{n}$ by $B_{n}$). Then $F_{n,m}$ satisfies the global Lipschitz and (global) linear growth estimates in (F1-B), hence by Proposition 6.3 there exists a unique global $B_{n}$-valued mild solution $X_{n,m}\in L^{p}(\Omega;C([0,T];B_{n}))$ to (5.1) with drift coefficient operator $F_{n,m}$. By the triangle inequality,
$\|X_{n,m}\|_{L^{p}(\Omega;C([0,T];B_{n}))}$ (6.27)
$\qquad\leq\|S_{n}\otimes\xi_{n}+S_{n}*F_{n,m}(\,\cdot\,,X_{n,m})\|_{L^{p}(\Omega;C([0,T];B_{n}))}+\|W_{A_{n}}\|_{L^{p}(\Omega;C([0,T];B_{n}))}.$
As shown in the proof of [56, Theorem 4.3], $F_{n,m}$ inherits the dissipativity estimate (F1′′-B), with the same constants $a^{\prime}$, $b^{\prime}$ and $N$, from $F_{n}$. Thus, by [56, Lemma 4.4],
$\|S_{n}\otimes\xi_{n}+S_{n}*F_{n,m}(\,\cdot\,,X_{m})\|_{C([0,T];B_{n})}$ (6.28)
$\qquad\leq e^{b^{\prime}T}\biggl{(}\|\xi\|_{B_{n}}+a^{\prime}\int_{0}^{T}(1+\|W_{A_{n}}(s)\|_{B_{n}})^{N}\mathop{}\!\mathrm{d}s\biggr{)}$ (6.29)
[…]
It follows that
$\|S_{n}\otimes\xi_{n}+S_{n}*F_{n,m}(\,\cdot\,,X_{m})\|_{L^{p}(\Omega;C([0,T];B_{n}))}$ (6.31)
$\qquad\lesssim_{(a^{\prime},b^{\prime},T,N)}1+\|\xi_{n}\|_{L^{p}(\Omega;B_{n})}+\|W_{A_{n}}\|_{L^{Np}(\Omega;C([0,T];B_{n}))}^{N}.$
Note that $W_{A_{n}}\in L^{Np}(\Omega;C([0,T];B_{n}))$ by Proposition 6.1 (with $Np$ taking the role of $p$). Combining these estimates, we find
$\sup_{m\in\mathbb{N}}\|X_{n,m}\|_{L^{p}(\Omega;C([0,T];B_{n}))}\lesssim_{(a^{\prime},b^{\prime},T,N)}1+\|\xi_{n}\|_{L^{p}(\Omega;B_{n})}+\|W_{A_{n}}\|_{L^{Np}(\Omega;C([0,T];B_{n}))}^{N},$ (6.32)
so the result follows by Theorem 5.10(iv)(v) applied to $Y_{m}\coloneqq X_{n,m}$. ∎
Combined with Theorem 6.7(ii)(iii), whose uniform-boundedness hypothesis (6.15) is verified by the combination of Lemma 6.9 and Proposition 6.1 under the assumption in the following corollary, we derive:
Corollary 6.10.
Suppose that (A1-B)(A4-B)(F1′′-B)(F2-B) and (IC-B) are satisfied with $\widetilde{M}_{S}=1$, $p\in(1,\infty)$ and $\theta+2\beta<1$. Then for any $p^{-}\in[1,p)$, the sequence $((X_{n})_{t\in[0,T]})_{n\in\overline{\mathbb{N}}}$ of $B_{n}$-valued global solutions to (6.1) satisfies
$\widetilde{X}_{n}\to X_{\infty}\quad\text{ in }L^{p^{-}}\!(\Omega;C([0,T];\widetilde{B}))\text{ as }n\to\infty.$ (6.33)
7. Outlook
In this section we suggest some possible extensions of both the convergence of graph-discretized Whittle–Matérn SPDEs shown in Section 3 as well as the underlying abstract results from Sections 46.
As discussed in Subsection 3.5, the approximation results from Theorem 3.10 for (3.16) might be extended to broader classes of domains $\mathcal{M}$, connectivity length regimes $(h_{n})_{n\in\mathbb{N}}$, coefficient functions $\tau,\kappa\colon\mathcal{M}\to[0,\infty)$ and (fractional) powers ${s\in(0,\infty)}$. Possible advancements to this end include the establishment of more general $L^{\infty}$-convergence results for graph Laplacian eigenfunctions (or convergence of Whittle–Matérn operators without using spectral convergence), as well as uniform $L^{\infty}$-boundedness of the semigroups (for instance via heat kernel estimates). Under more restrictive assumptions on the connectivity parameter regime, rates of convergence for the case of purely spatial (i.e., stationary) graph-discretized linear SPDEs were established in [71]. The same might be possible in the linear spatiotemporal setting since the discrete-to-continuum Trotter–Kato theorem can be extended to yield error estimates, see [51, Section 2.2]. The semilinear cases, however, appear to be out of reach for the methods used in this work.
The proofs of the abstract discrete-to-continuum approximation results from Sections 46 largely rely on incorporating projection and lifting operators into arguments from [55, 56] in an appropriate way. By adapting other proofs from these sources along similar lines, it is likely that our results can be extended further, in particular enabling us to relax the simplifying assumptions that the UMD Banach spaces $(E_{n})_{n\in\overline{\mathbb{N}}}$ have type $2$ and that the driving noise is additive. In fact, we expect more generally that many results asserting continuous dependence of stochastic evolution problems on their coefficients can be extended to discrete-to-continuum approximation theorems via this procedure.
One particular type of problem for which this would be interesting is the class of stochastic evolution inclusions, whose drift operators are allowed to be multi-valued; this occurs, for instance, in the Langevin setting if $F_{n}=\partial\varphi_{n}$ is the subdifferential of a convex and lower-semicontinuous but non-differentiable functional $\varphi_{n}$ on the state space, taking values in (,]($\infty$,$\infty$]( - ∞ , ∞ ]. Continuous dependence results for stochastic inclusion problems have been established in two different settings in [35, 73]; however, neither of these covers the class of semilinear inclusions driven by cylindrical (i.e., white) noise. Not much theory appears to be available for such problems, with even the question of well-posedness (for fixed $n\in\overline{\mathbb{N}}$) being highly nontrivial. In fact, to the best of our knowledge, the only results in this direction concern the (important) subclass of stochastic reflection problems (also known as Skorokhod problems in the scalar-valued case), given by
[…]
where $\Gamma$ is a convex subset of a (Hilbertian) state space $H$ and the indicator functional $I_{\Gamma}\colon H\to(-\infty,\infty]$ is defined to vanish on $\Gamma$ and equal $\infty$ outside of it. In the first work on this problem, Nualart and Pardoux [66] used a direct approach to show existence and uniqueness in a setting which corresponds to $H\coloneqq L^{2}(0,1)$, $A\coloneqq-\frac{\mathop{}\!\mathrm{d}^{2}}{\mathop{}\!\mathrm{d}x^{2}}$ with homogeneous Dirichlet or Neumann boundary conditions, and $\Gamma\coloneqq K_{0}$, where
$K_{\alpha}\coloneqq\{u\in L^{2}(0,1):u(x)\geq-\alpha\text{ for a.e.\ }x\in(0,1)\},\quad\alpha\in[0,\infty).$
In [70], the authors first use the theory of Dirichlet forms to establish well-posedness of (7.1) in the case that $\Gamma$ is a “regular” convex subset of a general Hilbert space $H$, a condition which includes $\Gamma\coloneqq K_{\alpha}$ for $\alpha>0$ but not $K_{0}$, which is treated separately using different techniques. Lastly, the work [4] describes a variational approach to study (7.1) in a similar setting under the assumption that 00 belongs to the interior of $\Gamma$, which excludes the choice $\Gamma\coloneqq K_{0}$. We point out that the argument used on [4, p. 362] to extract a weak-convergent sequence from the set $(u_{\varepsilon})_{\varepsilon>0}$ in the dual space of $L^{\infty}(0,T;H)$ appears to be flawed, as it seems to imply that the closed unit ball of the dual of this (non-separable) space is sequentially compact, which is not the case. Hence, the argument would need to be finished using a generalized subsequence (also known as a subnet) converging in the weak sense to some $u^{*}$. For this reason, and since the theory for convergence of Dirichlet forms and their associated processes is well established—see for instance [52] for general results and [88] for an application to Markov chain Monte Carlo scaling—the setting of [70] is perhaps the most promising for an attempt at establishing discrete-to-continuum convergence results for (7.1).
Conflict of interest and data availability statement
The authors declare that they have no conflict of interest. Data sharing is not applicable to this article as no datasets were generated or analyzed during the current study.
Appendix A Proofs of intermediate results in Section 3.3
We start with the proof of Lemma 3.11 which establishes spectral convergence rates for the Laplace–Beltrami operator on the torus, discretized by a uniform grid.
Proof of Lemma 3.11.
First suppose $m=1$. In this case, the continuum Laplace–Beltrami operator reduces to the second derivative $-\frac{\mathop{}\!\mathrm{d}^{2}}{\mathop{}\!\mathrm{d}x^{2}}$ with periodic boundary conditions. Its eigenvalues and $L^{2}(\mathcal{M})$-normalized eigenfunctions are respectively given, for all $j\in\mathbb{N}$ and $x\in[0,1]$, by
$\lambda_{\infty}^{(j)}=\begin{cases}j^{2}\pi^{2}&\text{if $j$ is even},\\ (j-1)^{2}\pi^{2}&\text{if $j$ is odd};\end{cases}\quad\psi_{\infty}^{(j)}(x)=\begin{cases}1&\text{if }j=1,\\ \sqrt{2}\sin(j\pi x)&\text{if }j\text{ is even},\\ \sqrt{2}\cos((j-1)\pi x)&\text{if }j\text{ is odd}.\end{cases}$ (A.1)
That is, 00 is an eigenvalue corresponding to the constant $1$ eigenfunction, and $(2k)^{2}\pi^{2}$ is an eigenvalue with eigenfunctions $x\mapsto\sin(2k\pi x),\cos(2k\pi x)$ for all $k\in\mathbb{N}$.
The eigenvalues $(\lambda_{n}^{(j)})_{j=1}^{n}$ of the corresponding graph Laplacian, which in this case reduces to the finite difference approximation of the second derivative on the grid with $n\in\mathbb{N}$ points, are given by
$\lambda_{n}^{(j)}=\begin{cases}4n^{2}\sin^{2}\Bigl{(}\frac{\pi j}{2n}\Bigr{)}&\text{if $j$ is even},\\ 4n^{2}\sin^{2}\Bigl{(}\frac{\pi(j-1)}{2n}\Bigr{)}&\text{if $j$ is odd}.\end{cases}$ (A.2)
The corresponding $L^{2}(\mathcal{M}_{n})$-normalized eigenfunctions are
$\psi_{n}^{(j)}(x_{n}^{(i)})=\begin{cases}1&\text{if }j=1,\\ (-1)^{i}&\text{if }j=n\text{ is even},\\ \sqrt{2}\sin\bigl{(}j\pi x_{n}^{(i)}\bigr{)}&\text{if }j\neq n\text{ and }j\text{ is even},\\ \sqrt{2}\cos\bigl{(}(j-1)\pi x_{n}^{(i)}\bigr{)}&\text{if }j\text{ is odd}.\end{cases}$ (A.3)
Let $j\in\{1,\dots,n\}$. Supposing that $j$ is even (the odd case being analogous), we can write
$\lambda_{\infty}^{(j)}-\lambda_{n}^{(j)}=j^{2}\pi^{2}\biggl{(}1-\frac{4n^{2}}{j^{2}\pi^{2}}\sin^{2}\Bigl{(}\frac{\pi j}{2n}\Bigr{)}\biggr{)}=j^{2}\pi^{2}\biggl{(}1-\biggl{[}\frac{2n}{j\pi}\sin\Bigl{(}\frac{\pi j}{2n}\Bigr{)}\biggr{]}^{2}\biggr{)},$ (A.4)
so that the estimate in (3.26) for $m=1$ follows from the elementary inequality $0\leq 1-(\sin(x)/x)^{2}\leq\frac{1}{3}x^{2},$ which is valid for all $x\in\mathbb{R}$. Estimate (3.27) is a consequence of the fact that the sine and cosine functions are $1$-Lipschitz; note that we only consider $j\in\{1,\dots,n-1\}$ to avoid the case where $j=n$ for even $n$.
The result for higher dimensions $m\in\mathbb{N}$ can be derived from the $m=1$ case. Indeed, by separation of variables in the continuum case, or by writing the discretized operator as a Kronecker sum of $m$ one-dimensional discretizations, one finds that the eigenvalues and eigenvectors of the $m$-dimensional operators are given by sums and products, respectively, of their $1$-dimensional counterparts. From this, one can deduce the desired result. ∎
Next we prove Theorem 3.14 regarding the convergence of the fractional resolvent operators $\widetilde{R}_{n}^{\beta^{\prime}}$ in various settings and norms.
Proof of Theorem 3.14.
Assertions ac can all be shown using analogous arguments. Thus, we only provide a detailed proof for part b, being the most involved case, and subsequently summarize the changes needed for a and c.
b Step 1 (Setup and notation). The operator $\widetilde{R}_{n}^{\beta^{\prime}}$ acts on functions $f\in L^{2}(\mathcal{M})$ in the following way:
$\widetilde{R}_{n}^{\beta^{\prime}}f$ $=\sum_{j=1}^{n}\bigl{(}1+[\lambda^{(j)}_{n}]^{s}\bigr{)}^{-{\beta^{\prime}}}\langle\Pi_{n}f,\psi^{(j)}_{n}\rangle_{L^{2}(\mathcal{M}_{n})}\Lambda_{n}\psi^{(j)}_{n}$ (A.5)
$=\sum_{j=1}^{n}\bigl{(}1+[\lambda^{(j)}_{n}]^{s}\bigr{)}^{-{\beta^{\prime}}}\langle f,\Lambda_{n}\psi^{(j)}_{n}\rangle_{L^{2}(\mathcal{M})}\Lambda_{n}\psi^{(j)}_{n}.$ (A.6)
Here, we used the fact that $\Pi_{n}^{*}=\Lambda_{n}$ (see (3.7)) on the second line. Using the tensor notation from Section 2.1 and denoting $\widetilde{\psi}^{(j)}_{n}\coloneqq\Lambda_{n}\psi^{(j)}_{n}$, we can write these operators more concisely as follows:
$U_{n}\coloneqq\widetilde{R}_{n}^{{\beta^{\prime}}}=\sum_{j=1}^{n}\bigl{(}1+[\lambda^{(j)}_{n}]^{s}\bigr{)}^{-{\beta^{\prime}}}\,\widetilde{\psi}^{(j)}_{n}\otimes\widetilde{\psi}^{(j)}_{n}.$ (A.7)
By Assumption 3.5(ii), there exists a sequence of natural numbers $(k_{n})_{n\in\mathbb{N}}$ which satisfies the conditions of Theorem 3.13b, as well as the relation
$k_{n}\gg n^{\frac{m}{4s\beta}}\gtrsim n^{\frac{m}{4s\beta^{\prime}}}.$ (A.8)
We use the sequence $(k_{n})_{n\in\mathbb{N}}$ to define the following approximations for $n\in\mathbb{N}$:
$U_{n}^{1}$ $\coloneqq\sum_{j=1}^{k_{n}}\bigl{(}1+[\lambda^{(j)}_{\infty}]^{s}\bigr{)}^{-\beta^{\prime}}\psi^{(j)}_{\infty}\otimes\psi^{(j)}_{\infty},$ (A.9)
$U_{n}^{2}$ $\coloneqq\sum_{j=1}^{k_{n}}\bigl{(}1+[\lambda^{(j)}_{\infty}]^{s}\bigr{)}^{-\beta^{\prime}}\widetilde{\psi}^{(j)}_{n}\otimes\widetilde{\psi}^{(j)}_{n},$ (A.10)
$U_{n}^{3}$ $\coloneqq\sum_{j=1}^{k_{n}}\bigl{(}1+[\lambda^{(j)}_{n}]^{s}\bigr{)}^{-\beta^{\prime}}\widetilde{\psi}^{(j)}_{n}\otimes\widetilde{\psi}^{(j)}_{n}.$ (A.11)
For notational convenience, we will abbreviate the $\mathscr{L}(L^{2}(\mathcal{M});L^{\infty}(\mathcal{M}))$-norm by $\|\,\cdot\,\|_{2\to\infty}$ throughout this proof. We will make repeated use of the following estimate: Given an operator of the form $U=\sum_{j}\alpha_{j}e_{j}\otimes f_{j}$ for some scalars $(\alpha_{j})_{j}\subseteq\mathbb{R}$, an orthonormal system $(e_{j})_{j}\subseteq L^{2}(\mathcal{M})$ and some functions $(f_{j})_{j}\subseteq L^{\infty}(\mathcal{M})$, we have by the Cauchy–Schwarz inequality:
$\|U\|_{2\to\infty}\leq\sup\nolimits_{j}\|f_{j}\|_{L^{\infty}(\mathcal{M})}\biggl{(}\sum\nolimits_{j}\lvert\alpha_{j}\rvert^{2}\biggr{)}^{\frac{1}{2}}.$ (A.12)
Moreover, it is immediate from the definition of the $\|\,\cdot\,\|_{2\to\infty}$-norm that
$\|h\otimes f\|_{2\to\infty}=\|h\|_{L^{2}(\mathcal{M})}\|f\|_{L^{\infty}(\mathcal{M})}\quad\text{for all }h\in L^{2}(\mathcal{M})\text{ and }f\in L^{\infty}(\mathcal{M}).$ (A.13)
Step 2 ($U_{\infty}-U_{n}^{1}$ and $U_{n}^{3}-U_{n}$). For the difference between $U_{\infty}$ and $U_{n}^{1}$, we find using (A.12) and Assumption 3.9:
$\|U_{\infty}-U_{n}^{1}\|_{2\to\infty}^{2}\leq M^{2}_{\psi,\infty}\sum_{j=k_{n}+1}^{\infty}\bigl{(}1+[\lambda^{(j)}_{\infty}]^{s}\bigr{)}^{-2\beta^{\prime}}.$ (A.14)
Recalling Weyl’s law (3.12), which implies that $\bigl{(}1+[\lambda^{(j)}_{\infty}]^{s}\bigr{)}^{-2\beta^{\prime}}\eqsim_{\mathcal{M}}j^{-4s\beta^{\prime}/m}$, we observe that the series on the right-hand side is convergent precisely when $\beta^{\prime}>\frac{m}{4s}$. Since $k_{n}\to\infty$ as $n\to\infty$, this implies $U_{\infty}\to U_{n}^{1}$ in $\mathscr{L}(L^{2}(\mathcal{M});L^{\infty}(\mathcal{M}))$.
Similarly, for the difference between $U_{n}^{3}$ and $U_{n}$, we have
$\|U_{n}^{3}-U_{n}\|_{2\to\infty}^{2}$ $\leq M_{\psi,\infty}^{2}\sum_{j=k_{n}+1}^{n}\bigl{(}1+[\lambda^{(j)}_{n}]^{s}\bigr{)}^{-2\beta^{\prime}}\leq M_{\psi,\infty}^{2}\bigl{(}n-k_{n}\bigr{)}\bigl{(}1+[\lambda^{(k_{n})}_{n}]^{s}\bigr{)}^{-2\beta^{\prime}},$ (A.15)
where the second inequality is due to the non-decreasing order of $(\lambda_{n}^{(j)})_{j=1}^{n}$. Moreover, as a consequence of Theorem 3.12, there exists a constant $C^{\prime}>0$ such that, $\widetilde{\mathbb{P}}$-a.s., for all $n\in\mathbb{N}$ and $j\in\{1,\dots,k_{n}\}$, we have $\lambda_{n}^{(j)}\geq C^{\prime}\lambda_{\infty}^{(j)}$. In particular, $\lambda_{n}^{(k_{n})}\gtrsim\lambda_{\infty}^{(k_{n})}$. Together with Weyl’s law, we find $1+[\lambda^{(k_{n})}_{n}]^{s}\gtrsim_{\mathcal{M}}k_{n}^{{2s}/{m}}$, so that the convergence of this difference is due to (A.8):
$\|U_{n}^{3}-U_{n}\|_{2\to\infty}^{2}\lesssim_{\mathcal{M}}M_{\psi,\infty}^{2}nk_{n}^{-{4s\beta^{\prime}}/{m}}\to 0\quad\text{as }n\to\infty.$ (A.16)
Step 3 ($U_{n}^{1}-U_{n}^{2}$). In order to show that
$U_{n}^{1}-U_{n}^{2}\to 0\quad\text{in }L^{0}(\widetilde{\Omega};\mathscr{L}(L^{2}(\mathcal{M});L^{\infty}(\mathcal{M})))\quad\text{as }n\to\infty,$ (A.17)
we first fix an arbitrary $\varepsilon>0$. Then, for all $\ell,n\in\mathbb{N}$ such that $k_{n}>\ell$, we split off the first $\ell$ terms and use the triangle inequality to obtain
$\|U_{n}^{1}-U_{n}^{2}\|_{2\to\infty}\leq\sum_{j=1}^{\ell}\bigl{(}1+[\lambda^{(j)}_{\infty}]^{s}\bigr{)}^{-\beta^{\prime}}\|\psi^{(j)}_{\infty}\otimes\psi^{(j)}_{\infty}-\widetilde{\psi}^{(j)}_{n}\otimes\widetilde{\psi}^{(j)}_{n}\|_{2\to\infty}$ (A.18)
$+\bigg{\lVert}\sum_{j=\ell+1}^{k_{n}}\bigl{(}1+[\lambda^{(j)}_{\infty}]^{s}\bigr{)}^{-\beta^{\prime}}\psi^{(j)}_{\infty}\otimes\psi^{(j)}_{\infty}\bigg{\rVert}_{2\to\infty}+\bigg{\lVert}\sum_{j=\ell+1}^{k_{n}}\bigl{(}1+[\lambda^{(j)}_{\infty}]^{s}\bigr{)}^{-\beta^{\prime}}\widetilde{\psi}^{(j)}_{n}\otimes\widetilde{\psi}^{(j)}_{n}\bigg{\rVert}_{2\to\infty}.$
Using the triangle inequality once more, followed by (A.13) and Assumption 3.9, the norms in the summation over $j\in\{1,\dots,\ell\}$ can be bounded by
$\|\psi^{(j)}_{\infty}\otimes\psi^{(j)}_{\infty}-\widetilde{\psi}^{(j)}_{n}\otimes\widetilde{\psi}^{(j)}_{n}\|_{2\to\infty}$ (A.19)
$\quad\leq\|\psi_{\infty}^{(j)}\otimes(\psi_{\infty}^{(j)}-\widetilde{\psi}_{n}^{(j)})\|_{2\to\infty}+\|(\psi_{\infty}^{(j)}-\widetilde{\psi}_{n}^{(j)})\otimes\widetilde{\psi}_{n}^{(j)}\|_{2\to\infty}$ (A.20)
$\quad=\|\psi_{\infty}^{(j)}-\widetilde{\psi}_{n}^{(j)}\|_{L^{\infty}(\mathcal{M})}+\|\psi_{\infty}^{(j)}-\widetilde{\psi}_{n}^{(j)}\|_{L^{2}(\mathcal{M})}\|\widetilde{\psi}_{n}^{(j)}\|_{L^{\infty}(\mathcal{M})}$ (A.21)
$\quad\leq(1+M_{\psi,\infty})\|\psi_{\infty}^{(j)}-\widetilde{\psi}_{n}^{(j)}\|_{L^{\infty}(\mathcal{M})},$ (A.22)
whereas the remaining two summations can be treated by arguing as for (A.14). Together, this yields
$\|U_{n}^{1}-U_{n}^{2}\|_{2\to\infty}\leq\begin{aligned} (1+M_{\psi,\infty})&\sum_{j=1}^{\ell}\bigl{(}1+[\lambda^{(j)}_{\infty}]^{s}\bigr{)}^{-\beta^{\prime}}\|\psi_{\infty}^{(j)}-\widetilde{\psi}_{n}^{(j)}\|_{L^{\infty}(\mathcal{M})}\\ &+2\biggl{(}\sum_{j=\ell+1}^{\infty}\bigl{(}1+[\lambda^{(j)}_{\infty}]^{s}\bigr{)}^{-2\beta^{\prime}}\biggr{)}^{\frac{1}{2}}.\end{aligned}$ (A.23)
Since we have already seen in Step 2 that the latter series converges, we can fix $\ell\in\mathbb{N}$ so large that the second sum on the right-hand side is less than $\frac{1}{2}\varepsilon$. Moreover, it follows from Theorem 3.13b that $\|\psi_{\infty}^{(j)}-\widetilde{\psi}_{n}^{(j)}\|_{L^{\infty}(\mathcal{M})}\to 0$ in $L^{0}(\widetilde{\Omega},\widetilde{\mathbb{P}})$ as $n\to\infty$ for every $j\in\{1,\dots,\ell\}$. In particular, there exists $N\in\mathbb{N}$ such that, for all $n\geq N$, the first sum on the right-hand side is less than $\frac{1}{2}\varepsilon$, and thus the whole right-hand side is less than $\varepsilon$, with probability ~$\widetilde{\mathbb{P}}\geq 1-\varepsilon$. This shows $\|U_{n}^{1}-U_{n}^{2}\|_{2\to\infty}\to 0$ in probability, as desired.
Step 4 ($U_{n}^{2}-U_{n}^{3}$). Finally, the difference $U_{n}^{2}-U_{n}^{3}$ can be treated in the same manner as Step 3, namely by writing, for all $\ell,n\in\mathbb{N}$ such that $k_{n}>\ell$,
[…]
Using the fact that, $\widetilde{\mathbb{P}}$-a.s., we have $\lambda_{n}^{(j)}\gtrsim\lambda_{\infty}^{(j)}$ for all $j\in\{1,\dots,k_{n}\}$ (see Step 2), the latter two summations can be bounded, up to a multiplicative constant, by the convergent series $\sum_{j=\ell+1}^{\infty}\bigl{(}1+[\lambda^{(j)}_{\infty}]^{s}\bigr{)}^{-2\beta^{\prime}}.$ Combined with the eigenvalue convergence asserted by Theorem 3.12, which can be applied to the remaining summation, we obtain $\|U_{n}^{2}-U_{n}^{3}\|_{2\to\infty}\to 0$ as $n\to\infty$, $\widetilde{\mathbb{P}}$-a.s., by arguing as in Step 3. Thus, we have shown part b.
a Replace (A.12) by the identity $\|U\|_{\mathscr{L}_{2}(L^{2}(\mathcal{M}))}^{2}=\sum\nolimits_{j}\lvert\alpha_{j}\rvert^{2}\|f_{j}\|_{L^{2}(\mathcal{M})}^{2},$ which follows directly from the definition of the Hilbert–Schmidt norm. Assumption 3.9 is not required since all the eigenfunctions are $L^{2}$-normalized. The sufficiency of Assumption 3.5(i) and the $\widetilde{\mathbb{P}}$-a.s. convergence in the conclusion are due to the use of Theorem 3.13a instead of Theorem 3.13b.
c Recall from Setting 3.2 that $h_{n}\coloneqq n^{-\frac{1}{m}}$ by definition. In view of Lemma 3.11, we can take $k_{n}\coloneqq n-1$, hence neither of the bounds on $h_{n}$ from Assumption 3.5 is needed. Indeed, in the proof of b, the lower bound (A.8) on $k_{n}$ was only used in (A.16), which becomes $\|U_{n}^{3}-U_{n}\|_{2\to\infty}^{2}\lesssim_{\mathcal{M}}M_{\psi,\infty}^{2}k_{n}^{-{4s\beta^{\prime}}/{m}}$ in the current situation, and this tends to zero since, trivially, $k_{n}\to\infty$ as $n\to\infty$. ∎
Lastly, we prove Lemma 3.15 asserting the uniform ultracontractivity of the semigroups $(S_{n}(t))_{t\geq 0}$ associated to the (discretized) generalized Whittle–Matérn operators $-(\mathcal{L}_{n}^{\kappa,\tau})^{s}$.
Proof of Lemma 3.15.
a For $p=\infty$, the statement holds by Assumption 3.6(i). For $p=2$, we note that, for all $n\in\overline{\mathbb{N}}$, $t>0$, and $f\in L^{2}(\mathcal{M}_{n})$,
$\|S_{n}(t)f\|_{L^{\infty}(\mathcal{M}_{n})}=\|R_{n}^{\beta}(\operatorname{Id}_{n}+A_{n})^{\beta}S_{n}(t)f\|_{L^{\infty}(\mathcal{M}_{n})}\leq t^{-\beta}\|R_{n}^{\beta}\|_{2\to\infty}\|f\|_{L^{2}(\mathcal{M}_{n})},$ (A.25)
where we used (C.3), as $(S_{n}(t))_{t\geq 0}$ is a contractive analytic semigroup on $L^{2}(\mathcal{M}_{n})$. In the proof of Theorem 3.14b, we found $\|R_{n}^{\beta}\|_{2\to\infty}\leq M_{\psi,\infty}\sum_{j=1}^{n}(1+[\lambda_{n}^{(j)}]^{s})^{-2\beta}$ under Assumption 3.9. Since Assumption 3.5(ii) implies 3.5(i) with the same $\beta$, and since the estimate of $\|R_{n}^{\beta}\|_{2\to\infty}$ only involves the eigenvalues (and not the eigenfunctions), we can in both cases argue as in Theorem 3.14a, under Assumption 3.4(i), to deduce that, $\widetilde{\mathbb{P}}$-a.s., the right-hand side can be bounded independently of $n$. This proves the statement for $p=2$, hence by the Riesz–Thorin interpolation theorem [37, Theorem 1.3.4], the lemma holds for all $p\in[2,\infty]$, with $M_{S,p}\leq M_{S,2}^{\frac{2}{p}}M_{S,\infty}^{1-\frac{2}{p}}$.
b The differences with part a are the use of Theorem 3.14c instead of 3.14b, and the fact that Assumption 3.9 is automatically satisfied. ∎
Appendix B Fractional parabolic integration
Let $(E,\|\,\cdot\,\|_{E})$ be a Banach space. Suppose that $-A\colon\mathsf{D}(A)\subseteq E\to E$ generates a strongly continuous semigroup $(S(t))_{t\geq 0}$. This implies the existence of constants $M\in[1,\infty)$ and $w\in\mathbb{R}$ such that
$\|S(t)\|_{\mathscr{L}(E)}\leq Me^{wt}\quad\text{for all }t\in[0,\infty).$ (B.1)
Fixing $T\in(0,\infty)$, we define $k_{s}\colon\mathbb{R}\to\mathscr{L}(E)$ by $k_{s}(\tau)\coloneqq\frac{1}{\Gamma(s)}\tau^{s-1}S(s)\mathbf{1}_{[0,T]}(\tau)$ for $\tau\in\mathbb{R}$. For any function $f\colon(0,T)\to E$ such that the following Bochner integral converges in $E$ for $s\in(0,\infty)$ and a.e. $t\in(0,T)$, we set
$\mathfrak{I}_{A}^{s}f(t)\coloneqq k_{s}*f(t)=\frac{1}{\Gamma(s)}\int_{0}^{t}(t-\tau)^{s-1}S(t-\tau)f(\tau)\mathop{}\!\mathrm{d}\tau.$ (B.2)
For $s=0$ we set $\mathfrak{I}_{A}^{0}\coloneqq\operatorname{Id}_{E}$. The following properties of the fractional parabolic integration operator $\mathfrak{I}_{A}^{s}$ (for a single operator $A$) are well known, see [20, Proposition 5.9], and used throughout the main text. We will state them here for the sake of self-containedness.
Proposition B.1.
Suppose that $-A\colon\mathsf{D}(A)\subseteq E\to E$ generates a strongly continuous semigroup $(S(t))_{t\geq 0}\subseteq\mathscr{L}(E)$ satisfying (B.1). Then, for every $s\in[0,\infty)$, $p\in[1,\infty]$ and $T\in(0,\infty)$, we have:
  1. (a)
    $\mathfrak{I}_{A}^{s}$ is bounded from $L^{p}(0,T;E)$ to itself, with an operator norm depending only on $s$, $p$, $T$, $w$ and $M$.
If $(F,\|\,\cdot\,\|_{F})$ is a Banach space for which there exist $M^{\prime}\in[1,\infty)$ and $\alpha\in[0,\infty)$ such that
$S(t)\in\mathscr{L}(E;F)\quad\text{with}\quad\|S(t)\|_{\mathscr{L}(E;F)}\leq M^{\prime}t^{-\alpha/2}\quad\text{for all }t\in[0,\infty),$ (B.3)
and in addition we have either $p=1,\,s\geq 1+\frac{\alpha}{2}$ or $p>1,\,s>\frac{1}{p}+\frac{\alpha}{2}$, then
  1. (b)
    $\mathfrak{I}_{A}^{s}$ is bounded from $L^{p}(0,T;E)$ to $C([0,T];F)$, with an operator norm depending only on $s$, $p$, $T$ and $M^{\prime}$.
For a sequence of operators $(A_{n})_{n\in\overline{\mathbb{N}}}$ on Banach spaces which satisfy the appropriate discrete-to-continuum assumptions from the main text, the proposition above implies the following corollary regarding uniform boundedness of the sequence (~$(\widetilde{\mathfrak{I}}_{A_{n}}^{s})_{n\in\overline{\mathbb{N}}}$, where ~$\widetilde{\mathfrak{I}}_{A_{n}}^{s}\coloneqq\Lambda_{n}\mathfrak{I}_{A_{n}}^{s}\Pi_{n}$ for all $n\in\overline{\mathbb{N}}$. From this, in turn, one can derive Proposition B.3 below asserting the strong convergence of these operators.
Corollary B.2.
Let the Banach spaces $(E_{n},\|\,\cdot\,\|_{E_{n}})_{n\in\overline{\mathbb{N}}}$, $(\widetilde{E},\|\,\cdot\,\|_{\widetilde{E}})$ and the linear operators $(A_{n})_{n\in\overline{\mathbb{N}}}$ satisfy the assumptions of Theorem 2.3, and suppose that ${p\in[1,\infty]}$ and $s\in[0,\infty)$. The following assertions hold:
  1. (a)
    The sequence (~$(\widetilde{\mathfrak{I}}^{s}_{A_{n}})_{n\in\overline{\mathbb{N}}}$ is uniformly bounded in $\mathscr{L}(L^{p}(0,T;\widetilde{E}))$.
  2. (b)
    The sequence (~$(\widetilde{\mathfrak{I}}^{s}_{A_{n}})_{n\in\overline{\mathbb{N}}}$ is uniformly bounded in $\mathscr{L}(L^{p}(0,T;\widetilde{E});C([0,T];\widetilde{E}))$ if, either, $p=1$ and $s\geq 1$, or $p>1$ and $s>\frac{1}{p}$.
If the spaces $(E_{n})_{n\in\overline{\mathbb{N}}}$, $(B_{n})_{n\in\overline{\mathbb{N}}}$ and $\widetilde{B}$ are as in Assumptions (A1-B), (A2-B) and (A4-B), and we have $s>\frac{1}{p}+\frac{\theta}{2}$, then
  1. (c)
    the sequence (~$(\widetilde{\mathfrak{I}}^{s}_{A_{n}})_{n\in\overline{\mathbb{N}}}$ is uniformly bounded in $\mathscr{L}(L^{p}(0,T;E_{\infty});C([0,T];\widetilde{B}))$.
Proposition B.3.
Let the Banach spaces $(E_{n},\|\,\cdot\,\|_{E_{n}})_{n\in\overline{\mathbb{N}}}$, $(\widetilde{E},\|\,\cdot\,\|_{\widetilde{E}})$ and the linear operators $(A_{n})_{n\in\overline{\mathbb{N}}}$ satisfy the assumptions of Theorem 2.3. Let ${p\in[1,\infty]}$ and $s\in[0,\infty)$. The following assertions hold:
  1. (a)
    If either $p=1$ and $s\geq 1$, or $p>1$ and $s>\frac{1}{p}$, then we have ~$\widetilde{\mathfrak{I}}^{s}_{A_{n}}f\to\mathfrak{I}^{s}_{A_{\infty}}f$ in $C([0,T];\widetilde{E})$, as $n\to\infty$, for every $f\in L^{p}(0,T;E_{\infty})$.
Moreover, let Assumptions (A1-B)(A2-B)(A3-B) and (A4-B) hold.
  1. (b)
    If $s>\frac{1}{p}+\frac{\theta}{2}$, then we have ~$\widetilde{\mathfrak{I}}^{s}_{A_{n}}f\to\mathfrak{I}^{s}_{A_{\infty}}f$ in $C([0,T];\widetilde{B})$ as $n\to\infty$ for every $f\in L^{p}(0,T;E_{\infty})$.
Proof.
We only present the details of the argument for part b, the proof of a being similar.
Let $p\in[1,\infty)$, $s\in(\frac{1}{p}+\frac{\theta}{2},\infty)$, $f\in L^{p}(0,T;E_{\infty})$ and fix an arbitrary $\varepsilon>0$. By the density of $B_{\infty}$ in $E_{\infty}$ (see (A4-B)), and that of $B_{\infty}$-valued simple functions in $L^{p}(0,T;B_{\infty})$, there exists a function $g\colon[0,T]\to B_{\infty}$ of the form
$g=\sum_{j=1}^{K}\mathbf{1}_{(a_{j},b_{j})}\otimes x_{j},\quad K\in\mathbb{N};\>0\leq a_{j}<b_{j}\leq T,\;x_{j}\in B_{\infty}\text{ for all }j\in\{1,\dots,K\}$ (B.4)
such that
$\|f-g\|_{L^{p}(0,T;E_{\infty})}<\frac{\varepsilon}{4}\Bigl{(}\sup\nolimits_{n\in\overline{\mathbb{N}}}\|\widetilde{\mathfrak{I}}_{A_{n}}^{s}\|_{\mathscr{L}(L^{p}(0,T;E_{\infty});C([0,T];\widetilde{B}))}\Bigr{)}^{-1}.$ (B.5)
Note that the expression between the parentheses is finite by Corollary B.2c and can be assumed to be nonzero without loss of generality, as otherwise ~$\widetilde{\mathfrak{I}}_{A_{n}}^{s}=0$ for all $n\in\overline{\mathbb{N}}$ and the asserted convergence would be trivial. Thus, for every $n\in\mathbb{N}$,
~$\|\widetilde{\mathfrak{I}}_{A_{n}}^{s}f-\mathfrak{I}_{A}^{s}f\|_{C([0,T];\widetilde{B})}$ (B.6)
~$\quad\leq\|\widetilde{\mathfrak{I}}_{A_{n}}^{s}(f-g)\|_{C([0,T];\widetilde{B})}+\|\widetilde{\mathfrak{I}}_{A_{n}}^{s}g-\mathfrak{I}_{A}^{s}g\|_{C([0,T];\widetilde{B})}+\|\mathfrak{I}_{A}^{s}(g-f)\|_{C([0,T];\widetilde{B})}$ (B.7)
$\quad<\tfrac{1}{2}\varepsilon+\|\widetilde{\mathfrak{I}}_{A_{n}}^{s}g-\mathfrak{I}_{A}^{s}g\|_{C([0,T];\widetilde{B})}.$ (B.8)
For any $j\in\{1,\dots,K\}$, by (A3-B) and the discrete-to-continuum Trotter–Kato approximation theorem, we can choose $N_{j}\in\mathbb{N}$ so large that
$\|\widetilde{S}_{n}\otimes x_{j}-S\otimes x_{j}\|_{C([0,T];\widetilde{B})}<\frac{s\Gamma(s)}{2T^{s}K}\varepsilon\quad\text{for all }n\geq N_{j}.$ (B.9)
Thus, setting $N\coloneqq\max_{j=1}^{K}N_{j}$, we find for all $n\geq N$ and $t\in[0,T]$:
~$\|\widetilde{\mathfrak{I}}_{A_{n}}^{s}g(t)-\mathfrak{I}_{A}^{s}g(t)\|_{\widetilde{B}}$ $\leq\frac{1}{\Gamma(s)}\sum_{j=1}^{K}\int_{a_{j}}^{b_{j}}(t-r)^{s-1}\|\widetilde{S}_{n}(t-r)x_{j}-S(t-r)x_{j}\|_{\widetilde{B}}\mathop{}\!\mathrm{d}r$ (B.10)
$\leq\frac{1}{\Gamma(s)}\sum_{j=1}^{K}\int_{0}^{T}r^{s-1}\|\widetilde{S}_{n}(r)x_{j}-S(r)x_{j}\|_{\widetilde{B}}\mathop{}\!\mathrm{d}r$ (B.11)
$\leq\frac{T^{s}}{s\Gamma(s)}\sum_{j=1}^{K}\|\widetilde{S}_{n}\otimes x_{j}-S\otimes x_{j}\|_{C([0,T];\widetilde{B})}<\frac{\varepsilon}{2}.$ (B.12)
Since $t\in[0,T]$ was arbitrary, we conclude that ~$\|\widetilde{\mathfrak{I}}_{A_{n}}^{s}g-\mathfrak{I}_{A}^{s}g\|_{C([0,T];\widetilde{B})}<\frac{1}{2}\varepsilon$, and therefore ~$\|\widetilde{\mathfrak{I}}_{A_{n}}^{s}f-\mathfrak{I}_{A}^{s}f\|_{C([0,T];\widetilde{B})}<\varepsilon$ for all $n\geq N$. ∎
Appendix C Uniformly sectorial sequences of operators
We first recall the concept of sectorial operators. Given $\omega\in(0,\pi)$, we say that a linear operator $A\colon\mathsf{D}(A)\subseteq E\to E$ on a (real or complex) Banach space $E$, with spectrum $\sigma(A)\coloneqq\mathbb{C}\setminus\rho(A)$, is said to be $\omega$-sectorial if
$\sigma(A)\subseteq\overline{\Sigma}_{\omega}\quad\text{ and }\quad M(\omega,A)\coloneqq\sup\{\|\lambda R(\lambda,A)\|_{\mathscr{L}(E)}:\lambda\in\mathbb{C}\setminus\overline{\Sigma}_{\omega}\}<\infty,$ (C.1)
where $\Sigma_{\omega}$ is as in (3.14) and $M(\omega,A)$ is called the $\omega$-sectoriality constant. Its angle of sectoriality $\omega(A)\in[0,\pi)$ is defined as the infimum of all $\omega$ for which (C.1) holds.
If $A$ is closed and densely defined, then by [81, Theorem 13.30], there exists $\omega\in(0,\frac{1}{2}\pi)$ such that $A$ is $\omega$-sectorial if and only if there exists $\eta\in(0,\frac{1}{2}\pi)$ such that $-A$ generates a bounded analytic semigroup $(S(t))_{t\geq 0}$ on $\Sigma_{\eta}$. The latter means that the mapping $[0,\infty)\ni t\mapsto S(t)\in\mathscr{L}(E)$ extends to a bounded holomorphic function $\Sigma_{\eta}\ni z\mapsto S(z)\in\mathscr{L}(E)$. Inspecting the proof of the cited theorem reveals that, whenever these equivalent conditions hold, we have
$\sup\nolimits_{z\in\Sigma_{\eta}}\|S(z)\|_{\mathscr{L}(E)}\eqsim_{(\omega,\eta)}M(\omega,A).$ (C.2)
This theorem also asserts that the supremum of the set of $\eta\in(0,\frac{1}{2}\pi)$ for which $(S(t))_{t\geq 0}$ extends to a bounded analytic semigroup on $\Sigma_{\eta}$ equals $\frac{1}{2}\pi-\omega(A)$.
Moreover, by [41, Propositions 3.4.1 and 3.4.3] we have
$\|A^{\alpha}S(t)\|_{\mathscr{L}(E)}\lesssim_{(\omega,\alpha)}M(\omega,A)\,t^{-\alpha},$ (C.3)
for all $\omega\in(\omega(A),\frac{1}{2}\pi)$ and $t\in(0,\infty)$, where the implicit constant is non-decreasing in $\alpha$ for any fixed $\omega$.
We say that a sequence $(A_{n})_{n\in\mathbb{N}}$ of linear operators $A_{n}\colon\mathsf{D}(A_{n})\subseteq E_{n}\to E_{n}$ is uniformly sectorial of angle $\omega\in[0,\pi)$ if $A_{n}$ is sectorial of angle $\omega$ for all $n\in\mathbb{N}$ and
$M_{\text{Unif}}(\omega^{\prime},A)\coloneqq\sup\nolimits_{n\in\mathbb{N}}M(\omega^{\prime},A_{n})<\infty\quad\text{for all }\omega^{\prime}\in(\omega,\pi).$ (C.4)
Lemma C.1 complements Theorem 2.3 in the situation where the semigroup generators are uniformly sectorial of angle less than $\frac{1}{2}\pi$, in which case we obtain the uniform convergence $\Lambda_{n}A_{n}^{\alpha}S(\,\cdot\,)\Pi_{n}x\to A_{\infty}^{\alpha}S_{\infty}(\,\cdot\,)x$ on compact subsets of $(0,\infty)$. It is an analog to [55, Lemma 4.1(2)] in the discrete-to-continuum setting and for general $\alpha\in(0,\infty)$ (instead of $\alpha=1$).
Lemma C.1.
Let the linear operators $A_{n}\colon\mathsf{D}(A_{n})\subseteq E_{n}\to E_{n}$ on the Banach spaces $(E_{n},\|\,\cdot\,\|_{E_{n}})_{n\in\overline{\mathbb{N}}}$ be uniformly sectorial of angle $\omega\in[0,\frac{1}{2}\pi)$, and denote by $(S_{n}(t))_{t\geq 0}$ the bounded analytic $C_{0}$-semigroups generated by $-A_{n}$. Let Assumptions 2.1 and 2.2 be satisfied with $w=0$, and suppose that the equivalent statements a and b in Theorem 2.3 hold. Then we have, for all $\alpha\in(0,\infty)$, $x\in E_{\infty}$ and $0<a<b<\infty$,
$\sup_{t\in[a,b]}\|\Lambda_{n}A_{n}^{\alpha}S_{n}(t)\Pi_{n}x-A_{\infty}^{\alpha}S_{\infty}(t)x\|_{\widetilde{E}}\to 0\quad\text{as }n\to\infty.$ (C.5)
Proof.
Fix $\omega^{\prime}\in(\omega,\frac{1}{2}\pi)$, $n\in\overline{\mathbb{N}}$ and $t\in(0,\infty)$. We begin by sketching the functional calculus argument (see [50, Chapter 15] or [41] for a more comprehensive overview of this topic) which shows that we have the following Cauchy integral representation:
$A_{n}^{\alpha}S_{n}(t)=\frac{1}{2\pi i}\int_{\partial\Sigma_{\omega^{\prime}}}z^{\alpha}e^{-tz}R(z,A_{n})\mathop{}\!\mathrm{d}z.$ (C.6)
To see this, define the functions $f_{\alpha},g_{t}\colon\Sigma_{\omega^{\prime}}\to\mathbb{C}$ by $f_{\alpha}(z)\coloneqq z^{\alpha}$ and $g_{t}(z)\coloneqq e^{-tz}$ for $z\in\Sigma_{\omega^{\prime}}$ and $t\in(0,\infty)$. Denote by $f_{\alpha}(A_{n})$ and $g_{t}(A_{n})$ the operators obtained via the extended Dunford calculus for sectorial operators as defined in [50, Definition 15.1.8]. Then $f_{\alpha}(A_{n})$ is the fractional power $A_{n}^{\alpha}$ in the sense of [50, Definition 15.2.2], which satisfies $f_{\alpha}(A_{n})x=f_{\alpha}(\lambda)x=\lambda^{\alpha}x$ if $A_{n}x=\lambda x$, hence this (more general) definition agrees with the spectrally defined fractional powers in the setting of (3.13). Moreover, $g_{t}(A_{n})=S_{n}(t)$ by [50, Theorem 15.1.7]. Since $S_{n}(t)$ is bounded, we have $(f_{\alpha}g_{t})(A_{n})=f_{\alpha}(A_{n})g_{t}(A_{n})$ by [50, Proposition 15.1.12]. Finally, the function $(f_{\alpha}g_{t})(z)=z^{\alpha}e^{-z}$ is holomorphic and has (super)polynomial decay at 00 and $\infty$, and thus belongs to the domain of the primary Dunford calculus [50, Definition 15.1.1], so that $(f_{\alpha}g_{t})(A_{n})=\frac{1}{2\pi i}\int_{\partial\Sigma_{\omega^{\prime}}}z^{\alpha}e^{-tz}R(z,A_{n})\mathop{}\!\mathrm{d}z$. Putting all these observations together yields (C.6). Applying the projection and lifting operators and parametrizing the complex integral yields, for all $x\in E_{\infty}$,
[…]
where we recall that $\Pi_{\infty}=\Lambda_{\infty}=\operatorname{Id}_{\widetilde{E}}$ for $n=\infty$. It follows that the above estimate implies the following uniform bound on the interval $[a,b]$:
$\sup_{t\in[a,b]}\|\Lambda_{n}A_{n}^{\alpha}S_{n}(t)\Pi_{n}x-A_{\infty}^{\alpha}S_{\infty}(t)x\|_{\widetilde{E}}$ (C.8)
[…]
Setting $\eta\coloneqq\frac{1}{2}\pi-\vartheta$ for some $\vartheta\in(\omega,\omega^{\prime})$, the uniform sectoriality of $(A_{n})_{n\in\overline{\mathbb{N}}}$ implies that the operators $(-A_{n}e^{\pm i\eta})_{n\in\overline{\mathbb{N}}}$ generate $C_{0}$-semigroups $(S_{n}(te^{\pm i\eta}))_{t\geq 0}$ which are uniformly bounded in $t$ and $n$. Therefore, we can apply Theorem 2.3 to these two sequences of semigroups to find that $\widetilde{R}(\lambda,A_{n})x\to R(\lambda,A_{\infty})x$ for all $\lvert\arg\lambda\rvert>\vartheta$ if this convergence holds for one such $\lambda$. We have in fact $\widetilde{R}(\lambda,A_{n})x\to R(\lambda,A_{\infty})x$ for all $\lambda$ such that $\lvert\arg\lambda\rvert>\frac{1}{2}\pi>\vartheta$ by our hypothesis that the operators $(A_{n})_{n\in\overline{\mathbb{N}}}$ satisfy statements a and b in Theorem 2.3, so we conclude $\widetilde{R}(re^{\pm i\omega^{\prime}},A_{n})x\to R(re^{\pm i\omega^{\prime}},A_{\infty})x$ for all $r\in(0,\infty)$. On the other hand, by (C.4) and Assumption 2.1 we have $\|\widetilde{R}(re^{\pm i\omega^{\prime}},A_{n})x\|_{\widetilde{E}}\leq M_{\Pi}M_{\Lambda}M_{\text{Unif}}(\vartheta,A)\|x\|_{\widetilde{E}}/r$ for all $n\in\overline{\mathbb{N}}$. Hence, we can bound the integrand in (C.8), up to $n$-independent constants, by the integrable function $r\mapsto r^{\alpha-1}\exp(-a\cos(\omega^{\prime})r)$, so that the integrals tend to zero by the dominated convergence theorem. ∎